### v1

# Bayes Adaptive Monte Carlo Tree Search for Offline Model-based Reinforcement Learning

Abstract
Offline reinforcement learning (RL) is a powerful approach for data-driven decision-making and control. Compared to model-free methods, offline model-based reinforcement learning (MBRL) explicitly learns world models from a static dataset and uses them as surrogate simulators, improving the data efficiency and enabling the learned policy to potentially generalize beyond the dataset support. However, there could be various MDPs that behave identically on the offline dataset and so dealing with the uncertainty about the true MDP can be challenging. In this paper, we propose modeling offline MBRL as a Bayes Adaptive Markov Decision Process (BAMDP), which is a principled framework for addressing model uncertainty. We further introduce a novel Bayes Adaptive Monte-Carlo planning algorithm capable of solving BAMDPs in continuous state and action spaces with stochastic transitions. This planning process is based on Monte Carlo Tree Search and can be integrated into offline MBRL as a policy improvement operator in policy iteration. Our “RL + Search" framework follows in the footsteps of superhuman AIs like AlphaZero, improving on current offline MBRL methods by incorporating more computation input. The proposed algorithm significantly outperforms state-of-the-art model-based and model-free offline RL methods on twelve D4RL MuJoCo benchmark tasks and three target tracking tasks in a challenging, stochastic tokamak control simulator.
1 Introduction
The success of reinforcement learning (RL) typically relies on large amounts of interactions with the environment. However, in real-world scenarios, such interactions can be unsafe or costly. As an alternative, offline RL (Levine et al. (2020)) leverages offline datasets of transitions, collected by a behavior policy, to train a policy that can transfer to an online task. To avoid overestimation of the value function for some (out-of-sample) states in the environment, which can mislead policy learning, model-free offline RL methods (Kumar et al. (2020); Wu et al. (2019)) often constrain the learned policy to remain close to the behavior policy or within the support of the offline dataset. However, collecting transitions that comprehensively cover possible task scenarios, or acquiring a large volume of demonstrations from a high-quality behavior policy, can be expensive. This challenge has led to the development of offline model-based reinforcement learning (MBRL) approaches, such as (Lu et al. (2022); Guo et al. (2022)). These methods train dynamics models from offline data and optimize policies using imaginary rollouts generated by the models. Notably, the dynamics modeling is independent of the behavior policy, making it possible to achieve high returns even with data collected from a random policy. Furthermore, with careful dynamics modeling and thorough simulation, the learned policy can more effectively handle the environmental stochasticity and generalize to states beyond the support of the offline dataset.
Given a dataset, there may be various potential MDPs that behave identically on the limited set of states and actions, but their dynamics and reward functions could differ, especially on out-of-sample states and actions. This implies that we are dealing with a distribution of possible world models underlying the dataset. A common strategy in offline MBRL is to learn an ensemble of world models and treat them equally. For instance, when determining the next state, a world model is uniformly sampled from the ensemble and generate its prediction. However, different ensemble members may perform better in different regions of the state-action space, making it necessary to adapt the belief over each ensemble member based on the experience accumulated since the start of the episode. The Bayes Adaptive Markov Decision Process (BAMDP, Duff (2002)) provides a principled framework for modelling such an adaptive process. We show in Section 4.1 that, despite the need for Bayesian posterior updates, BAMDPs can still be efficiently simulated using deep ensembles. BAMCP (Guez et al. (2013)) is an efficient online planning method for solving BAMDPs. However, BAMCP has several limitations: (1) it relies on a ground-truth world model for planning; (2) it is restricted to discrete state and action spaces; and (3) its outcome is an action choice at a particular state, rather than a policy function. To address these challenges: (1) we apply a reward penalty (defined with the adapting belief) to construct a pessimistic BAMDP, preventing overexploitation of inaccurate world models (learned from the offline dataset); (2) we propose a novel planning algorithm to solve BAMDPs in continuous state and action spaces by extending BAMCP with double progressive widening (Auger et al. (2013)); and (3) we integrate the planning component as a policy improvement operator within policy iteration RL methods (Sutton & Barto (2018)), enabling the derivation of a policy suitable for real-time execution from the planning results. Specifically, the planning process is carried out through Monte Carlo Tree Search on a BAMDP. Integrating search with RL allows for the use of significantly more computation input, thereby improving policy learning performance. Grounded in the “scaling law", this paradigm has seen tremendous success in sophisticated policy learning, as demonstrated in (Silver et al. (2017b); Schrittwieser et al. (2020); AlphaProof & AlphaGeometry (2024)). Its application to offline MBRL, particularly in continuous control tasks, is a promising area for exploration.
To summarize, the main contributions of this work include: (1) Introducing BAMDPs to handle model uncertainties in offline MBRL; (2) Proposing an efficient Bayes Adaptive Monte Carlo Tree Search method for planning in continuous, stochastic BAMDPs; (3) Developing the first algorithm to successfully integrate Bayesian RL, offline MBRL, and deep search for sophisticated policy learning in continuous control under highly stochastic environments; (4) Demonstrating the improvements brought by Bayesian RL and deep search across twelve D4RL MuJoCo tasks and three target tracking tasks in a stochastic tokamak control scenario (for nuclear fusion), highlighting the potential of our algorithm to tackle challenging, real-world problems.
2 Background
A Markov Decision Process (MDP, Puterman (2014)) is described as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$. $\mathcal{S}$ and $\mathcal{A}$ are the state space and action space, respectively. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ is the dynamics function and $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{[0,1]}$ is the reward function, where $\Delta_{\mathcal{X}}$ denotes the set of possible probability distributions on $\mathcal{X}$. $\gamma\in[0,1)$ is a discount factor. The goal of solving an MDP is to find a policy $\pi:\mathcal{S}\rightarrow\Delta_{\mathcal{A}}$ that maximizes the expected return, defined as the cumulative discounted reward over time.
A Bayes Adaptive MDP (BAMDP, Duff (2002)) can model scenarios where the precise MDP $\mathcal{M}_{\theta}=\langle\mathcal{S},\mathcal{A},\mathcal{P}_{\theta},\mathcal{R}_{\theta},\gamma\rangle$ is uncertain but is known to follow a prior distribution $b_{0}(\theta)$. During planning, a Bayes-optimal agent would update its belief over the MDP based on experience. Formally, a BAMDP can be described as a tuple $\mathcal{M}^{+}=\langle\mathcal{S}^{+},\mathcal{A},\mathcal{P}^{+},\mathcal{R}^{+},\gamma\rangle$. $\mathcal{S}^{+}$ denotes the space of information states $(s,b)$, which is a composition of the physical state and the current belief over the MDP. After each transition $(s,a,r,s^{\prime})$, the belief is updated to the corresponding Bayesian posterior: $b^{\prime}(\theta)\propto b(\theta)P((s,a,r,s^{\prime})|\theta)=b(\theta)\mathcal{P}_{\theta}(s^{\prime}|s,a)\mathcal{R}_{\theta}(r|s,a)$. Accordingly, $\mathcal{P}^{+}$ and $\mathcal{R}^{+}$ can be defined as follows:
[…]
The Q-function that satisfies the Bellman optimality equations: $(\forall x=(s,b)\in\mathcal{S}^{+},a\in\mathcal{A})$
[…]
is the Bayes-optimal Q-function and $\pi^{*}(s,b)=\arg\max_{a}Q^{*}((s,b),a)$ is the Bayes-optimal policy. Actions derived from $\pi^{*}$ are executed in the real MDP and constitute the best course of action for a Bayesian agent with respect to its prior belief $b_{0}$ over the environment (Guez et al. (2014)). A BAMDP can be cast into a partially observable MDP (POMDP) [^1] Such a POMDP can be defined as a tuple $\langle\mathcal{S}^{+},\mathcal{A},\mathcal{P}^{+},\mathcal{R}^{+},\mathcal{Z},\mathcal{S},\gamma\rangle$, according to its original definition in (Littman (2009)). by viewing $\mathcal{S}^{+}$ and $\mathcal{S}$ as the state and observation spaces, respectively, and setting the observation function as $\mathcal{Z}((s^{\prime},b^{\prime}),a,o)=\mathds{1}(s^{\prime}=o)$. ($\mathcal{Z}$ specifies the probability of observing $o$ under the previous action $a$ and current state $(s^{\prime},b^{\prime})$.) As a result, approaches developed for POMDPs can potentially be used to solve BAMDPs.
Bayesian Reinforcement Learning (BRL, Ghavamzadeh et al. (2015)), as introduced above, is a principled approach to dealing with uncertainty in the world model $\mathcal{M}_{\theta}$ and has two main advantages: (1) Domain knowledge can be injected by defining a proper prior belief; (2) A Bayes Adaptive policy solves the exploration-exploitation dilemma by explicitly including the belief in its state representation and incorporating belief updates into the planning process (Sorg et al. (2010)). However, Bayes-optimal planning is generally intractable, for which we introduce some approximate methods in the next section.
3 Related Works
Offline RL (Levine et al. (2020); Chen et al. (2024)) enables an agent to learn control policies from datasets of environment transitions pre-collected by a behavior policy $\mu$, i.e., $\mathcal{D}_{\mu}=\{[(s_{t}^{i},a_{t}^{i},r_{t}^{i})_{t=1}^{T}]_{i=1}^{N}\}$, circumventing the need for potentially expensive or unsafe online interactions. Model-free offline RL methods directly learn value/policy functions from $\mathcal{D}_{\mu}$ but restrict the policy to remain close to $\mu$ or to behave within the support of $\mathcal{D}_{\mu}$. Offline Model-based RL (MBRL) methods, on the other hand, explicitly learn world models $\mathcal{M}_{\theta}$ from $\mathcal{D}_{\mu}$ (through supervised learning) and adopt $\mathcal{M}_{\theta}$ as a surrogate simulator, enabling the learned policy to possibly generalize to states beyond $\mathcal{D}_{\mu}$. Specifically, both planning methods (Argenson & Dulac-Arnold (2021); Zhan et al. (2022); Diehl et al. (2023)), such as MPC (Garcia et al. (1989)), and RL methods (Yu et al. (2020); Kidambi et al. (2020); Lu et al. (2022); Yu et al. (2021); Guo et al. (2022)) can be applied on top of the learned $\mathcal{M}_{\theta}$. However, since $\mathcal{D}_{\mu}$ may not span the entire state-action space, $\mathcal{M}_{\theta}$ is unlikely to be globally accurate. Learning/Planning without any safeguards against such model inaccuracy can yield poor results. In this case, the authors of (Yu et al. (2020); Kidambi et al. (2020); Lu et al. (2022)) propose learning an ensemble of world models, using ensemble-based uncertainty estimations to construct a pessimistic MDP (P-MDP), and learning a near-optimal policy atop it. Ideally, for any policy, the performance in the real environment is lower-bounded by the performance in the corresponding P-MDP (with high probability), thus avoiding being overly optimistic about an inaccurate model. For practical implementation of the P-MDP, Lu et al. (2022) conduct a thorough empirical study to compare a series of heuristics. However, none of these offline MBRL methods have modeled the problem as a BAMDP, even though Bayesian RL/planning provides a principled framework for handling model uncertainty. Additionally, with the learned model $\mathcal{M}_{\theta}$, deep and broad search in the policy learning process is possible, but the offline RL methods introduced above have not attempted such search-based policy improvement.
Monte-Carlo Tree Search (MCTS, Browne et al. (2012)) has been successfully integrated with RL, as exemplified by AlphaZero (Silver et al. (2017a)) and MuZero (Schrittwieser et al. (2020)). These methods have achieved superhuman performance in domains requiring highly precise and sophisticated decision-making processes. AlphaZero relies on given world models, whereas MuZero learns the world model and policy simultaneously by interacting with the environment. Although there have been various extensions of MuZero (Hubert et al. (2021); Schrittwieser et al. (2021); Ye et al. (2021); Danihelka et al. (2022); Antonoglou et al. (2022); Oren et al. (2022); Chen et al. (2023); Zhao et al. (2024)), most algorithms are designed for online MBRL. According to Niu et al. (2023), the applications of MuZero in offline learning, especially for continuous control in highly stochastic environments, which is our focus, still require significant improvement. Although search-based, MuZero learns a dynamics model and a reward model in a latent state space rather than the real state space, does not apply uncertainty estimation of the learned models as in usual offline MBRL methods, and does not adopt the Bayesian RL framework, making it fundamentally different from our algorithm design.
As mentioned in Section 2, Bayes-optimal planning is typically intractable. Approximate methods, such as (Asmuth et al. (2009); Sorg et al. (2010); Castro & Precup (2010); Asmuth & Littman (2011); Wang et al. (2012); Fonteneau et al. (2013); Guez et al. (2013); Slade et al. (2020)) have been developed. As a representative work, BAMCP (Guez et al. (2013)) adopts MCTS for Bayes Adaptive planning and is shown to converge in probability to a near Bayes-optimal policy at the root node of the search tree. However, all these methods cannot be directly applied to large-scale MDPs with continuous state and action spaces. Guez et al. (2014) and Hoang & Vien (2020) attempt to extend the use of BAMCP to scenarios with continuous states and actions using value/policy function approximations, but they replace MCTS with simple Monte Carlo rollouts (Bertsekas & Castañón (1999)) and do not utilize UCT (Kocsis & Szepesvári (2006)). Tree search makes branches at each node in the tree, providing a more thorough look-ahead search than simple rollouts, and UCT, as an efficient exploration algorithm, is the basis for the convergence guarantee of BAMCP. Moreover, these planning algorithms are not designed for offline MBRL. How to incorporate search-based planning in the policy improvement stage of RL and how to handle the uncertainty of the learned world model require further exploration, which are shown in the following section.
4 Methodology
In this work, we propose a novel offline MBRL algorithm based on Bayes Adaptive MCTS. The core challenge is to design a Bayes Adaptive planning method that is efficient in large stochastic MDPs. In this case, we propose Continuous BAMCP in Section 4.2, which can be applied to continuous control tasks with high dynamics stochasticity. Then, in Section 4.3, we present a search-based policy iteration framework, where the search results are distilled into policy and value networks for policy improvement and policy evaluation, respectively, at each iteration. In this way, we integrate offline MBRL with Bayes Adaptive MCTS. Both components require the use of an ensemble of world models for practical implementation and uncertainty quantification, which is detailed in Section 4.1.
4.1 The Key Role of Deep Ensembles
Offline MBRL methods estimate world models $\mathcal{M}_{\theta}$ from a static dataset $\mathcal{D}_{\mu}$, which would inevitably induce epistemic uncertainty about the identity of the real MDP $\mathcal{M}^{*}$. Specifically, there could be various potential MDPs that behave identically on the limited set of states and actions in $\mathcal{D}_{\mu}$, but their dynamics and reward functions may differ, especially on out-of-sample states and actions. Thus, we are actually dealing with a distribution of world models that follow a prior distribution $b_{0}(\theta)\triangleq P(\mathcal{M}_{\theta}|\mathcal{D}_{\mu})$. As introduced in Section 2, Bayesian RL based on BAMDP is a principled framework for handling model uncertainty by explicitly including the belief over the models in its state representation. Essentially, the belief is updated with experience, providing a measure of how the models’ uncertainty has changed since the beginning of the episode. As a result, the agent can adjust its behavior upon receiving new information that reduces the epistemic uncertainty. Such an adaptive policy is necessary to act optimally in offline RL, as demonstrated in (Ghosh et al. (2022)).
The idea of deep ensembles (Lakshminarayanan et al. (2017)) is to train multiple deep neural networks as approximations of a function, each using a different weight initialization and optimized with a different mini-batch sequence. For offline MBRL, we can learn an ensemble of dynamics models $\{\mathcal{P}_{\theta}^{1},\cdots,\mathcal{P}_{\theta}^{K}\}$ and reward models $\{\mathcal{R}_{\theta}^{1},\cdots,\mathcal{R}_{\theta}^{K}\}$ [^2] In some MBRL scenarios, a certain reward function is available, for instance, as defined by domain experts. Otherwise, the reward and dynamics function $(\mathcal{R}_{\theta}^{i},\mathcal{P}_{\theta}^{i})$ are usually trained as a unified probabilistic model $\mathcal{N}(\mu_{\theta}^{i},\sigma_{\theta}^{i})$, since the reward $r$ can be viewed as an element of the next state $s^{\prime}$. from the dataset $\mathcal{D}_{\mu}$ by minimizing the following supervised learning loss: ($i=1,\cdots,K$)
$\mathcal{L}(\mathcal{P}_{\theta}^{i})=-\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{\mu}}\left[\log\mathcal{P}_{\theta}^{i}(s^{\prime}|s,a)\right],\ \mathcal{L}(\mathcal{R}_{\theta}^{i})=-\mathbb{E}_{(s,a,r)\sim\mathcal{D}_{\mu}}\left[\log\mathcal{R}_{\theta}^{i}(r|s,a)\right]$ (3)
$\{(\mathcal{P}_{\theta}^{i},\mathcal{R}_{\theta}^{i})_{i=1}^{K}\}$ can be viewed as a set of independent and identically distributed (IID) samples from the prior $P(\mathcal{M}_{\theta}|\mathcal{D}_{\mu})$ and constitute a finite approximation of the space of world models. With such an ensemble, the belief over the world models can be converted to a mass function over a set of $K$ items, where the $i$-th element denotes the probability of being in the MDP $(\mathcal{P}_{\theta}^{i},\mathcal{R}_{\theta}^{i})$. In this case, a reasonable prior distribution is $b_{0}(\theta)=[1/K,\cdots,1/K]$, since these models are IID prior samples. After receiving a transition $(s,a,r,s^{\prime})$, the belief can be updated as follows:
$b^{\prime}(\theta)(i)=x^{i}/\sum_{j=1}^{K}x^{j},\ x^{i}=b(\theta)(i)\mathcal{P}_{\theta}^{i}(s^{\prime}|s,a)\mathcal{R}_{\theta}^{i}(r|s,a)$ (4)
This update requires a single inference from each ensemble member, but can be parallelized for computational efficiency. Equation (4) is a practical implementation of the Bayesian posterior update based on deep ensembles, and the simplification of the definition of $b(\theta)$ makes the transitions in Bayesian RL (i.e., Equation (1)) both practical and efficient.
The ensemble can also be used for uncertainty quantification. As aforementioned, our algorithm relies on thorough search on the learned world models. Without any constraints on the search process, the learned policy may overfit to an inaccurate model (by overestimating the expected return) and fail in the true MDP. Although the agent could adapt its belief and follow more reliable ensemble members in the Bayesian RL framework, there could be regions in the state-action space where none of the members generalize well, as they are all learned from a static offline dataset. A typical solution is to construct a P-MDP (see Section 3), which lower-bounds the true MDP and discourages the policy from regions where there is large discrepancy between the true and learned world models. We construct the P-MDP by modifying each reward estimation $r$ into $\tilde{r}$: ($\mu_{\theta}(s,a)=\sum_{i=1}^{K}b(\theta)(i)\mu_{\theta}^{i}(s,a)$)
$\tilde{r}(s,a,r,b(\theta))=r-\lambda\sqrt{\sum_{i=1}^{K}b(\theta)(i)(\sigma_{\theta}^{i}(s,a)^{2}+\mu_{\theta}^{i}(s,a)^{2})-\mu_{\theta}(s,a)^{2}}$ (5)
The reward penalty is weighted by a hyperparameter $\lambda>0$ and corresponds to the standard deviation (std) of the mixture of Gaussian dynamics models, where $\mu_{\theta}^{i}$ and $\sigma_{\theta}^{i}$ are the mean and std from the ensemble member $i$. This penalty design combines epistemic and aleatoric model uncertainty and has been shown to be successful at capturing errors in predicted dynamics (Lu et al. (2022)) [^3] In the original literature (Lakshminarayanan et al. (2017); Lu et al. (2022)), the ensemble is treated as a uniformly-weighted mixture model, i.e., $b(\theta)(i)=1/K,\ (i=1,\cdots K)$, since their belief is not adaptive. Equation (5) fits into the Bayesian RL framework by adapting the belief $b(\theta)$, which is part of our novelty..
Algorithm 1 Continuous BAMCP
Input: $\pi$, $V$, $E$, $d_{\text{max}}$, $\gamma$, $\alpha$, $\beta$, $\mathcal{P}_{\theta}^{1:K}$, $\mathcal{R}_{\theta}^{1:K}$
procedure Search($(s,h),b(\theta)$)
     for $e=1\cdots E$ do
         Simulate($(s,h),b(\theta),d_{\text{max}}$)
     end for
     $v_{\text{ret}}=\sum_{a\in C((s,h))}\frac{N((s,h),a)}{N((s,h))}Q((s,h),a)$
      return $\pi_{\text{ret}},v_{\text{ret}}$
end procedure
procedure Simulate($(s,h),b(\theta),d$)
     if $d==0$ then return $V((s,h))$
     $a\leftarrow\textsc{ActionPW}((s,h))$
     $r,s^{\prime},b^{\prime}(\theta)\leftarrow\textsc{StatePW}((s,h),b(\theta),a)$
[…]
     if $N((s,h))>1$ then
         $R\leftarrow\textsc{Simulate}((s^{\prime},hars^{\prime}),b^{\prime}(\theta),d-1)$
     else
         $R\leftarrow V((s^{\prime},hars^{\prime}))$
     end if
     Access $\tilde{r}$ or calculate $\tilde{r}$ using Eq. (5)
     $R\leftarrow\tilde{r}+\gamma R$, cache $\tilde{r}$
     $Q((s,h),a)\mathrel{{+}{=}}\frac{R-Q((s,h),a)}{N((s,h),a)}$
     return $R$
end procedure
procedure ActionPW($(s,h)$)
     if first visit then $C((s,h))\leftarrow\emptyset$
     if $\left\lfloor{N((s,h))^{\alpha}}\right\rfloor\geq|C((s,h))|$ then
         $a\sim\pi(\cdot|(s,h))$
         $C((s,h))\leftarrow C((s,h))\cup\{a\}$
         $N((s,h),a),Q((s,h),a)\leftarrow 0,0$
     else
         $a\leftarrow\arg\max_{x\in C((s,h))}\tilde{Q}((s,h),x)$
     end if
     return $a$
end procedure
procedure StatePW($(s,h),b(\theta),a$)
     if first visit then $C((s,h),a)\leftarrow\emptyset$
     if $\left\lfloor{N((s,h),a)^{\beta}}\right\rfloor\geq|C((s,h),a)|$ then
         $r\sim\sum_{i=1}^{K}b(\theta)(i)\mathcal{R}_{\theta}^{i}(\cdot|s,a)$
         $s^{\prime}\sim\sum_{i=1}^{K}b(\theta)(i)\mathcal{P}_{\theta}^{i}(\cdot|s,a)$
         Update $b(\theta)$ to $b^{\prime}(\theta)$ using Eq. (4)
         $C((s,h),a)\leftarrow C((s,h),a)\cup\{(r,s^{\prime},b^{\prime}(\theta))\}$
         $N((s^{\prime},hars^{\prime}))\leftarrow 0$
         return $r,s^{\prime},b^{\prime}(\theta)$
     end if
     return the least visited node in $C((s,h),a)$
end procedure
4.2 Bayes Adaptive MCTS in Continuous State and Action Spaces
BAMCP (Guez et al. (2013)) has been successful in solving large-scale BAMDPs, as detailed in Appendix A, but it is limited to scenarios with discrete state and action spaces. In this subsection, we introduce an online planning method to approximate the Bayes-optimal policy at a decision point $(s,h)$. ($h$ denotes the transition history that ends at $s$.) In particular, this method can be used to solve BAMDPs with continuous states and actions. Specifically, we adopt double progressive widening (DPW, Couëtoux et al. (2011); Auger et al. (2013)), which maintains a finite list of chance nodes to be searched at each decision point and incrementally adds a new chance node to the list based on the visitation counts. For a node $(s,h)$, a new action $a$ is sampled (with the current policy) and added to its children set $C((s,h))$, if $\left\lfloor{N((s,h))^{\alpha}}\right\rfloor\geq|C((s,h))|$, where $\alpha\in(0,1)$ is a hyperparameter that controls the growth rate and $N$ denotes the visitation counts. Otherwise, an action will be sampled from $C((s,h))$ according to the UCT (Kocsis & Szepesvári (2006)) rule. Similarly, to handle the infinitely many possible transitions, a new next state $s^{\prime}$ is added to the children set $C((s,h),a)$ only if $\left\lfloor{N((s,h),a)^{\beta}}\right\rfloor\geq|C((s,h),a)|$ ($\beta\in(0,1)$). Otherwise, the least visited child in $C((s,h),a)$ will be selected as the next state. With DPW, the sets of possible actions or next states for search are all finite, allowing deep search as in discrete scenarios, and the more promising states and actions (with higher $N$) have more subsequent branches, thereby reducing corresponding estimation uncertainty. However, directly integrating DPW and BAMCP (i.e., Algorithm 3) likely cannot solve BAMDPs with continuous state and action spaces. As introduced in Appendix A, BAMCP relies on root sampling, which samples dynamics functions (from $b(\theta)$) only at the root node and follows a specific dynamics function throughout a simulation procedure. However, the rationale of root sampling (i.e., Lemma A.1) does not hold when applying DPW [^4] The last equality of Eq. (6) does not hold, since $\tilde{b}(\theta|has^{\prime})\propto\tilde{b}(\theta|ha)\widetilde{\mathcal{P}}_{\theta}(s^{\prime}|s,a)\neq\tilde{b}(\theta|ha)\mathcal{P}_{\theta}(s^{\prime}|s,a)$. $\widetilde{\mathcal{P}}_{\theta}(s^{\prime}|s,a)$ represents the state transition distribution when applying DPW, which differs from the true distribution $\mathcal{P}_{\theta}(s^{\prime}|s,a)$, as dictated by the DPW rule.. Also, BAMCP assumes that the reward function is certain rather than a function of $\theta$.
On the other hand, Polynomial Upper Confidence Trees (PUCT, Auger et al. (2013)), based on DPW, is a provably consistent planning method for solving MDPs with infinite-scale state/action spaces and arbitrarily stochastic transition kernels. In this case, we propose casting BAMDPs into MDPs (i.e., $\mathcal{M}^{+}$ defined in Section 2) and solving them with PUCT. The pseudo code is shown as Algorithm 1. Ideally, as the number of simulations $E\rightarrow\infty$, PUCT can find a near-optimal solution of $\mathcal{M}^{+}$, which is also a near Bayes-optimal solution for the true environment. Each simulation follows a path that begins at the root node and ends at a leaf node, utilizing progressive widening for sampling both actions and next states, as illustrated in the ActionPW and StatePW procedures. Compared to PUCT, the significant modifications include: (1) replacing $\langle\mathcal{S},\mathcal{P},\mathcal{R}\rangle$ with their extended definitions in BAMDPs, $\langle\mathcal{S}^{+},\mathcal{P}^{+},\mathcal{R}^{+}\rangle$, and (2) applying reward penalties to account for model uncertainty. To be specific, in StatePW, $r$ and $s^{\prime}$ are sampled from a mixture of ensemble members, which is a practical implementation of sampling from $\mathcal{R}^{+}$ and $\mathcal{P}^{+}$ as shown in Eq. (1). After receiving the transition $(s,a,r,s^{\prime})$, the belief vector $b(\theta)$ is updated to $b^{\prime}(\theta)$ following Eq. (4), finishing the transition in $\mathcal{S}^{+}$ from $(s,b(\theta))$ to $(s,b^{\prime}(\theta))$. Moreover, the reward $r$ is adjusted with a penalty term as defined in Eq. (5), which is then used to calculate the return $R$. As aforementioned, applying such a reward penalty can effectively mitigate the issue of model exploitation.
[…]
4.3 The Overall Framework: Search-based Policy Iteration
In this subsection, we present a framework (inspired by MuZero [^6] The novelty of our algorithm in comparison to MuZero is detailed in Section 3.) that integrates continuous BAMCP (i.e., Algorithm 1) into policy improvement and policy evaluation. By iteratively running these procedures, we can approach a near Bayes-optimal policy, i.e., $\pi$, that can be directly referred to during execution in the true environment. The pseudo code of the overall framework is shown as Algorithm 2. For efficiency, a learner and a number of actors execute in parallel, reading from and sending data to the replay buffer $\mathcal{D}$ respectively. The actors update their copies of policy and value functions every $E_{l}$ learner steps.
In particular, each actor would interact with the learned world models to sample trajectories $\tau$ [^7] In Algorithm 2, $\hat{\rho}(s)$ denotes the empirical distribution of initial states in the offline dataset $\mathcal{D}_{\mu}$. At each time step of the trajectory, a Search procedure (defined in Algorithm 1) is executed using the actor’s current copy of $\pi$ and $V$ for planning at the current decision point. The search result $\pi_{\text{ret}}$ is then used to indicate the action choice, i.e., $a\sim\pi_{\text{ret}}(\cdot|(s,h))$. The subsequent transition process follows a BAMDP, where $r\sim\mathcal{R}^{+}(\cdot|(s,h),a)$, $s^{\prime}\sim\mathcal{P}^{+}(\cdot|(s,h),a)$, and the belief $b({\theta})$ adapts with experience, as described in StatePW. The collected trajectories are used in the learning process, where $\pi$ is trained to mimic the search result $\pi_{\text{ret}}$ by minimizing a cross-entropy loss (i.e., $\mathcal{L}(\pi,\{\tau_{i}\}_{i=1}^{B})$), while $V$ is trained to predict the $n$-step pessimistic return $z$ [^8] In the definition of $z$, $\tilde{r}^{(0)}=\tilde{r}$ is the pessimistic reward associated with $(s,h)$ and the subscript $(i)$ denotes that the corresponding quantity is collected in $i$ time steps within that trajectory. As noted in (Hubert et al. (2021)), $\pi_{\text{ret}}$ improves $\pi$ at each decision point, so repeatedly applying continuous BAMCP to obtain $\pi_{\text{ret}}$ and projecting the search results to the parameter space of $\pi$ (by minimizing $\mathcal{L}(\pi,\{\tau_{i}\}_{i=1}^{B})$) represents the policy improvement process. Meanwhile, value backups to update the Q estimates (i.e., $R\leftarrow V((s^{\prime},hars^{\prime}))$ in Algorithm 1) and the temporal difference learning for the value function $V$ constitute the policy evaluation process. In this way, the search algorithm is used as a powerful improvement operator to iteratively enhance the performance of the learned policy $\pi$.
Algorithm 2 Search-based Policy Iteration
Input: $T$, $E_{l}$, $\mathcal{P}_{\theta}^{1:K}$, $\mathcal{R}_{\theta}^{1:K}$
Initialize $\pi$ and $V$, $\mathcal{D}\leftarrow\emptyset$
procedure Learner
     $e\leftarrow 0$
     while true do
         $\{\tau_{i}\}_{i=1}^{B}\sim\mathcal{D}$
         $\min_{\pi}\mathcal{L}(\pi,\{\tau_{i}\}_{i=1}^{B})$, $\min_{V}\mathcal{L}(V,\{\tau_{i}\}_{i=1}^{B})$
[…]
         Update $\pi,V$ in Actor if $e\%E_{l}==0$
     end while
end procedure
$\mathcal{L}(\pi,\{\tau_{i}\}_{i=1}^{B})=-\sum_{((s,h),\pi_{\text{ret}})}\pi_{\text{ret}}^{T}\log\pi(\cdot|(s,h))/(BT)$
$\mathcal{L}(V,\{\tau_{i}\}_{i=1}^{B})=\sum_{((s,h),z)}(V(s,h)-z)^{2}/(BT)$,
$z=\tilde{r}^{(0)}+\gamma\tilde{r}^{(1)}+\cdots+\gamma^{n-1}\tilde{r}^{(n-1)}+\gamma^{n}v_{\text{ret}}^{(n)}$
procedure Actor
     while true do
         $s\leftarrow\hat{\rho}_{0}(\cdot)$, $h\leftarrow s$, $\tau\leftarrow[]$
         $b(\theta)\leftarrow[1/K,\cdots,1/K]$
         for $t=1\cdots T$ do
              $\pi_{\text{ret}},v_{\text{ret}}\leftarrow\textsc{Search}((s,h),b(\theta))$
              $a\sim\pi_{\text{ret}}(\cdot|(s,h))$
              Acquire $r,s^{\prime},b^{\prime}(\theta)$ as in StatePW
              Calculate $\tilde{r}$ using Eq. (5)
              Append $\tau$ with $((s,h),\tilde{r},\pi_{\text{ret}},v_{\text{ret}})$
              $s,h,b(\theta)\leftarrow s^{\prime},hars^{\prime},b^{\prime}(\theta)$
         end for
         $\mathcal{D}\leftarrow\mathcal{D}\cup\{\tau\}$
     end while
end procedure
5 Evaluation
To evaluate the effectiveness of each component in our algorithm design, we introduce three variants: (1) BA-MBRL leverages learned world models as surrogate simulators, applying reward penalties to collected transitions and using standard online RL algorithms (e.g., SAC (Haarnoja et al. (2018))) to learn a policy. While following existing offline MBRL methods, it models the problem as a BAMDP (rather than an MDP), with environment transitions defined by Equations (1) and (4) and the reward penalty by Equation (5), and is designed to evaluate the effectiveness of Bayesian RL. (2) BA-MCTS builds on BA-MBRL by introducing Continuous BAMCP (Algorithm 1) to plan at decision points, rather than inferring directly from the policy, to generate trajectories for downstream SAC, demonstrating the impact of deep search on policy learning. (3) BA-MCTS-SL, described in Algorithm 2, replaces the policy learning algorithm in BA-MCTS from policy gradient methods (as in SAC) with supervised learning (SL), allowing us to compare which approach offers a more efficient policy update mechanism, particularly for continuous control tasks.
Data Type Environment BA-MCTS -SL (ours) BA-MCTS (ours) BA-MBRL (ours) Optimized COMBO MOReL MOPO
random HalfCheetah 29.20 $\pm$ 2.00 36.23 $\pm$ 1.04 32.76 $\pm$ 1.16 31.7 38.8 25.6 35.4
random Hopper 33.83 $\pm$ 0.10 31.56 $\pm$ 0.12 31.47 $\pm$ 0.03 12.1 17.9 53.6 11.7
random Walker2d 21.89 $\pm$ 0.07 21.59 $\pm$ 0.32 21.45 $\pm$ 0.53 21.7 7.0 37.3 13.6
medium HalfCheetah 70.47 $\pm$ 3.52 75.84 $\pm$ 3.81 56.54 $\pm$ 5.20 45.7 54.2 42.1 42.3
medium Hopper 97.75 $\pm$ 7.09 96.70 $\pm$ 14.0 98.25 $\pm$ 3.42 69.3 97.2 95.4 28.0
medium Walker2d 82.24 $\pm$ 1.85 74.73 $\pm$ 3.25 75.41 $\pm$ 4.17 79.7 81.9 77.8 17.8
med-replay HalfCheetah 61.16 $\pm$ 1.60 65.45 $\pm$ 0.81 62.50 $\pm$ 0.18 58.0 55.1 40.2 53.1
med-replay Hopper 106.3 $\pm$ 0.13 101.8 $\pm$ 3.46 93.91 $\pm$ 4.25 90.8 89.5 93.6 67.5
med-replay Walker2d 92.13 $\pm$ 5.13 95.06 $\pm$ 2.11 97.54$\pm$ 1.93 65.8 56.0 49.8 39.0
med-expert HalfCheetah 80.53 $\pm$ 6.63 76.16 $\pm$ 10.3 90.52 $\pm$ 4.13 104.2 90.0 53.3 63.3
med-expert Hopper 112.2 $\pm$ 0.29 108.3 $\pm$ 0.22 107.8 $\pm$ 0.37 105.8 111.1 108.7 23.7
med-expert Walker2d 107.7 $\pm$ 0.82 110.0 $\pm$ 1.74 84.71 $\pm$ 0.87 97.1 103.3 95.6 44.6
Average Score 74.62 74.45 71.06 65.16 66.83 64.42 36.67
Table 1: Comparisons between the proposed algorithms and SOTA offline model-based RL methods on the D4RL benchmark suite. Each value represents the normalized score, as proposed in (Fu et al. (2020)), of the policy trained by the corresponding algorithm. These scores are undiscounted returns normalized to approximately range between 0 and 100, where a score of 0 corresponds to a random policy and a score of 100 corresponds to an expert-level policy. For our algorithms, we report the average score of the final ten policy learning epochs and its standard deviation across three random seeds. Results in the last four columns are taken from the original papers (Lu et al. (2022); Yu et al. (2021); Kidambi et al. (2020); Yu et al. (2020)), respectively.
We first evaluate our algorithms on a widely-used continuous control benchmark for offline RL methods – D4RL MuJoCo (Fu et al. (2020)). The evaluation results for three types of robotic agents, each with offline datasets of four different qualities, are presented in Table 1. (1) Compared to SOTA offline MBRL methods, our algorithms achieve superior performance on nine out of twelve tasks. In terms of average performance, BA-MBRL significantly outperforms the baselines, demonstrating the effectiveness of using BAMDPs to handle model uncertainties in offline MBRL. (2) Both BA-MCTS and BA-MCTS-SL further improve upon BA-MBRL, highlighting the enhancement brought by deep search in policy learning. Notably, we apply Continuous BAMCP to only 10% of states when collecting training trajectories, while for the remaining states, we sample actions directly from the policy, i.e., $a\sim\pi(\cdot|s)$. Increasing the search ratio could further enhance policy performance at the cost of increased computation. (3) BA-MCTS-SL performs similarly to BA-MCTS, validating the effectiveness of both policy update mechanisms. However, BA-MCTS-SL struggles on Walker2d, where a warm-up training phase (using BA-MBRL) is required to establish a better initial policy. On the other hand, the advantage of the SL-based policy update is evident in the training plots of our algorithms in Figure 3, where BA-MCTS-SL exhibits much smoother learning curves compared to the other two algorithms, indicating greater robustness in model selection. (4) We further compare our algorithms with model-free offline policy learning methods, as shown in Appendix D. The performance improvement is even greater than that over model-based methods, highlighting the necessity of model-based learning. Particularly, when data quality is low, merely mimicking or staying close to the behavior policy would result in an underperforming policy. (5) We provide an ablation study in Appendix F to demonstrate the necessity of incorporating the reward penalty in offline MBRL to prevent the overexploitation of inaccurate world models. Additionally, we find that the SL-based policy update is less sensitive to model inaccuracies.
(a) hc-med-expert
(b) hc-med-replay
(c) hc-medium
(d) hc-random
(e) hp-med-expert
(f) hp-med-replay
(g) hp-medium
(h) hp-random
(i) wk-med-expert
(j) wk-med-replay
(k) wk-medium
(l) wk-random
Figure 1: Performance of Sampled EfficientZero on D4RL MuJoCo tasks. The results for HalfCheetah, Hopper, and Walker2d are presented in the three rows, respectively. Each subfigure depicts the change in undiscounted episodic return as a function of the number of training samples. Experiments are repeated three times with different random seeds, with the solid line representing the mean and the shaded area indicating the 95% confidence interval. For reference, the expert-level episodic returns for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively.
Figure 2: Evaluation results for the tokamak control tasks. The figure shows the change in episodic returns over training epochs for the proposed algorithms and baselines across three target tracking tasks in the nuclear fusion scenario. Solid lines represent the average performance, while shaded areas indicate the 95% confidence intervals.
For fair comparisons and real-time execution, we do not perform test-time search and adopt the same policy architecture as the baselines, i.e., a feedforward neural network, rather than an RNN that incorporates transition history as input. These alternative designs have the potential to further improve our algorithms. For implementation, we build on the codebase of Optimized (Lu et al. (2022)), which thoroughly explores design choices in offline MBRL, making minimal changes to the code and hyperparameter settings [^9] The detailed hyperparameter setups of our algorithms are provided in Appendix C.. Therefore, we believe the performance improvements stem from the Bayesian RL framework and deep search component. Further, both components can be seamlessly integrated with other advancements in offline MBRL, such as more accurate world model learning and improved uncertainty quantification for constructing pessimistic MDPs.
MuZero also applies deep search to MBRL. To evaluate its performance on D4RL MuJoCo tasks, we use the open-source implementation and hyperparameter configurations of Sampled EfficientZero (Ye et al. (2021)) provided by LightZero (Niu et al. (2023)). Benchmarking results from LightZero indicate that Sampled EfficientZero, equipped with a Gaussian policy, achieves the best performance on (online) MuJoCo locomotion tasks compared to other MuZero variants. To adapt Sampled EfficientZero for offline learning, we employ the reanalyse technique proposed by (Schrittwieser et al. (2021)). The evaluation results are presented in Figure 1. For reference, the expert-level episodic returns (corresponding to scores of 100) for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively. As shown, the results are significantly worse compared to the performance of offline RL methods listed in Table 1, despite Sampled EfficientZero’s higher computational cost (detailed in Appendix E). Notably, both Sampled EfficientZero and BA-MCTS-SL rely on supervised learning for policy improvement. However, for continuous control tasks, the agent can only sample a finite number of actions at a decision point, and the search result (e.g., $\pi_{\text{ret}}$ in Algorithm 1) is a distribution over this finite set, which could be a poor approximation of the optimal action distribution. Thus, purely mimicking the search result may be less sample-efficient than policy gradient methods, as it fails to account for the continuous nature of the action space. Furthermore, world model learning is the foundation of MBRL and can be particularly challenging in continuous control and offline learning settings, where the state-action space is vast but training data is limited. Sampled EfficientZero integrates model learning and policy training into a single stage, which significantly increases the learning difficulty (compared to BA-MCTS-SL).
Finally, we evaluate our algorithms on three target tracking tasks in tokamak control. The tokamak is one of the most promising confinement devices for achieving controllable nuclear fusion, where the primary challenge lies in confining the plasma, i.e., an ionized gas of hydrogen isotopes, while heating it and increasing its pressure to initiate and sustain fusion reactions (Pironti & Walker (2005)). Tokamak control involves applying a series of direct actuators (e.g., neutral beam, ECH power, magnetic field) and indirect actuators (e.g., setting targets for the plasma shape and density) to confine the plasma to achieve a desired state or track a given target. This sophisticated physical process is an ideal test bed for our algorithms. Specifically, we use a well-trained data-driven dynamics model provided by Char et al. (2024) as a “ground truth" simulator for the nuclear fusion process during evaluation, and generate a dataset containing 725270 transitions using this model for offline RL. We select a reference shot (i.e., an episode of a fusion process) from DIII-D [^10] DIII-D is a tokamak device located in San Diego, California, operated by General Atomics., and use its trajectories of Ion Rotation, Electron Temperature, and $\beta_{n}$ as targets for three tracking tasks. These are critical quantities in tokamak control, particularly $\beta_{n}$, which serves as an economic indicator of the efficiency of nuclear fusion. The tracking tasks have a 28-dimensional state space and a 14-dimensional action space, both continuous. Moreover, these tasks are highly stochastic, as the underlying dynamics model is a probabilistic neural network and each state transition is a sample from this model. For details on the simulator, and the design of the state/action spaces and reward functions, please refer to Appendix G. We compare our algorithms with SOTA model-free and model-based offline RL methods, specifically CQL and Optimized. The learning performance on the three tracking tasks is shown in Figure 2, where the x-axis and y-axis represent the training epochs and (negative) full-shot tracking errors, respectively. Our algorithms consistently outperform the baselines. Notably, the offline dataset does not include the reference shot or any similar, nearby shots. Therefore, restricting the policy to stay close to the behavior policy, as done in model-free methods, can be problematic. Also, learning dynamics models for MBRL is quite challenging in this nuclear fusion scenario. Our algorithms share the same ensemble of dynamics models with “Optimized" for policy learning, and the comparisons can demonstrate the superiority of Bayesian RL and deep search. Figure 2 has been smoothed for visualization [^11] The episodic return is plotted every 10 training epochs, with the y-axis representing the average value of a sliding window of length 5.. We further report the average return over the final 10 training epochs in Table 2, and the conclusions align with those from the D4RL MuJoCo tasks, showing the robustness of our proposed algorithms.
Task BA-MCTS -SL (ours) BA-MCTS (ours) BA-MBRL (ours) CQL Optimized
Temperature -21.16 $\pm$ 5.00 -23.83 $\pm$ 9.66 -29.35 $\pm$ 4.72 -59.62 $\pm$ 1.57 -83.55 $\pm$ 10.56
Rotation -14.14 $\pm$ 1.88 -19.07 $\pm$ 5.85 -31.33 $\pm$ 11.54 -85.48 $\pm$ 2.72 -71.54 $\pm$ 9.88
$\beta_{n}$37.03 $\pm$ 17.98 -18.93 $\pm$ 1.75 -23.4 $\pm$ 10.77 -36.37 $\pm$ 1.17 -57.84 $\pm$ 10.27
Average -24.11 -20.61 -28.03 -60.49 -70.98
Table 2: Comparisons between the proposed algorithms and offline RL baselines on the target tracking tasks. For each algorithm, we report the average return of the final ten policy learning epochs and its standard deviation across three different random seeds.
6 Conclusion and Discussions
In this work, we propose framing offline model-based reinforcement learning (MBRL) as a Bayes Adaptive Markov Decision Process (BAMDP) to better address uncertainties in the world models learned from offline datasets. We also introduce a novel planning method for solving BAMDPs in continuous state and action spaces using Monte Carlo Tree Search. This planning process is integrated into a policy iteration framework, enabling the derivation of a policy suitable for real-time execution from the planning results. In our evaluation, we test several variants of our algorithms to separately highlight the effectiveness of Bayesian RL and deep search. Additionally, we compare two different approaches for policy updates (based on the search results) in continuous control tasks: supervised learning and policy gradient methods. Our findings demonstrate that: (1) adapting beliefs over an ensemble of world models based on experience yields more accurate model approximations for MBRL; (2) deep search improves learning performance by incorporating planning and additional computation input; and (3) while supervised-learning-based policy updates result in smoother learning curves, they may struggle in complex continuous control tasks due to their approximation of the continuous action space as a finite set of action samples. For future work, our algorithms can be further improved by integrating advancements in offline MBRL and Bayesian RL, such as Bayesian RL methods that do not rely on deep ensembles, techniques to address sparse rewards in MBRL, and more principled approaches to construct pessimistic MDPs beyond those based on ensemble discrepancy.
Appendix A BAMCP
Algorithm 3 BAMCP
Input: $\pi$, $E$, $d_{\text{max}}$, $\mathcal{R}$, $\gamma$, $\mathcal{A}$, $c$
procedure Search($(s,h),b(\theta)$)
     for $e=1\cdots E$ do
         $\theta\sim b(\cdot)$
         Simulate($(s,h),\theta,d_{\text{max}}$)
     end for
     return $\arg\max_{a}Q((s,h),a)$
end procedure
procedure Rollout($(s,h),\theta,d$)
     if $d==0$ then
         return 0
     end if
     $a\sim\pi(\cdot|(s,h))$
     $s^{\prime}\sim\mathcal{P}_{\theta}(\cdot|s,a)$
     $r\leftarrow\mathcal{R}(s,a)$
     $R\leftarrow r+\gamma\textsc{Rollout}((s^{\prime},has^{\prime}),\theta,d-1)$
     return $R$
end procedure
procedure Simulate($(s,h),\theta,d$)
     if $d==0$ then
         return 0
     end if
     if $N((s,h))==0$ then
         for $a\in\mathcal{A}$ do
              $N((s,h),a),Q((s,h),a)\leftarrow 0,0$
         end for
         $a\sim\pi(\cdot|(s,h))$
         $s^{\prime}\sim\mathcal{P}_{\theta}(\cdot|s,a),r\leftarrow\mathcal{R}(s,a)$
         $R\leftarrow r+\gamma\textsc{Rollout}((s^{\prime},has^{\prime}),\theta,d-1)$
         $N((s,h)),N((s,h),a)\leftarrow 1,1$
         $Q((s,h),a)\leftarrow R$
         return $R$
     end if
     $a\leftarrow\arg\max_{x}Q((s,h),x)+c\sqrt{\frac{\log N((s,h))}{N((s,h),x)}}$
     $s^{\prime}\sim\mathcal{P}_{\theta}(\cdot|s,a),r\leftarrow\mathcal{R}(s,a)$
     $R\leftarrow r+\gamma\textsc{Simulate}((s^{\prime},has^{\prime}),\theta,d-1)$
[…]
     $Q((s,h),a)\mathrel{{+}{=}}\frac{R-Q((s,h),a)}{N((s,h),a)}$
     return $R$
end procedure
Bayes Adaptive Monte Carlo Planning (BAMCP, Guez et al. (2013)) is a sample-based online planning method, aiming to find the action $a^{*}$ that approximately maximizes the expected return at a decision point $(s,h)$ under the BAMDP. Its detailed pseudo code is shown as Algorithm 3. BAMCP has demonstrated success in solving BAMDPs with large-scale discrete state and action spaces. Its key algorithmic ideas include: (1) applying MCTS with an efficient exploration strategy – UCT (Kocsis & Szepesvári (2006)) to the BAMDP in order to simulate the outcomes of different action choices; (2) utilizing root sampling to avoid frequent Bayesian posterior updates. Specifically, the UCT rule is used for selecting actions at non-leaf nodes, i.e., $a\leftarrow\arg\max_{x}Q((s,h),x)+c\sqrt{\frac{\log N((s,h))}{N((s,h),x)}}$, managing the tradeoff between exploration and exploitation. Root sampling refers to sampling the dynamics model only at the root node (i.e., $\theta\sim b(\cdot)$) and not adapting the belief $b(\cdot)$ according to the Bayes rule during the search process, of which the rationality is justified in the following lemma.
Lemma A.1.
For all suffix histories $h^{\prime}$ of $h$, $b(\theta|h^{\prime})=\tilde{b}(\theta|h^{\prime})$. Here, $b(\theta|h^{\prime})$ is the true posterior probability of $\theta$ at the decision point $h^{\prime}$, while $\tilde{b}(\theta|h^{\prime})$ is the probability of experiencing $\theta$ at $h^{\prime}$ when using root sampling.
Proof.
This lemma can be proved by induction.
Base case: When $h^{\prime}=h$, $b(\theta|h^{\prime})=\tilde{b}(\theta|h^{\prime})=b(\theta)$.
Step case:
$b(\theta|has^{\prime})$ $=P(has^{\prime}|\theta)P(\theta)/P(has^{\prime})$ (6)
$=P(h|\theta)\mathcal{P}_{\theta}(s^{\prime}|s,a)P(\theta)/P(has^{\prime})$
$=b(\theta|h)P(h)\mathcal{P}_{\theta}(s^{\prime}|s,a)/P(has^{\prime})$
$=Zb(\theta|h)\mathcal{P}_{\theta}(s^{\prime}|s,a)$
$=Z\tilde{b}(\theta|h)\mathcal{P}_{\theta}(s^{\prime}|s,a)$
$=Z\tilde{b}(\theta|ha)\mathcal{P}_{\theta}(s^{\prime}|s,a)=\tilde{b}(\theta|has^{\prime})$
[…]
Appendix B Alternative Design Choices for Continuous BAMCP
The terms labeled in blue in Algorithm 1 have alternative design choices. Empirical comparisons among these alternatives are reserved for future work.
$\tilde{Q}((s,h),x)$ in the exploration strategy, i.e., $a\leftarrow\arg\max_{x\in C((s,h))}\tilde{Q}((s,h),x)$, could take various forms. For instance, in (Couëtoux et al. (2011); Guez et al. (2013); Lee et al. (2020)), $\tilde{Q}((s,h),x)=Q((s,h),x)+c\sqrt{\frac{\log N((s,h))}{N((s,h),x)}}$; in PUCT (Auger et al. (2013)), $\tilde{Q}((s,h),x)=Q((s,h),x)+\sqrt{\frac{N((s,h))^{e(d)}}{N((s,h),x)}}$, where $e(d)$ is a schedule of coefficients related to the search depth $d$; in Sampled MuZero (a variant of MuZero that can be applied in continuous action spaces (Hubert et al. (2021))), $\tilde{Q}((s,h),x)=Q((s,h),x)+\hat{\pi}(x|(s,h))\frac{\sqrt{N((s,h))}}{1+N((s,h),x)}\left(c_{1}+\log\left(\frac{N((s,h))+c_{2}+1}{c_{2}}\right)\right)$. Here, $c$, $c_{1}$, $c_{2}$ are hyperparameters, $\hat{\pi}=\hat{\beta}\pi^{1-1/\tau}$ is a sample policy defined upon the real policy $\pi$. In particular, at each decision point $(s,h)$, Sampled MuZero would sample $M$ actions $\{a_{i}\}$ from the distribution $\pi^{1/\tau}$ and accordingly define $\hat{\beta}(a|(s,h))=\sum_{i}\mathds{1}_{a_{i}=a}/M$, where $\tau>0$ is a temperature hyperparameter. Thus, Sampled MuZero does not adopt progressive widening like ours. Following BAMCP, we adopt the first definition of $\tilde{Q}((s,h),x)$, though it could potentially be improved with other choices. In addition, as an implementation trick (Hamrick et al. (2021)), the Q estimates are usually normalized into $\bar{Q}\in[0,1]$ before being used to calculate $\tilde{Q}$ as above. The normalized estimates can be computed as $\bar{Q}((s,h),x)=\frac{Q((s,h),x)-Q_{\text{min}}}{Q_{\text{max}}-Q_{\text{min}}}$, where $Q_{\text{max}}$ and $Q_{\text{min}}$ are the maximum and minimum Q values observed in the search tree so far.
As the planning/search result, $\pi_{\text{ret}}$ can take multiple forms. In (Guez et al. (2013); Sunberg & Kochenderfer (2018); Lee et al. (2020)), $\pi_{\text{ret}}((s,h))=\arg\max_{a\in C((s,h))}Q((s,h),a)$; in (Sampled) MuZero, $\pi_{\text{ret}}(a|(s,h))=\frac{N((s,h),a)^{1/\tau}}{\sum_{x\in C((s,h))}N((s,h),x)^{1/\tau}}$; in ROSMO (a variant of MuZero with improved performance in offline scenarios (Liu et al. (2023))), $\pi_{\text{ret}}(a|(s,h))\propto\pi(a|(s,h))\exp(Q((s,h),a)-V((s,h)))$. Here, $\tau\in(0,1]$ is a temperature parameter and decays with the training process, ensuring the action selection becomes greedier. We select the second form for $\pi_{\text{ret}}$ in Algorithm 1. This is because (1) as described in PUCT, the returned action should be the most visited one, which is not necessarily the one with the highest Q value, and (2) ROSMO adopts one-step look-ahead rather than deep tree search at each root node, which does not align with our approach.
As for the conditions of double progressive widening, PUCT designs $\alpha$ and $\beta$ to be functions of the search depth $d$, while UCT-DPW (Couëtoux et al. (2011)) utilizes a different set of conditions: $\left\lceil{K_{a}N((s,h))^{\alpha}}\right\rceil\geq|C((s,h))|$, $\left\lceil{K_{s}N((s,h),a)^{\beta}}\right\rceil\geq|C((s,h),a)|$, where $K_{a},K_{s},\alpha,\beta$ are all constant hyperparameters. When the progressive widening condition for sampling the next state is not satisfied, either the least visited node in $C((s,h),a)$ can be selected (following PUCT), or a random node can be sampled from $C((s,h),a)$ following a distribution proportional to the number of visits (following UCT-DPW). As shown in Algorithm 1, we follow the designs of PUCT, but keep $\alpha$ and $\beta$ as constants for simplicity in hyperparameter fine-tuning.
Finally, the condition for continuing the simulation procedure, i.e., $N((s,h))>1$, could potentially be replaced with $N((s,h),a)>1$ or $N((s^{\prime},hars^{\prime}))>0$. These conditions indicate that the nodes $(s,h)$, $(s,h,a)$, and $(s^{\prime},hars^{\prime})$ have been visited before, respectively. At the end of the simulation procedure, we can either apply rollouts, i.e., simulating a single path until the end of an episode, to estimate the expected value for a leaf node $(s,h)$, or directly use $V((s,h))$ as the estimation. The former approach is widely used in online planning algorithms (Guez et al. (2013); Sunberg & Kochenderfer (2018); Lee et al. (2020)), while the latter is used in iterative frameworks like MuZero.
Appendix C Key Hyperparameter Setup
Data Type Environment BA-MBRL BA-MCTS BA-MCTS-SL
$K$ $\lambda$ $H$ $N$ $K$ $\lambda$ $H$ $N$ $K$ $\lambda$ $H$ $N$
random HalfCheetah 10 7 6 200 10 7 6 800 10 7 6 500
random Hopper 6 50 47 700 6 50 47 800 6 50 47 500
random Walker2d 10 0.5 20 700 10 0.5 20 800 10 0.5 20 500
medium HalfCheetah 12 6 6 300 12 6 5 800 12 6 5 500
medium Hopper 12 40 42 200 12 40 42 800 12 40 42 200
medium Walker2d 8 5 20 700 8 5 20 800 8 5 20 500
med-replay HalfCheetah 11 40 10 300 11 40 10 800 11 40 10 500
med-replay Hopper 7 5 5 700 7 5 5 800 7 5 5 500
med-replay Walker2d 13 2.5 47 1000 13 2.5 47 800 13 2.5 47 500
med-expert HalfCheetah 7 100 5 1000 7 100 5 800 7 100 5 1100
med-expert Hopper 12 40 43 600 12 40 43 800 12 40 43 500
med-expert Walker2d 6 20 37 400 6 20 37 800 6 20 37 500
Table 3: Key hyperparameters of the proposed algorithms for each evaluation task. $K$: ensemble size, $\lambda$: reward penalty coefficient, $H$: rollout horizon, $N$: number of training epochs.
In Table 3, we list the key hyperparameters of the proposed algorithms. For each task, an ensemble of $K$ dynamics and reward models is trained using the provided offline dataset. These learned models are then utilized as a simulator to train a control policy using off-the-shelf RL methods, such as SAC. The policy is trained for $N$ epochs. At each epoch, $50000H$ transitions are sampled by interacting with the simulator, followed by $1000$ RL training iterations. In particular, $50000$ states are randomly sampled from the offline dataset, with each state followed by a rollout lasting $H$ time steps. To mitigate overestimation, a reward penalty based on the discrepancy among the ensemble members is applied with a coefficient $\lambda$, as shown in Equation (5). The setups for $K$, $\lambda$, and $H$ are almost the same across the three algorithms and primarily inherited from the baseline – “Optimized" (Lu et al. (2022)), to make sure the improvements are brought by the Bayesian RL and deep search components.
The policy is evaluated on the ground truth environment for 10 episodes at the end of each training epoch. We report the average scores across the final 10 training epochs of our algorithms in Tables 1 and 5. It is important to note that increasing the number of training epochs $N$ does not necessarily lead to better policy performance, since the training is based on learned dynamics and reward models rather than the ground truth. According to (Lu et al. (2022)) and our experiments, the hyperparameters listed above can significantly influence the performance of model-based RL. Adjusting these hyperparameters could either enhance or impair the learning performance of our algorithms. We also suspect that the performance of the baselines listed in Tables 1 and 5, which are from their original papers, could be further improved by fine-tuning the relevant hyperparameters.
Data Type Environment BA-MCTS BA-MCTS-SL
$\rho$ $\alpha$ $c$ $\eta$ $n_{s}$ $n_{a}$ $\rho$ $\alpha$ $c$ $\eta$ $n_{s}$ $n_{a}$ $N_{SL}$ $N_{P}$
random HalfCheetah 0.1 0.5 2.5 0.3 1 20 0.1 0.5 2.5 0.3 1 20 5 0
random Hopper 0.1 0.5 2.5 0.3 1 20 0.1 0.5 2.5 0.3 1 20 5 0
random Walker2d 0.1 0.5 2.5 0.3 1 20 0.1 0.5 2.5 0.3 1 20 5 100
medium HalfCheetah 0.1 0.5 1.0 0.1 5 10 0.1 0.5 1.0 0.1 5 10 20 0
medium Hopper 0.1 0.5 2.5 0.3 1 20 0.1 0.5 1.0 0.3 1 20 5 0
medium Walker2d 0.1 0.5 2.5 0.3 1 20 0.1 0.5 2.5 0.3 1 20 5 100
med-replay HalfCheetah 0.1 0.8 1.0 0.3 5 10 0.1 0.8 2.5 0.3 1 20 5 0
med-replay Hopper 0.1 0.8 1.0 0.1 1 20 0.1 0.8 1.0 0.1 1 20 15 0
med-replay Walker2d 0.1 0.8 2.5 0.3 1 20 0.1 0.8 2.5 0.3 1 20 5 200
med-expert HalfCheetah 0.1 0.8 1.0 0.3 5 10 0.1 0.8 2.5 0.3 1 20 5 0
med-expert Hopper 0.1 0.8 1.0 0.3 1 20 0.1 0.8 1.0 0.3 1 20 5 0
med-expert Walker2d 0.1 0.8 2.5 0.3 1 20 0.1 0.8 2.5 0.3 1 20 15 100
Table 4: Important hyperparameters used in the search process.
(a) hc-med-expert
(b) hc-med-replay
(c) hc-medium
(d) hc-random
(e) hp-med-expert
(f) hp-med-replay
(g) hp-medium
(h) hp-random
(i) wk-med-expert
(j) wk-med-replay
(k) wk-medium
(l) wk-random
Figure 3: Performance of our proposed algorithms on D4RL MuJoCo tasks. The results for HalfCheetah, Hopper, and Walker2d are presented in the three rows, respectively. Each subfigure depicts the change in the undiscounted episodic return as a function of training epochs. Experiments are repeated three times with different random seeds, with the solid line representing the mean and the shaded area indicating the 95% confidence interval. For reference, the expert-level episodic returns for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively. Note that the training epochs for each algorithm, as listed in Table 3, have been linearly scaled to 800 for better visualization.
BA-MCTS and BA-MCTS-SL utilize Bayes Adaptive Monte Carlo Tree Search to collect samples for offline model-based RL. Instead of performing a tree search at every state, we randomly select a proportion (i.e., $\rho$) of states from the available 50000 states at each rollout time step as root nodes for tree search. For the remaining states, actions are sampled directly from the policy, i.e., $a\sim\pi(\cdot|s)$. The tree search procedure is detailed in Algorithm 1, with the number of MCTS iterations, $E$, set to 50. Increasing $\rho$ and $E$ can potentially enhance performance, but it will also linearly increase the computational cost. Table 4 outlines the key hyperparameters related to the search process for each algorithm and task. (1) As described in Algorithm 1, the parameters $\alpha$ and $\beta$ control the rate of double progressive widening. ($\beta$ is set as 0.5 across all tasks.) To encourage deeper search, we limit the number of actions sampled from a state under $n_{a}$ and the number of next states sampled from an action under $n_{s}$, respectively. Action selection follows the UCT rule, as discussed in Appendix A, where $c>0$ balances the exploration and exploitation. Additionally, inspired by the success of MuZero in enhancing exploration, we introduce Dirichlet noise $x_{d}$ at the root nodes, where actions are sampled from a mixture of distributions: $a\sim\eta x_{d}+(1-\eta)\pi(\cdot|s)$ and $\eta$ controls the mixture rate. Notably, for $(c,\eta,n_{a},n_{s})$, we explore the set of possible combinations: $\{(2.5,0.3,20,1),(1.0,0.3,20,1),(1.0,0.1,20,1),(1.0,0.3,10,5),(1.0,0.1,10,5)\}$ during hyperparameter fine-tuning. We believe there are likely more optimal search settings yet to be discovered. (2) In BA-MCTS-SL, policy improvement is achieved through supervised learning. We find that, rather than learning solely from samples collected within the current epoch, incorporating a buffer of samples from the past $N_{SL}$ epochs helps to stabilize the learning process. Also, in the Walker2d environment, BA-MCTS-SL requires a warm-up training phase of $N_{P}$ epochs using BA-MBRL, allowing the initial policy to generate effective signals for supervised learning.
Appendix D Comparisons with Model-free Methods on D4RL MuJoCo
Data Type Environment BA-MCTS -SL (ours) BA-MCTS (ours) BA-MBRL (ours) CQL BEAR BRAC-v SAC BC
random HalfCheetah 29.20 $\pm$ 2.00 36.23 $\pm$ 1.04 32.76 $\pm$ 1.16 35.4 25.1 31.2 30.5 2.1
random Hopper 33.83 $\pm$ 0.10 31.56 $\pm$ 0.12 31.47 $\pm$ 0.03 10.8 11.4 12.2 11.3 1.6
random Walker2d 21.89 $\pm$ 0.07 21.59 $\pm$ 0.32 21.45 $\pm$ 0.53 7.0 7.3 1.9 4.1 9.8
medium HalfCheetah 70.47 $\pm$ 3.52 75.84 $\pm$ 3.81 56.54 $\pm$ 5.20 44.4 41.7 46.3 -4.3 36.1
medium Hopper 97.75 $\pm$ 7.09 96.70 $\pm$ 14.0 98.25 $\pm$ 3.42 86.6 52.1 31.1 0.8 29.0
medium Walker2d 82.24 $\pm$ 1.85 74.73 $\pm$ 3.25 75.41 $\pm$ 4.17 74.5 59.1 81.1 0.9 6.6
med-replay HalfCheetah 61.16 $\pm$ 1.60 65.45 $\pm$ 0.81 62.50 $\pm$ 0.18 46.2 38.6 47.7 -2.4 38.4
med-replay Hopper 106.3 $\pm$ 0.13 101.8 $\pm$ 3.46 93.91 $\pm$ 4.25 48.6 33.7 0.6 3.5 11.8
med-replay Walker2d 92.13 $\pm$ 5.13 95.06 $\pm$ 2.11 97.54$\pm$ 1.93 32.6 19.2 0.9 1.9 11.3
med-expert HalfCheetah 80.53 $\pm$ 6.63 76.16 $\pm$ 10.3 90.52 $\pm$ 4.13 62.4 53.4 41.9 1.8 35.8
med-expert Hopper 112.2 $\pm$ 0.29 108.3 $\pm$ 0.22 107.8 $\pm$ 0.37 111 96.3 0.8 1.6 111.9
med-expert Walker2d 107.7 $\pm$ 0.82 110.0 $\pm$ 1.74 84.71 $\pm$ 0.87 98.7 40.1 81.6 -0.1 6.4
Average Score 74.62 74.45 71.06 54.85 39.83 31.44 4.13 25.07
Table 5: Comparisons between the proposed algorithms and model-free offline policy learning methods on the D4RL benchmark suite. Each value represents the normalized score, as proposed in (Fu et al. (2020)), of the policy trained by the corresponding algorithm. These scores are undiscounted returns normalized to approximately range between 0 and 100, where a score of 0 corresponds to a random policy and a score of 100 corresponds to an expert-level policy. For our algorithms, we report the average score of the final ten policy learning epochs and its standard deviation across three different random seeds.
As a complement to Table 1, we compare our algorithms with a series of model-free offline policy learning (Chen et al. (2024)) methods. We include SOTA model-free offline RL methods: CQL (Kumar et al. (2020)), BEAR (Kumar et al. (2019)), and BRAC-v (Wu et al. (2019)). Additionally, we show the performance of directly applying SAC or Behavioral Cloning (BC, Chen et al. (2024)) to the provided offline dataset in the last two columns. The mean performance of the baselines are taken from related works (Yu et al. (2020); Kidambi et al. (2020); Fu et al. (2020)). Our algorithms show significantly better performance, demonstrating the necessity of model-based learning in these environments. The training plots of our proposed algorithms in each environment is further detailed in Figure 3.
Appendix E Computation Cost of Sampled EfficientZero on D4RL MuJoCo
Here, we report the training time of Sampled EfficientZero on the D4RL MuJoCo tasks in Table 6. This set of experiments is run on a server with 40 Intel(R) Xeon(R) Gold 5215 CPUs and 4 Tesla V100-SXM2-32GB GPUs. Despite the substantial computational input, no corresponding performance improvement is observed. We welcome researchers to use our released code for further hyperparameter fine-tuning and to improve the results we have provided.
hc-med-expert hc-med-replay hc-medium hc-random
55.4 $\pm$ 1.3 54.8 $\pm$ 1.8 54.7 $\pm$ 1.9 54.8 $\pm$ 2.6
hp-med-expert hp-med-replay hp-medium hp-random
66.1 $\pm$ 0.9 126.6 $\pm$ 12.0 70.5 $\pm$ 1.0 153.7 $\pm$ 19.4
wk-med-expert wk-med-replay wk-medium wk-random
60.8 $\pm$ 1.8 134.9 $\pm$ 4.2 63.7 $\pm$ 1.5 175.7 $\pm$ 28.2
Table 6: Training time (in hours) of Sampled EfficientZero for each evaluation task. The mean and standard deviation from repeated experiments are reported.
Appendix F Ablation Study on the Reward Penalty
Data Type Environment BA-MCTS -SL BA-MCTS BA-MBRL BA-MCTS -SL ($\lambda=0$) BA-MCTS ($\lambda=0$) BA-MBRL ($\lambda=0$)
random HalfCheetah 29.20 $\pm$ 2.00 36.23 $\pm$ 1.04 32.76 $\pm$ 1.16 34.80 $\pm$ 1.39 38.78 $\pm$ 1.65 39.64 $\pm$ 2.86
random Hopper 33.83 $\pm$ 0.10 31.56 $\pm$ 0.12 31.47 $\pm$ 0.03 9.16 $\pm$ 0.16 7.44 $\pm$ 0.14 6.97 $\pm$ 0.07
random Walker2d 21.89 $\pm$ 0.07 21.59 $\pm$ 0.32 21.45 $\pm$ 0.53 17.53 $\pm$ 6.16 21.53 $\pm$ 0.42 21.41 $\pm$ 0.64
medium HalfCheetah 70.47 $\pm$ 3.52 75.84 $\pm$ 3.81 56.54 $\pm$ 5.20 61.64 $\pm$ 4.58 60.84 $\pm$ 2.00 41.49 $\pm$ 2.29
medium Hopper 97.75 $\pm$ 7.09 96.70 $\pm$ 14.0 98.25 $\pm$ 3.42 102.8 $\pm$ 2.29 104.4 $\pm$ 1.88 93.68 $\pm$ 11.4
medium Walker2d 82.24 $\pm$ 1.85 74.73 $\pm$ 3.25 75.41 $\pm$ 4.17 82.61 $\pm$ 0.86 57.01 $\pm$ 7.24 57.97 $\pm$ 15.4
med-replay HalfCheetah 61.16 $\pm$ 1.60 65.45 $\pm$ 0.81 62.50 $\pm$ 0.18 42.10 $\pm$ 2.85 36.65 $\pm$ 2.39 44.03 $\pm$ 7.35
med-replay Hopper 106.3 $\pm$ 0.13 101.8 $\pm$ 3.46 93.91 $\pm$ 4.25 107.9 $\pm$ 0.07 84.11 $\pm$ 2.97 91.81 $\pm$ 11.5
med-replay Walker2d 92.13 $\pm$ 5.13 95.06 $\pm$ 2.11 97.54$\pm$ 1.93 88.61 $\pm$ 5.21 97.33 $\pm$ 3.51 98.19 $\pm$ 1.23
med-expert HalfCheetah 80.53 $\pm$ 6.63 76.16 $\pm$ 10.3 90.52 $\pm$ 4.13 51.76 $\pm$ 5.31 26.60 $\pm$ 1.46 29.88 $\pm$ 2.28
med-expert Hopper 112.2 $\pm$ 0.29 108.3 $\pm$ 0.22 107.8 $\pm$ 0.37 106.8 $\pm$ 6.34 81.76 $\pm$ 6.45 86.79 $\pm$ 18.7
med-expert Walker2d 107.7 $\pm$ 0.82 110.0 $\pm$ 1.74 84.71 $\pm$ 0.87 110.8 $\pm$ 1.72 110.2 $\pm$ 0.91 53.35 $\pm$ 38.0
Average Score 74.62 74.45 71.06 68.04 60.55 55.43
Table 7: Comparison of the proposed algorithms with their corresponding versions without the reward penalty (i.e., $\lambda=0$). The definitions of the values in this table are consistent with those in Table 5.
To demonstrate the necessity of incorporating the reward penalty in offline MBRL, we conduct an ablation study by setting $\lambda$ in Eq. (5) to 0, resulting in ablated versions of our proposed three algorithms. The results are presented in Table 7. First, the average performance of the algorithms with the reward penalty is consistently better, demonstrating the importance of using reward penalties in offline MBRL to prevent the overexploitation of the learned world models (which can be inaccurate). Second, the supervised-learning-based algorithm (i.e., BA-MCTS-SL ($\lambda=0$)) is less affected by the absence of the reward penalty, compared to the policy-gradient-based methods. Notably, BA-MCTS-SL ($\lambda=0$) and BA-MCTS-SL achieve comparable performance in the Hopper and Walker2d tasks. This shows an additional advantage of BA-MCTS-SL – its reduced sensitivity to model inaccuracies. Lastly, there are instances where superior performance is achieved with $\lambda$ set to 0. For example, BA-MCTS-SL ($\lambda=0$) performs better than BA-MCTS-SL in 5 out of 12 tasks. This suggests that the performance of our algorithms in Tables 1 and 5 could be further improved by adjusting hyperparameters such as $\lambda$ [^12] As mentioned in Section 5, we retain most of the hyperparameter settings from Optimized (Lu et al. (2022)) to ensure that the performance improvements are attributed to our algorithm design..
Appendix G Details of the Tokamak Control Tasks
STATE SPACE
Scalar States $\beta_{n}$, Internal Inductance, Line Averaged Density, Loop Voltage, Stored Energy
Profile States Electron Density, Electron Temperature, Pressure, Safety Factor, Ion Temperature, Ion Rotation
ACTION SPACE
Targets Current Target, Density Target
Shape Variables Elongation, Top Triangularity, Bottom Triangularity, Minor Radius, Radius and Vertical Locations of the Plasma Center
Direct Actuators Power Injected, Torque Injected, Total Deuterium Gas Injection, Total ECH Power, Magnitude and Sign of the Toroidal Magnetic Field
Table 8: The state and action spaces of the tokamak control tasks.
Nuclear fusion is a promising energy source to meet the world’s growing demand. It involves fusing the nuclei of two light atoms, such as hydrogen, to form a heavier nucleus, typically helium, releasing energy in the process. The primary challenge of fusion is confining a plasma, i.e., an ionized gas of hydrogen isotopes, while heating it and increasing its pressure to initiate and sustain fusion reactions. The tokamak is one of the most promising confinement devices. It uses magnetic fields acting on hydrogen atoms that have been ionized (given a charge) so that the magnetic fields can exert a force on the moving particles (Pironti & Walker (2005)).
Char et al. (2024) trained a deep recurrent network as a dynamics model for the DIII-D tokamak, a device located in San Diego, California, and operated by General Atomics, using a large dataset of operational data from that device. A typical shot (i.e., episode) on DIII-D lasts around 6-8 seconds, consisting of a one-second ramp-up phase, a multi-second flat-top phase, and a one-second ramp-down phase. The DIII-D also features several real-time and post-shot diagnostics that measure the magnetic equilibrium and plasma parameters with high temporal resolution. The authors demonstrate that the learned model predicts these measurements for entire shots with remarkable accuracy. Thus, we use this model as a “ground truth" simulator for tokamak control tasks. Specifically, we generate a dataset of 725270 transitions for offline RL and evaluate the learned policy using this data-driven simulator.
The state and action spaces for the tokamak control tasks are outlined in Table 8. For detailed physical explanations of their components, please refer to (Abbate et al. (2021); Char et al. (2023); Ariola et al. (2008)). The state space consists of five scalar values and six profiles which are discretized measurements of physical quantities along the minor radius of the toroid. After applying principal component analysis (Maćkiewicz & Ratajczak (1993)), the pressure profile is reduced to two dimensions, while the other profiles are reduced to four dimensions each. In total, the state space comprises 27 dimensions. The action space includes direct control actuators for neutral beam power, torque, gas, ECH power, current, and magnetic field, as well as target values for plasma density and plasma shape, which are managed through a lower-level control module. Altogether, the action space consists of 14 dimensions. While for certain tasks, it is possible to prune the state and action spaces to reduce the learning complexity, we have chosen not to apply any domain-specific knowledge in these evaluations for general RL algorithms. We reserve the domain-specific applications of our algorithms, which would require more domain knowledge and engineering efforts, as an important future work.
We select a reference shot from DIII-D, which spans 251 time steps, and use its trajectories of Ion Rotation, Electron Temperature, and $\beta_{n}$ as targets for three tracking tasks. Specifically, $\beta_{n}$ is the normalized ratio between plasma pressure and magnetic pressure, a key quantity serving as a rough economic indicator of efficiency. Since the tracking targets vary over time, we include the time step as part of the policy input. The reward function for each task is defined as the negative squared tracking error of the corresponding component (i.e., temperature, rotation, or $\beta_{n}$) at each time step, and the reward is normalized by the episode horizon (i.e., 251 time steps). Notably, for policy learning, the reward function is provided rather than learned from the offline dataset as in D4RL tasks; and the dataset does not include the reference shot or any nearby, similar shots.

### v4

@@
 Compared to model-free methods, offline model-based reinforcement learning (MBRL) explicitly learns world models from a static dataset and uses them as surrogate simulators, improving the data efficiency and enabling the learned policy to potentially generalize beyond the dataset support.
-However, there could be various MDPs that behave identically on the offline dataset and so dealing with the uncertainty about the true MDP can be challenging.
+However, there could be various MDPs that behave identically on the offline dataset and dealing with the uncertainty about the true MDP can be challenging.
 In this paper, we propose modeling offline MBRL as a Bayes Adaptive Markov Decision Process (BAMDP), which is a principled framework for addressing model uncertainty.
-We further introduce a novel Bayes Adaptive Monte-Carlo planning algorithm capable of solving BAMDPs in continuous state and action spaces with stochastic transitions.
+We further propose a novel Bayes Adaptive Monte-Carlo planning algorithm capable of solving BAMDPs in continuous state and action spaces with stochastic transitions.
 This planning process is based on Monte Carlo Tree Search and can be integrated into offline MBRL as a policy improvement operator in policy iteration.
-Our “RL + Search" framework follows in the footsteps of superhuman AIs like AlphaZero, improving on current offline MBRL methods by incorporating more computation input.
-The proposed algorithm significantly outperforms state-of-the-art model-based and model-free offline RL methods on twelve D4RL MuJoCo benchmark tasks and three target tracking tasks in a challenging, stochastic tokamak control simulator.
+Our “RL + Search” framework follows in the footsteps of superhuman AIs like AlphaZero, improving on current offline MBRL methods by incorporating more computation input.
+The proposed algorithm significantly outperforms state-of-the-art offline RL methods on twelve D4RL MuJoCo tasks and three challenging, stochastic tokamak control tasks.
 1 Introduction
-The success of reinforcement learning (RL) typically relies on large amounts of interactions with the environment.
+The success of RL typically relies on large amounts of interactions with the environment.
 However, in real-world scenarios, such interactions can be unsafe or costly.
-As an alternative, offline RL (Levine et al.
-(2020)) leverages offline datasets of transitions, collected by a behavior policy, to train a policy that can transfer to an online task.
-To avoid overestimation of the value function for some (out-of-sample) states in the environment, which can mislead policy learning, model-free offline RL methods (Kumar et al.
-(2020); Wu et al.
-(2019)) often constrain the learned policy to remain close to the behavior policy or within the support of the offline dataset.
-However, collecting transitions that comprehensively cover possible task scenarios, or acquiring a large volume of demonstrations from a high-quality behavior policy, can be expensive.
-This challenge has led to the development of offline model-based reinforcement learning (MBRL) approaches, such as (Lu et al.
+As an alternative, offline RL (Levine et al., 2020) leverages offline datasets of transitions, collected by a behavior policy, to train a policy.
+To avoid overestimation of the expected return for out-of-distribution states, which can mislead policy learning, model-free offline RL methods (Kumar et al., 2020; Wu et al., 2019) often constrain the learned policy to remain close to the behavior policy.
+However, acquiring a large volume of demonstrations from a high-quality behavior policy, can be expensive.
+This challenge has led to the development of offline model-based reinforcement learning (MBRL) approaches, such as Lu et al.
 (2022); Guo et al.
-(2022)).
-These methods train dynamics models from offline data and optimize policies using imaginary rollouts generated by the models.
-Notably, the dynamics modeling is independent of the behavior policy, making it possible to achieve high returns even with data collected from a random policy.
-Furthermore, with careful dynamics modeling and thorough simulation, the learned policy can more effectively handle the environmental stochasticity and generalize to states beyond the support of the offline dataset.
-Given a dataset, there may be various potential MDPs that behave identically on the limited set of states and actions, but their dynamics and reward functions could differ, especially on out-of-sample states and actions.
-This implies that we are dealing with a distribution of possible world models underlying the dataset.
+(2022).
+These methods train dynamics models from offline data and optimize policies using imaginary rollouts simulated by the models.
+A key advantage of this approach is that the dynamics modeling process is decoupled from the behavior policy’s objectives.
+This makes it possible, in principle, to learn effective policies even from sub-optimal or exploratory data, so long as the dataset provides adequate coverage of the relevant state-action space.
+Multiple MDPs can exhibit identical behaviors on an offline dataset of states and actions, yet their underlying dynamics and reward functions may differ, particularly on out-of-sample regions.
+This implies that we are dealing with a distribution of possible world models underlying a dataset.
 A common strategy in offline MBRL is to learn an ensemble of world models and treat them equally.
-For instance, when determining the next state, a world model is uniformly sampled from the ensemble and generate its prediction.
-However, different ensemble members may perform better in different regions of the state-action space, making it necessary to adapt the belief over each ensemble member based on the experience accumulated since the start of the episode.
-The Bayes Adaptive Markov Decision Process (BAMDP, Duff (2002)) provides a principled framework for modelling such an adaptive process.
-We show in Section 4.1 that, despite the need for Bayesian posterior updates, BAMDPs can still be efficiently simulated using deep ensembles.
-BAMCP (Guez et al.
-(2013)) is an efficient online planning method for solving BAMDPs.
+For instance, when determining the next state, a world model would be uniformly sampled from the ensemble to generate its prediction.
+However, different ensemble members may perform better in different regions of the state-action space, making it necessary to adapt the belief over each ensemble based on the experience.
+The Bayes Adaptive Markov Decision Process (BAMDP) (Duff, 2002) provides a principled framework for modeling such adaptations, and BAMCP (Guez et al., 2013) is an efficient online planning method for solving BAMDPs.
 However, BAMCP has several limitations: (1) it relies on a ground-truth world model for planning; (2) it is restricted to discrete state and action spaces; and (3) its outcome is an action choice at a particular state, rather than a policy function.
-To address these challenges: (1) we apply a reward penalty (defined with the adapting belief) to construct a pessimistic BAMDP, preventing overexploitation of inaccurate world models (learned from the offline dataset); (2) we propose a novel planning algorithm to solve BAMDPs in continuous state and action spaces by extending BAMCP with double progressive widening (Auger et al.
-(2013)); and (3) we integrate the planning component as a policy improvement operator within policy iteration RL methods (Sutton & Barto (2018)), enabling the derivation of a policy suitable for real-time execution from the planning results.
-Specifically, the planning process is carried out through Monte Carlo Tree Search on a BAMDP.
-Integrating search with RL allows for the use of significantly more computation input, thereby improving policy learning performance.
-Grounded in the “scaling law", this paradigm has seen tremendous success in sophisticated policy learning, as demonstrated in (Silver et al.
-(2017b); Schrittwieser et al.
-(2020); AlphaProof & AlphaGeometry (2024)).
-Its application to offline MBRL, particularly in continuous control tasks, is a promising area for exploration.
-To summarize, the main contributions of this work include: (1) Introducing BAMDPs to handle model uncertainties in offline MBRL; (2) Proposing an efficient Bayes Adaptive Monte Carlo Tree Search method for planning in continuous, stochastic BAMDPs; (3) Developing the first algorithm to successfully integrate Bayesian RL, offline MBRL, and deep search for sophisticated policy learning in continuous control under highly stochastic environments; (4) Demonstrating the improvements brought by Bayesian RL and deep search across twelve D4RL MuJoCo tasks and three target tracking tasks in a stochastic tokamak control scenario (for nuclear fusion), highlighting the potential of our algorithm to tackle challenging, real-world problems.
+To address these challenges: (1) we learn an ensemble of world models from the offline dataset and construct a pessimistic BAMDP as an estimation of the true MDP; (2) we propose a novel planning algorithm to solve BAMDPs in continuous state and action spaces based on double progressive widening (Auger et al., 2013); and (3) we distill the planning outcomes into a policy function via RL, enabling real-time execution.
+Notably, we provide a theoretical proof establishing the consistency of our planner in continuous, Bayes-adaptive MDP settings.
+Our main contributions include: (1) Firstly using BAMDPs to handle model uncertainties in offline MBRL; (2) Proposing theoretically-grounded Continuous BAMCP for planning in continuous, stochastic BAMDPs; (3) Developing the first algorithm to successfully integrate Bayesian RL, offline MBRL, and deep search for data-driven continuous control; (4) Demonstrating the state-of-the-art performance on twelve D4RL MuJoCo tasks and three target tracking tasks in tokamak control (for nuclear fusion).
 2 Background
-A Markov Decision Process (MDP, Puterman (2014)) is described as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$.
+An MDP (Puterman, 2014) can be described as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$.
 $\mathcal{S}$ and $\mathcal{A}$ are the state space and action space, respectively.
-$\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ is the dynamics function and $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{[0,1]}$ is the reward function, where $\Delta_{\mathcal{X}}$ denotes the set of possible probability distributions on $\mathcal{X}$.
+$\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ is the dynamics function and $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{[0,1]}$ is the reward function, where $\Delta_{\mathcal{X}}$ denotes the set of probability distributions on $\mathcal{X}$.
 $\gamma\in[0,1)$ is a discount factor.
-The goal of solving an MDP is to find a policy $\pi:\mathcal{S}\rightarrow\Delta_{\mathcal{A}}$ that maximizes the expected return, defined as the cumulative discounted reward over time.
-A Bayes Adaptive MDP (BAMDP, Duff (2002)) can model scenarios where the precise MDP $\mathcal{M}_{\theta}=\langle\mathcal{S},\mathcal{A},\mathcal{P}_{\theta},\mathcal{R}_{\theta},\gamma\rangle$ is uncertain but is known to follow a prior distribution $b_{0}(\theta)$.
-During planning, a Bayes-optimal agent would update its belief over the MDP based on experience.
+A Bayes Adaptive MDP (BAMDP) (Duff, 2002) can model scenarios where the precise MDP $\mathcal{M}_{\theta}=\langle\mathcal{S},\mathcal{A},\mathcal{P}_{\theta},\mathcal{R}_{\theta},\gamma\rangle$ is uncertain but is known to follow a prior distribution $b_{0}(\theta)$.
+During planning, a Bayes-optimal agent would update its belief over the MDP based on its experience.
 Formally, a BAMDP can be described as a tuple $\mathcal{M}^{+}=\langle\mathcal{S}^{+},\mathcal{A},\mathcal{P}^{+},\mathcal{R}^{+},\gamma\rangle$.
@@
 After each transition $(s,a,r,s^{\prime})$, the belief is updated to the corresponding Bayesian posterior: $b^{\prime}(\theta)\propto b(\theta)P((s,a,r,s^{\prime})|\theta)=b(\theta)\mathcal{P}_{\theta}(s^{\prime}|s,a)\mathcal{R}_{\theta}(r|s,a)$.
-Accordingly, $\mathcal{P}^{+}$ and $\mathcal{R}^{+}$ can be defined as follows:
-[…]
-The Q-function that satisfies the Bellman optimality equations: $(\forall x=(s,b)\in\mathcal{S}^{+},a\in\mathcal{A})$
-[…]
-is the Bayes-optimal Q-function and $\pi^{*}(s,b)=\arg\max_{a}Q^{*}((s,b),a)$ is the Bayes-optimal policy.
-Actions derived from $\pi^{*}$ are executed in the real MDP and constitute the best course of action for a Bayesian agent with respect to its prior belief $b_{0}$ over the environment (Guez et al.
-(2014)).
-A BAMDP can be cast into a partially observable MDP (POMDP) [^1] Such a POMDP can be defined as a tuple $\langle\mathcal{S}^{+},\mathcal{A},\mathcal{P}^{+},\mathcal{R}^{+},\mathcal{Z},\mathcal{S},\gamma\rangle$, according to its original definition in (Littman (2009)).
-by viewing $\mathcal{S}^{+}$ and $\mathcal{S}$ as the state and observation spaces, respectively, and setting the observation function as $\mathcal{Z}((s^{\prime},b^{\prime}),a,o)=\mathds{1}(s^{\prime}=o)$.
-($\mathcal{Z}$ specifies the probability of observing $o$ under the previous action $a$ and current state $(s^{\prime},b^{\prime})$.) As a result, approaches developed for POMDPs can potentially be used to solve BAMDPs.
-Bayesian Reinforcement Learning (BRL, Ghavamzadeh et al.
-(2015)), as introduced above, is a principled approach to dealing with uncertainty in the world model $\mathcal{M}_{\theta}$ and has two main advantages: (1) Domain knowledge can be injected by defining a proper prior belief; (2) A Bayes Adaptive policy solves the exploration-exploitation dilemma by explicitly including the belief in its state representation and incorporating belief updates into the planning process (Sorg et al.
-(2010)).
-However, Bayes-optimal planning is generally intractable, for which we introduce some approximate methods in the next section.
+Accordingly, $\mathcal{P}^{+}$ and $\mathcal{R}^{+}$ are defined as follows:
+$\mathcal{P}^{+}((s^{\prime},b^{\prime\prime})|(s,b),a)=\mathds{1}(b^{\prime\prime}=b^{\prime})\int_{\theta}\mathcal{P}_{\theta}(s^{\prime}|s,a)b(\theta)d\theta$ (1)
+$\mathcal{R}^{+}((s,b),a)=\int_{\theta}\mathcal{R}_{\theta}(s,a)b(\theta)d\theta$
+The Q-function that satisfies the Bellman optimality equations (i.e., Eq.
+(2)) is the Bayes-optimal Q-function and $\pi^{*}(s,b)=\arg\max_{a}Q^{*}((s,b),a)$ is the Bayes-optimal policy.
+Actions derived from $\pi^{*}$ are executed in the true MDP and constitute the best course of actions for a Bayesian agent with a prior belief $b_{0}$ over the underlying MDP (Guez et al., 2014).
+Bayesian RL (Ghavamzadeh et al., 2015) seeks to compute the optimal policy, $\pi^{*}$, and provides a principled framework for handling uncertainty in the world model $\mathcal{M}_{\theta}$.
+However, the requirement for Bayesian posterior updates (i.e., Eq.
+(1)) often renders it computationally intractable, and we discuss several approximate methods in Section 3.
+$Q^{*}(x,a)=\mathcal{R}^{+}(x,a)+\gamma\int_{x^{\prime}}V^{*}(x^{\prime})\mathcal{P}^{+}(x^{\prime}|x,a)dx^{\prime}$ (2)
+$V^{*}(x^{\prime})=\max_{a}Q^{*}(x^{\prime},a),\ \forall x=(s,b)\in\mathcal{S}^{+},\ a\in\mathcal{A}$
 3 Related Works
-Offline RL (Levine et al.
-(2020); Chen et al.
-(2024)) enables an agent to learn control policies from datasets of environment transitions pre-collected by a behavior policy $\mu$, i.e., $\mathcal{D}_{\mu}=\{[(s_{t}^{i},a_{t}^{i},r_{t}^{i})_{t=1}^{T}]_{i=1}^{N}\}$, circumventing the need for potentially expensive or unsafe online interactions.
-Model-free offline RL methods directly learn value/policy functions from $\mathcal{D}_{\mu}$ but restrict the policy to remain close to $\mu$ or to behave within the support of $\mathcal{D}_{\mu}$.
-Offline Model-based RL (MBRL) methods, on the other hand, explicitly learn world models $\mathcal{M}_{\theta}$ from $\mathcal{D}_{\mu}$ (through supervised learning) and adopt $\mathcal{M}_{\theta}$ as a surrogate simulator, enabling the learned policy to possibly generalize to states beyond $\mathcal{D}_{\mu}$.
-Specifically, both planning methods (Argenson & Dulac-Arnold (2021); Zhan et al.
-(2022); Diehl et al.
-(2023)), such as MPC (Garcia et al.
-(1989)), and RL methods (Yu et al.
+Offline model-based RL: Offline RL (Chen et al., 2024) enables an agent to learn control policies from datasets of environment transitions pre-collected by a behavior policy $\mu$, i.e., $\mathcal{D}_{\mu}=\{[(s_{t}^{i},a_{t}^{i},r_{t}^{i})_{t=1}^{T}]_{i=1}^{N}\}$.
+Offline Model-based RL (MBRL) methods explicitly learn world models $\mathcal{M}_{\theta}$ from $\mathcal{D}_{\mu}$ and adopt $\mathcal{M}_{\theta}$ as a surrogate simulator, enabling the learned policy to possibly generalize to states beyond $\mathcal{D}_{\mu}$.
+Both planning methods (Argenson and Dulac-Arnold, 2021; Zhan et al., 2022; Diehl et al., 2023) and RL methods (Yu et al., 2020; Kidambi et al., 2020; Lu et al., 2022) can use the learned $\mathcal{M}_{\theta}$ to obtain a policy.
+However, since $\mathcal{D}_{\mu}$ may not span the entire state-action space, $\mathcal{M}_{\theta}$ is unlikely to be globally accurate.
+Learning/Planning without any safeguards against such model inaccuracy can yield poor results.
+Yu et al.
 (2020); Kidambi et al.
 (2020); Lu et al.
-(2022); Yu et al.
-(2021); Guo et al.
-(2022)) can be applied on top of the learned $\mathcal{M}_{\theta}$.
-However, since $\mathcal{D}_{\mu}$ may not span the entire state-action space, $\mathcal{M}_{\theta}$ is unlikely to be globally accurate.
-Learning/Planning without any safeguards against such model inaccuracy can yield poor results.
-In this case, the authors of (Yu et al.
-(2020); Kidambi et al.
-(2020); Lu et al.
-(2022)) propose learning an ensemble of world models, using ensemble-based uncertainty estimations to construct a pessimistic MDP (P-MDP), and learning a near-optimal policy atop it.
+(2022) propose learning an ensemble of world models, using ensemble-based uncertainty estimations to construct a pessimistic MDP (P-MDP), and learning a near-optimal policy atop it.
 Ideally, for any policy, the performance in the real environment is lower-bounded by the performance in the corresponding P-MDP (with high probability), thus avoiding being overly optimistic about an inaccurate model.
-For practical implementation of the P-MDP, Lu et al.
-(2022) conduct a thorough empirical study to compare a series of heuristics.
-However, none of these offline MBRL methods have modeled the problem as a BAMDP, even though Bayesian RL/planning provides a principled framework for handling model uncertainty.
-Additionally, with the learned model $\mathcal{M}_{\theta}$, deep and broad search in the policy learning process is possible, but the offline RL methods introduced above have not attempted such search-based policy improvement.
-Monte-Carlo Tree Search (MCTS, Browne et al.
-(2012)) has been successfully integrated with RL, as exemplified by AlphaZero (Silver et al.
-(2017a)) and MuZero (Schrittwieser et al.
-(2020)).
-These methods have achieved superhuman performance in domains requiring highly precise and sophisticated decision-making processes.
-AlphaZero relies on given world models, whereas MuZero learns the world model and policy simultaneously by interacting with the environment.
-Although there have been various extensions of MuZero (Hubert et al.
-(2021); Schrittwieser et al.
-(2021); Ye et al.
-(2021); Danihelka et al.
-(2022); Antonoglou et al.
-(2022); Oren et al.
-(2022); Chen et al.
-(2023); Zhao et al.
-(2024)), most algorithms are designed for online MBRL.
-According to Niu et al.
-(2023), the applications of MuZero in offline learning, especially for continuous control in highly stochastic environments, which is our focus, still require significant improvement.
-Although search-based, MuZero learns a dynamics model and a reward model in a latent state space rather than the real state space, does not apply uncertainty estimation of the learned models as in usual offline MBRL methods, and does not adopt the Bayesian RL framework, making it fundamentally different from our algorithm design.
-As mentioned in Section 2, Bayes-optimal planning is typically intractable.
-Approximate methods, such as (Asmuth et al.
+Notably, none of these offline MBRL methods have modeled the problem as a BAMDP, although Bayesian RL provides a principled framework for handling model uncertainty.
+Bayes-adaptive planning: Approximate methods for solving BAMDPs, such as Asmuth et al.
 (2009); Sorg et al.
-(2010); Castro & Precup (2010); Asmuth & Littman (2011); Wang et al.
-(2012); Fonteneau et al.
+(2010); Castro and Precup (2010); Asmuth and Littman (2011); Wang et al.
 (2012); Guez et al.
 (2013); Slade et al.
-(2020)) have been developed.
-As a representative work, BAMCP (Guez et al.
-(2013)) adopts MCTS for Bayes Adaptive planning and is shown to converge in probability to a near Bayes-optimal policy at the root node of the search tree.
-However, all these methods cannot be directly applied to large-scale MDPs with continuous state and action spaces.
-Guez et al.
-(2014) and Hoang & Vien (2020) attempt to extend the use of BAMCP to scenarios with continuous states and actions using value/policy function approximations, but they replace MCTS with simple Monte Carlo rollouts (Bertsekas & Castañón (1999)) and do not utilize UCT (Kocsis & Szepesvári (2006)).
-Tree search makes branches at each node in the tree, providing a more thorough look-ahead search than simple rollouts, and UCT, as an efficient exploration algorithm, is the basis for the convergence guarantee of BAMCP.
-Moreover, these planning algorithms are not designed for offline MBRL.
-How to incorporate search-based planning in the policy improvement stage of RL and how to handle the uncertainty of the learned world model require further exploration, which are shown in the following section.
+(2020) have been developed.
+As a representative work, BAMCP (Guez et al., 2013) adopts Monte-Carlo Tree Search (MCTS) (Browne et al., 2012) for Bayes-adaptive planning and is shown to converge in probability to a near Bayes-optimal policy at the root node of the search tree.
+However, all these methods cannot be applied to large-scale MDPs with continuous state and action spaces.
+Also, these planning algorithms are not designed for offline MBRL.
+How to incorporate planning outcomes for policy improvement in RL and how to handle the inaccuracy of the learned world model during planning still require exploration.
 4 Methodology
-In this work, we propose a novel offline MBRL algorithm based on Bayes Adaptive MCTS.
-The core challenge is to design a Bayes Adaptive planning method that is efficient in large stochastic MDPs.
-In this case, we propose Continuous BAMCP in Section 4.2, which can be applied to continuous control tasks with high dynamics stochasticity.
-Then, in Section 4.3, we present a search-based policy iteration framework, where the search results are distilled into policy and value networks for policy improvement and policy evaluation, respectively, at each iteration.
+In this section, we present a novel offline MBRL algorithm based on Bayes Adaptive MCTS.
+In Section 4.2, we propose a Bayes Adaptive planning method (i.e., Continuous BAMCP) that can be applied to stochastic continuous control tasks.
+Then, in Section 4.3, we present a search-based policy iteration framework, where the search results (from Continuous BAMCP) are distilled into policy and value networks for policy improvement/evaluation at each iteration.
 In this way, we integrate offline MBRL with Bayes Adaptive MCTS.
-Both components require the use of an ensemble of world models for practical implementation and uncertainty quantification, which is detailed in Section 4.1.
+Both components require the use of an ensemble of world models for either practical implementation or uncertainty quantification, as detailed in Section 4.1.
 4.1 The Key Role of Deep Ensembles
-Offline MBRL methods estimate world models $\mathcal{M}_{\theta}$ from a static dataset $\mathcal{D}_{\mu}$, which would inevitably induce epistemic uncertainty about the identity of the real MDP $\mathcal{M}^{*}$.
-Specifically, there could be various potential MDPs that behave identically on the limited set of states and actions in $\mathcal{D}_{\mu}$, but their dynamics and reward functions may differ, especially on out-of-sample states and actions.
+Offline MBRL methods estimate world models $\mathcal{M}_{\theta}$ from a static dataset $\mathcal{D}_{\mu}$.
+However, there could be various MDPs that behave identically on the limited set of states and actions in $\mathcal{D}_{\mu}$, but their dynamics and reward functions may differ, especially on out-of-sample states and actions.
 Thus, we are actually dealing with a distribution of world models that follow a prior distribution $b_{0}(\theta)\triangleq P(\mathcal{M}_{\theta}|\mathcal{D}_{\mu})$.
-As introduced in Section 2, Bayesian RL based on BAMDP is a principled framework for handling model uncertainty by explicitly including the belief over the models in its state representation.
-Essentially, the belief is updated with experience, providing a measure of how the models’ uncertainty has changed since the beginning of the episode.
-As a result, the agent can adjust its behavior upon receiving new information that reduces the epistemic uncertainty.
-Such an adaptive policy is necessary to act optimally in offline RL, as demonstrated in (Ghosh et al.
-(2022)).
-The idea of deep ensembles (Lakshminarayanan et al.
-(2017)) is to train multiple deep neural networks as approximations of a function, each using a different weight initialization and optimized with a different mini-batch sequence.
-For offline MBRL, we can learn an ensemble of dynamics models $\{\mathcal{P}_{\theta}^{1},\cdots,\mathcal{P}_{\theta}^{K}\}$ and reward models $\{\mathcal{R}_{\theta}^{1},\cdots,\mathcal{R}_{\theta}^{K}\}$ [^2] In some MBRL scenarios, a certain reward function is available, for instance, as defined by domain experts.
-Otherwise, the reward and dynamics function $(\mathcal{R}_{\theta}^{i},\mathcal{P}_{\theta}^{i})$ are usually trained as a unified probabilistic model $\mathcal{N}(\mu_{\theta}^{i},\sigma_{\theta}^{i})$, since the reward $r$ can be viewed as an element of the next state $s^{\prime}$.
-from the dataset $\mathcal{D}_{\mu}$ by minimizing the following supervised learning loss: ($i=1,\cdots,K$)
-$\mathcal{L}(\mathcal{P}_{\theta}^{i})=-\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{\mu}}\left[\log\mathcal{P}_{\theta}^{i}(s^{\prime}|s,a)\right],\ \mathcal{L}(\mathcal{R}_{\theta}^{i})=-\mathbb{E}_{(s,a,r)\sim\mathcal{D}_{\mu}}\left[\log\mathcal{R}_{\theta}^{i}(r|s,a)\right]$ (3)
+As introduced in Section 2, Bayesian RL handles such model uncertainty by explicitly including the belief over the models in its state representation.
+The belief is updated with experience, providing a measure of how the models’ uncertainty has changed since the beginning of the episode.
+The idea of deep ensembles (Lakshminarayanan et al., 2017) is to train multiple deep neural networks as approximations of a function (for uncertainty quantification), each using a different weight initialization and optimized with a different mini-batch sequence.
+For offline MBRL, we can learn an ensemble of dynamics models $\{\mathcal{P}_{\theta}^{1},\cdots,\mathcal{P}_{\theta}^{K}\}$ and reward models $\{\mathcal{R}_{\theta}^{1},\cdots,\mathcal{R}_{\theta}^{K}\}$ from the dataset $\mathcal{D}_{\mu}$ by minimizing the following supervised learning loss: ($i=1,\cdots,K$)
+$\mathcal{L}(\mathcal{P}_{\theta}^{i})=-\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{\mu}}\left[\log\mathcal{P}_{\theta}^{i}(s^{\prime}|s,a)\right]$ (3)
+$\mathcal{L}(\mathcal{R}_{\theta}^{i})=-\mathbb{E}_{(s,a,r)\sim\mathcal{D}_{\mu}}\left[\log\mathcal{R}_{\theta}^{i}(r|s,a)\right]$
 $\{(\mathcal{P}_{\theta}^{i},\mathcal{R}_{\theta}^{i})_{i=1}^{K}\}$ can be viewed as a set of independent and identically distributed (IID) samples from the prior $P(\mathcal{M}_{\theta}|\mathcal{D}_{\mu})$ and constitute a finite approximation of the space of world models.
@@
 After receiving a transition $(s,a,r,s^{\prime})$, the belief can be updated as follows:
-$b^{\prime}(\theta)(i)=x^{i}/\sum_{j=1}^{K}x^{j},\ x^{i}=b(\theta)(i)\mathcal{P}_{\theta}^{i}(s^{\prime}|s,a)\mathcal{R}_{\theta}^{i}(r|s,a)$ (4)
-This update requires a single inference from each ensemble member, but can be parallelized for computational efficiency.
-Equation (4) is a practical implementation of the Bayesian posterior update based on deep ensembles, and the simplification of the definition of $b(\theta)$ makes the transitions in Bayesian RL (i.e., Equation (1)) both practical and efficient.
-The ensemble can also be used for uncertainty quantification.
-As aforementioned, our algorithm relies on thorough search on the learned world models.
-Without any constraints on the search process, the learned policy may overfit to an inaccurate model (by overestimating the expected return) and fail in the true MDP.
+$b^{\prime}(\theta)(i)\propto b(\theta)(i)\mathcal{P}_{\theta}^{i}(s^{\prime}|s,a)\mathcal{R}_{\theta}^{i}(r|s,a),\ i=1,\cdots,K$ (4)
+This is a practical implementation of the Bayesian posterior (originally defined in Eq.
+(1)) update based on deep ensembles, where $b(\theta)$, $b^{\prime}(\theta)$, and $\mathcal{P}_{\theta}^{i}(s^{\prime}|s,a)\mathcal{R}_{\theta}^{i}(r|s,a)$ denote the prior, posterior distributions, and likelihood, respectively.
 Although the agent could adapt its belief and follow more reliable ensemble members in the Bayesian RL framework, there could be regions in the state-action space where none of the members generalize well, as they are all learned from a static offline dataset.
-A typical solution is to construct a P-MDP (see Section 3), which lower-bounds the true MDP and discourages the policy from regions where there is large discrepancy between the true and learned world models.
-We construct the P-MDP by modifying each reward estimation $r$ into $\tilde{r}$: ($\mu_{\theta}(s,a)=\sum_{i=1}^{K}b(\theta)(i)\mu_{\theta}^{i}(s,a)$)
-$\tilde{r}(s,a,r,b(\theta))=r-\lambda\sqrt{\sum_{i=1}^{K}b(\theta)(i)(\sigma_{\theta}^{i}(s,a)^{2}+\mu_{\theta}^{i}(s,a)^{2})-\mu_{\theta}(s,a)^{2}}$ (5)
-The reward penalty is weighted by a hyperparameter $\lambda>0$ and corresponds to the standard deviation (std) of the mixture of Gaussian dynamics models, where $\mu_{\theta}^{i}$ and $\sigma_{\theta}^{i}$ are the mean and std from the ensemble member $i$.
-This penalty design combines epistemic and aleatoric model uncertainty and has been shown to be successful at capturing errors in predicted dynamics (Lu et al.
-(2022)) [^3] In the original literature (Lakshminarayanan et al.
-(2017); Lu et al.
-(2022)), the ensemble is treated as a uniformly-weighted mixture model, i.e., $b(\theta)(i)=1/K,\ (i=1,\cdots K)$, since their belief is not adaptive.
-Equation (5) fits into the Bayesian RL framework by adapting the belief $b(\theta)$, which is part of our novelty..
+A typical solution for such uncertainty is to construct a pessimistic MDP, which discourages the agent from exploiting regions with high uncertainty.
+Another use of the deep ensemble is to construct a Pessimistic Bayes-Adaptive MDP (P-BAMDP).
+In particular, we apply a reward penalty to each transition $(s,a,r)$ and get a penalized reward $\tilde{r}$ as follows:
+$r-\lambda\cdot\text{std}\left[r^{i}+\gamma\mathbb{E}_{\begin{subarray}{c}s^{\prime i}\sim\mathcal{P}_{\theta}^{i},a^{\prime}\sim\pi(s^{\prime i})\end{subarray}}[Q_{\psi^{-}}(s^{\prime i},a^{\prime})]\right]_{i=1}^{K}$ (5)
+Here, $s^{\prime i}\sim\mathcal{P}_{\theta}^{i}(\cdot|s,a)$ and $r^{i}=\mathbb{E}[\mathcal{R}_{\theta}^{i}(\cdot|s,a)]$ denote the next state and expected reward from the $i$-th ensemble member.
+$\pi$ and $Q_{\psi^{-}}$ represent the policy and target Q-network.
+This penalty term is the standard deviation of the one-step lookahead Q-value targets, computed across all ensemble members.
+The underlying intuition is that a high variation in these predictions indicates a high degree of uncertainty at this region.
+As demonstrated by Sun et al.
+(2023), employing this penalty to define a pessimistic MDP yields an agent performance estimate that more closely matches real-world outcomes, compared to penalties based solely on ensemble disagreement in next-state prediction, as proposed in Yu et al.
+(2020); Kidambi et al.
+(2020).
+4.2 Continuous BAMCP
+BAMCP (Guez et al., 2013) has been successful in solving large-scale BAMDPs, as detailed in Appendix A, but it is limited to scenarios with discrete state and action spaces.
+In this subsection, we introduce a novel planning method to approximate the Bayes-optimal policy at a decision point $(s,h)$ ($h$ denotes the transition history that ends at $s$), which can be used to solve BAMDPs with continuous states/actions and stochastic transition kernels.
+Double Progressive Widening (DPW): DPW (Couëtoux et al., 2011; Auger et al., 2013) is a technique to extend the use of MCTS to continuous state and action spaces.
+Instead of exploring all possible actions and next states, DPW maintains a finite list of options to search at each decision node, incrementally adding new options to the list based on the visitation counts of that decision node.
+To be specific, for a node $(s,h)$, a new action $a$ is sampled (with the current policy $\pi$) and added to its children set $C((s,h))$, if $\left\lfloor{N((s,h))^{\alpha}}\right\rfloor\geq|C((s,h))|$, where $\alpha\in(0,1)$ is a hyperparameter that controls the growth rate and $N$ denotes the visitation counts of $(s,h)$.
+Otherwise, an action is selected from existing options in $C((s,h))$ according to the UCT (Kocsis and Szepesvári, 2006) rule.
+Similarly, to handle the infinitely many possible state transitions in stochastic environments, a new next state $s^{\prime}$ is added to the children set $C((s,h),a)$ only if $\left\lfloor{N((s,h),a)^{\beta}}\right\rfloor\geq|C((s,h),a)|$ ($\beta\in(0,1)$).
+Otherwise, the least visited child in $C((s,h),a)$ will be selected as the next state.
+With DPW, the sets of possible actions or next states to explore are finite, allowing deep tree search as in discrete scenarios.
+Integration of DPW and BAMCP: Directly combining DPW and BAMCP (i.e., Algorithm 3) cannot solve BAMDPs with continuous state and action spaces.
+As introduced in Appendix A, BAMCP relies on root sampling, which samples dynamics functions only at the root node and follows a specific dynamics function throughout a simulation rollout.
+However, the rationale of root sampling (i.e., Lemma A.1) does not hold when applying DPW [^1] The last equality of Eq.
+(6) does not hold, since $\tilde{b}(\theta|has^{\prime})\propto\tilde{b}(\theta|ha)\widetilde{\mathcal{P}}_{\theta}(s^{\prime}|s,a)\neq\tilde{b}(\theta|ha)\mathcal{P}_{\theta}(s^{\prime}|s,a)$.
+$\widetilde{\mathcal{P}}_{\theta}(s^{\prime}|s,a)$ represents the distribution of next states when applying DPW, which differs from the true distribution $\mathcal{P}_{\theta}(s^{\prime}|s,a)$, as dictated by the DPW rule..
+As an alternative design, Polynomial Upper Confidence Tree (PUCT) (Auger et al., 2013), built upon DPW, is a provably consistent planning method for solving MDPs with infinite-scale state and action spaces and highly stochastic transition dynamics.
+Thus, we propose to recast P-BAMDPs as MDPs and solve them using PUCT, as a novel Bayes planning method for continuous state and action spaces.
 Algorithm 1 Continuous BAMCP
 Input: $\pi$, $V$, $E$, $d_{\text{max}}$, $\gamma$, $\alpha$, $\beta$, $\mathcal{P}_{\theta}^{1:K}$, $\mathcal{R}_{\theta}^{1:K}$
-procedure Search($(s,h),b(\theta)$)
+[…]
   for $e=1\cdots E$ do
-         Simulate($(s,h),b(\theta),d_{\text{max}}$)
+[…]
   end for
+  $\pi_{\text{ret}}(a|(s,h))\propto N((s,h),a)),a\in C((s,h)$
   $v_{\text{ret}}=\sum_{a\in C((s,h))}\frac{N((s,h),a)}{N((s,h))}Q((s,h),a)$
@@
 end procedure
-procedure Simulate($(s,h),b(\theta),d$)
+[…]
   if $d==0$ then return $V((s,h))$
@@
   $r,s^{\prime},b^{\prime}(\theta)\leftarrow\textsc{StatePW}((s,h),b(\theta),a)$
-[…]
+  $N((s,h))\mathrel{{+}{=}}1,N((s,h),a)\mathrel{{+}{=}}1$
   if $N((s,h))>1$ then
@@
 end procedure
-procedure StatePW($(s,h),b(\theta),a$)
+[…]
   if first visit then $C((s,h),a)\leftarrow\emptyset$
@@
 end procedure
-4.2 Bayes Adaptive MCTS in Continuous State and Action Spaces
-BAMCP (Guez et al.
-(2013)) has been successful in solving large-scale BAMDPs, as detailed in Appendix A, but it is limited to scenarios with discrete state and action spaces.
-In this subsection, we introduce an online planning method to approximate the Bayes-optimal policy at a decision point $(s,h)$.
-($h$ denotes the transition history that ends at $s$.) In particular, this method can be used to solve BAMDPs with continuous states and actions.
-Specifically, we adopt double progressive widening (DPW, Couëtoux et al.
-(2011); Auger et al.
-(2013)), which maintains a finite list of chance nodes to be searched at each decision point and incrementally adds a new chance node to the list based on the visitation counts.
-For a node $(s,h)$, a new action $a$ is sampled (with the current policy) and added to its children set $C((s,h))$, if $\left\lfloor{N((s,h))^{\alpha}}\right\rfloor\geq|C((s,h))|$, where $\alpha\in(0,1)$ is a hyperparameter that controls the growth rate and $N$ denotes the visitation counts.
-Otherwise, an action will be sampled from $C((s,h))$ according to the UCT (Kocsis & Szepesvári (2006)) rule.
-Similarly, to handle the infinitely many possible transitions, a new next state $s^{\prime}$ is added to the children set $C((s,h),a)$ only if $\left\lfloor{N((s,h),a)^{\beta}}\right\rfloor\geq|C((s,h),a)|$ ($\beta\in(0,1)$).
-Otherwise, the least visited child in $C((s,h),a)$ will be selected as the next state.
-With DPW, the sets of possible actions or next states for search are all finite, allowing deep search as in discrete scenarios, and the more promising states and actions (with higher $N$) have more subsequent branches, thereby reducing corresponding estimation uncertainty.
-However, directly integrating DPW and BAMCP (i.e., Algorithm 3) likely cannot solve BAMDPs with continuous state and action spaces.
-As introduced in Appendix A, BAMCP relies on root sampling, which samples dynamics functions (from $b(\theta)$) only at the root node and follows a specific dynamics function throughout a simulation procedure.
-However, the rationale of root sampling (i.e., Lemma A.1) does not hold when applying DPW [^4] The last equality of Eq.
-(6) does not hold, since $\tilde{b}(\theta|has^{\prime})\propto\tilde{b}(\theta|ha)\widetilde{\mathcal{P}}_{\theta}(s^{\prime}|s,a)\neq\tilde{b}(\theta|ha)\mathcal{P}_{\theta}(s^{\prime}|s,a)$.
-$\widetilde{\mathcal{P}}_{\theta}(s^{\prime}|s,a)$ represents the state transition distribution when applying DPW, which differs from the true distribution $\mathcal{P}_{\theta}(s^{\prime}|s,a)$, as dictated by the DPW rule..
-Also, BAMCP assumes that the reward function is certain rather than a function of $\theta$.
-On the other hand, Polynomial Upper Confidence Trees (PUCT, Auger et al.
-(2013)), based on DPW, is a provably consistent planning method for solving MDPs with infinite-scale state/action spaces and arbitrarily stochastic transition kernels.
-In this case, we propose casting BAMDPs into MDPs (i.e., $\mathcal{M}^{+}$ defined in Section 2) and solving them with PUCT.
-The pseudo code is shown as Algorithm 1.
-Ideally, as the number of simulations $E\rightarrow\infty$, PUCT can find a near-optimal solution of $\mathcal{M}^{+}$, which is also a near Bayes-optimal solution for the true environment.
-Each simulation follows a path that begins at the root node and ends at a leaf node, utilizing progressive widening for sampling both actions and next states, as illustrated in the ActionPW and StatePW procedures.
-Compared to PUCT, the significant modifications include: (1) replacing $\langle\mathcal{S},\mathcal{P},\mathcal{R}\rangle$ with their extended definitions in BAMDPs, $\langle\mathcal{S}^{+},\mathcal{P}^{+},\mathcal{R}^{+}\rangle$, and (2) applying reward penalties to account for model uncertainty.
-To be specific, in StatePW, $r$ and $s^{\prime}$ are sampled from a mixture of ensemble members, which is a practical implementation of sampling from $\mathcal{R}^{+}$ and $\mathcal{P}^{+}$ as shown in Eq.
+Proposed algorithm – Continuous BAMCP: In Algorithm 1, each simulation follows a path from the root node to an unvisited node, utilizing progressive widening when sampling actions or next states, as detailed in the ActionPW and StatePW procedures.
+Compared to PUCT, the main modifications include: (1) replacing $\langle\mathcal{S},\mathcal{P},\mathcal{R}\rangle$ in MDPs with their extended definitions in BAMDPs, i.e., $\langle\mathcal{S}^{+},\mathcal{P}^{+},\mathcal{R}^{+}\rangle$, and (2) applying reward penalties to account for model uncertainty.
+To be specific, in StatePW, $r$ and $s^{\prime}$ are sampled from the distribution predicted by all ensemble members, which is a practical implementation of sampling from $\mathcal{R}^{+}$ and $\mathcal{P}^{+}$ as outlined in Eq.
 (1).
 After receiving the transition $(s,a,r,s^{\prime})$, the belief vector $b(\theta)$ is updated to $b^{\prime}(\theta)$ following Eq.
-(4), finishing the transition in $\mathcal{S}^{+}$ from $(s,b(\theta))$ to $(s,b^{\prime}(\theta))$.
-Moreover, the reward $r$ is adjusted with a penalty term as defined in Eq.
-(5), which is then used to calculate the return $R$.
-As aforementioned, applying such a reward penalty can effectively mitigate the issue of model exploitation.
+(4), finishing the transition in $\mathcal{S}^{+}$ from $(s,b(\theta))$ to $(s^{\prime},b^{\prime}(\theta))$.
+The transition history $h$ is then updated to $h^{\prime}=hars^{\prime}$.
+Secondly, in Simulate, a penalized reward $\tilde{r}$ (i.e., Eq.
+(5)) is used, which effectively mitigates overfitting to inaccurate ensemble members by incorporating pessimism, as discussed in Section 4.1.
+The following theorem establishes that our planner is a reliable and consistent solver for P-BAMDPs defined in Section 4.1.
+The detailed proof is provided in Appendix B.
+This consistency guarantee is crucial, as it ensures that with sufficient computation, our planner’s output converges to the true optimal value within the pessimistic model, providing a reliable value target for our overall learning framework.
+Theorem 4.1.
+Let $\hat{V}(z_{0})$ be the value of the root node $z_{0}$ estimated by Continuous BAMCP after $E$ simulations.
+Under the regularity and value function conditions defined in Appendix B.1, the error of this estimate with respect to the P-BAMDP’s optimal value $V^{*}_{p}(z_{0})$ is bounded as follows, exponentially surely in $E$:
+$\left|\hat{V}(z_{0})-V^{*}_{p}(z_{0})\right|\leq C_{0}E^{-\gamma_{0}}+\gamma^{d_{max}}\epsilon_{V},$
+where $\gamma_{0},C_{0}>0$ are constants, $\gamma$ is the discount factor, $d_{max}$ is the maximum search depth, and $\epsilon_{V}$ is the uniform bound on the value approximation error at the leaf nodes.
 […]
-4.3 The Overall Framework: Search-based Policy Iteration
-In this subsection, we present a framework (inspired by MuZero [^6] The novelty of our algorithm in comparison to MuZero is detailed in Section 3.) that integrates continuous BAMCP (i.e., Algorithm 1) into policy improvement and policy evaluation.
-By iteratively running these procedures, we can approach a near Bayes-optimal policy, i.e., $\pi$, that can be directly referred to during execution in the true environment.
-The pseudo code of the overall framework is shown as Algorithm 2.
-For efficiency, a learner and a number of actors execute in parallel, reading from and sending data to the replay buffer $\mathcal{D}$ respectively.
-The actors update their copies of policy and value functions every $E_{l}$ learner steps.
-In particular, each actor would interact with the learned world models to sample trajectories $\tau$ [^7] In Algorithm 2, $\hat{\rho}(s)$ denotes the empirical distribution of initial states in the offline dataset $\mathcal{D}_{\mu}$.
-At each time step of the trajectory, a Search procedure (defined in Algorithm 1) is executed using the actor’s current copy of $\pi$ and $V$ for planning at the current decision point.
-The search result $\pi_{\text{ret}}$ is then used to indicate the action choice, i.e., $a\sim\pi_{\text{ret}}(\cdot|(s,h))$.
-The subsequent transition process follows a BAMDP, where $r\sim\mathcal{R}^{+}(\cdot|(s,h),a)$, $s^{\prime}\sim\mathcal{P}^{+}(\cdot|(s,h),a)$, and the belief $b({\theta})$ adapts with experience, as described in StatePW.
-The collected trajectories are used in the learning process, where $\pi$ is trained to mimic the search result $\pi_{\text{ret}}$ by minimizing a cross-entropy loss (i.e., $\mathcal{L}(\pi,\{\tau_{i}\}_{i=1}^{B})$), while $V$ is trained to predict the $n$-step pessimistic return $z$ [^8] In the definition of $z$, $\tilde{r}^{(0)}=\tilde{r}$ is the pessimistic reward associated with $(s,h)$ and the subscript $(i)$ denotes that the corresponding quantity is collected in $i$ time steps within that trajectory.
-As noted in (Hubert et al.
-(2021)), $\pi_{\text{ret}}$ improves $\pi$ at each decision point, so repeatedly applying continuous BAMCP to obtain $\pi_{\text{ret}}$ and projecting the search results to the parameter space of $\pi$ (by minimizing $\mathcal{L}(\pi,\{\tau_{i}\}_{i=1}^{B})$) represents the policy improvement process.
-Meanwhile, value backups to update the Q estimates (i.e., $R\leftarrow V((s^{\prime},hars^{\prime}))$ in Algorithm 1) and the temporal difference learning for the value function $V$ constitute the policy evaluation process.
-In this way, the search algorithm is used as a powerful improvement operator to iteratively enhance the performance of the learned policy $\pi$.
-Algorithm 2 Search-based Policy Iteration
+Algorithm 2 BA-MCTS
 Input: $T$, $E_{l}$, $\mathcal{P}_{\theta}^{1:K}$, $\mathcal{R}_{\theta}^{1:K}$
-Initialize $\pi$ and $V$, $\mathcal{D}\leftarrow\emptyset$
+Initialize the actor $\pi$ and critic $Q_{\psi}$, $\mathcal{D}\leftarrow\emptyset$
 procedure Learner
@@
   while true do
-         $\{\tau_{i}\}_{i=1}^{B}\sim\mathcal{D}$
-         $\min_{\pi}\mathcal{L}(\pi,\{\tau_{i}\}_{i=1}^{B})$, $\min_{V}\mathcal{L}(V,\{\tau_{i}\}_{i=1}^{B})$
-[…]
-         Update $\pi,V$ in Actor if $e\%E_{l}==0$
+   $\{(s,h,\pi_{\text{ret}},\tilde{r},s^{\prime},h^{\prime})_{i}\}_{i=1}^{B}\sim\mathcal{D}$
+   Update $\pi$ and $Q_{\psi}$ using SAC
+   $e\mathrel{{+}{=}}1$
+   Update $\pi$ and $V$ (defined on $Q_{\psi}$) in Actor if
+   $e\%E_{l}==0$
   end while
 end procedure
-$\mathcal{L}(\pi,\{\tau_{i}\}_{i=1}^{B})=-\sum_{((s,h),\pi_{\text{ret}})}\pi_{\text{ret}}^{T}\log\pi(\cdot|(s,h))/(BT)$
-$\mathcal{L}(V,\{\tau_{i}\}_{i=1}^{B})=\sum_{((s,h),z)}(V(s,h)-z)^{2}/(BT)$,
-$z=\tilde{r}^{(0)}+\gamma\tilde{r}^{(1)}+\cdots+\gamma^{n-1}\tilde{r}^{(n-1)}+\gamma^{n}v_{\text{ret}}^{(n)}$
 procedure Actor
   while true do
-         $s\leftarrow\hat{\rho}_{0}(\cdot)$, $h\leftarrow s$, $\tau\leftarrow[]$
-         $b(\theta)\leftarrow[1/K,\cdots,1/K]$
-         for $t=1\cdots T$ do
+   Sample $s$ from $\mathcal{D}_{\mu}$, $h\leftarrow s$, $\tau\leftarrow[]$
+   Obtain the prior $b(\theta)$ at $s$
+   for $t=1\cdots H$ do
      $\pi_{\text{ret}},v_{\text{ret}}\leftarrow\textsc{Search}((s,h),b(\theta))$
@@
 end procedure
+4.3 Overall Framework: BA-MCTS
+Now, we present how to integrate planning outcomes of continuous BAMCP into a policy iteration framework, to obtain a near Bayes-optimal policy, i.e., $\pi$, that can be directly referred to during execution.
+The pseudo code is shown as Algorithm 2.
+For efficiency, a learner and a number of actors execute in parallel, reading from and sending data to the replay buffer $\mathcal{D}$ respectively.
+The actors update their copies of policy and value functions every $E_{l}$ learner steps.
+Each actor interacts with the learned world models to sample trajectory segments $\tau$.
+The starting states of these segments are sampled from the provided dataset $\mathcal{D}_{\mu}$.
+Notably: (1) the segment length $H$ is kept short to minimize error accumulation when interacting with the learned world models; and (2) the prior $b(\theta)$ at a starting state $s$ is obtained by performing the Bayesian posterior update (i.e., Eq.
+(4)) on the offline trajectory in $\mathcal{D}_{\mu}$ containing $s$.
+These belief updates are reliable, as the offline trajectories are from the real environment.
+At each decision node $(s,h)$ of the segment, a Search procedure (defined in Algorithm 1) is executed.
+The search result $\pi_{\text{ret}}$ then indicates the action choice, i.e., $a\sim\pi_{\text{ret}}(\cdot|(s,h))$.
+Given $a$, the transition to the next state follows a BAMDP, where $r\sim\mathcal{R}^{+}(\cdot|(s,h),a)$, $s^{\prime}\sim\mathcal{P}^{+}(\cdot|(s,h),a)$, and the belief $b({\theta})$ is adapted with the new transition.
+With the collected segments $\tau$, we update the policy $\pi$ and value function $V$.
+Our primary algorithm, BA-MCTS, employs an off-policy RL method: SAC (Haarnoja et al., 2018).
+The value function (critic) is updated using standard temporal difference learning based on the sampled transitions, while the policy (actor) is updated using the policy gradient derived from the critic.
+Additionally, we investigate an alternative approach, BA-MCTS-SL, where the policy is updated via supervised learning to mimic the search result $\pi_{\text{ret}}$ by minimizing a cross-entropy loss: $\mathcal{L}(\pi,\{(s,h,\pi_{\text{ret}})_{i}\}_{i=1}^{B})=-\sum_{((s,h),\pi_{\text{ret}})}\pi_{\text{ret}}^{T}\log\pi(\cdot|(s,h))/B$.
+We compare these two approaches in our evaluation.
+In both methods, we seek to distill the superior, non-parametric search outcomes from Continuous BAMCP into the parameterized policy $\pi$.
 5 Evaluation
-To evaluate the effectiveness of each component in our algorithm design, we introduce three variants: (1) BA-MBRL leverages learned world models as surrogate simulators, applying reward penalties to collected transitions and using standard online RL algorithms (e.g., SAC (Haarnoja et al.
-(2018))) to learn a policy.
-While following existing offline MBRL methods, it models the problem as a BAMDP (rather than an MDP), with environment transitions defined by Equations (1) and (4) and the reward penalty by Equation (5), and is designed to evaluate the effectiveness of Bayesian RL.
-(2) BA-MCTS builds on BA-MBRL by introducing Continuous BAMCP (Algorithm 1) to plan at decision points, rather than inferring directly from the policy, to generate trajectories for downstream SAC, demonstrating the impact of deep search on policy learning.
-(3) BA-MCTS-SL, described in Algorithm 2, replaces the policy learning algorithm in BA-MCTS from policy gradient methods (as in SAC) with supervised learning (SL), allowing us to compare which approach offers a more efficient policy update mechanism, particularly for continuous control tasks.
-Data Type Environment BA-MCTS -SL (ours) BA-MCTS (ours) BA-MBRL (ours) Optimized COMBO MOReL MOPO
-random HalfCheetah 29.20 $\pm$ 2.00 36.23 $\pm$ 1.04 32.76 $\pm$ 1.16 31.7 38.8 25.6 35.4
-random Hopper 33.83 $\pm$ 0.10 31.56 $\pm$ 0.12 31.47 $\pm$ 0.03 12.1 17.9 53.6 11.7
-random Walker2d 21.89 $\pm$ 0.07 21.59 $\pm$ 0.32 21.45 $\pm$ 0.53 21.7 7.0 37.3 13.6
-medium HalfCheetah 70.47 $\pm$ 3.52 75.84 $\pm$ 3.81 56.54 $\pm$ 5.20 45.7 54.2 42.1 42.3
-medium Hopper 97.75 $\pm$ 7.09 96.70 $\pm$ 14.0 98.25 $\pm$ 3.42 69.3 97.2 95.4 28.0
-medium Walker2d 82.24 $\pm$ 1.85 74.73 $\pm$ 3.25 75.41 $\pm$ 4.17 79.7 81.9 77.8 17.8
-med-replay HalfCheetah 61.16 $\pm$ 1.60 65.45 $\pm$ 0.81 62.50 $\pm$ 0.18 58.0 55.1 40.2 53.1
-med-replay Hopper 106.3 $\pm$ 0.13 101.8 $\pm$ 3.46 93.91 $\pm$ 4.25 90.8 89.5 93.6 67.5
-med-replay Walker2d 92.13 $\pm$ 5.13 95.06 $\pm$ 2.11 97.54$\pm$ 1.93 65.8 56.0 49.8 39.0
-med-expert HalfCheetah 80.53 $\pm$ 6.63 76.16 $\pm$ 10.3 90.52 $\pm$ 4.13 104.2 90.0 53.3 63.3
-med-expert Hopper 112.2 $\pm$ 0.29 108.3 $\pm$ 0.22 107.8 $\pm$ 0.37 105.8 111.1 108.7 23.7
-med-expert Walker2d 107.7 $\pm$ 0.82 110.0 $\pm$ 1.74 84.71 $\pm$ 0.87 97.1 103.3 95.6 44.6
-Average Score 74.62 74.45 71.06 65.16 66.83 64.42 36.67
-Table 1: Comparisons between the proposed algorithms and SOTA offline model-based RL methods on the D4RL benchmark suite.
-Each value represents the normalized score, as proposed in (Fu et al.
-(2020)), of the policy trained by the corresponding algorithm.
-These scores are undiscounted returns normalized to approximately range between 0 and 100, where a score of 0 corresponds to a random policy and a score of 100 corresponds to an expert-level policy.
-For our algorithms, we report the average score of the final ten policy learning epochs and its standard deviation across three random seeds.
-Results in the last four columns are taken from the original papers (Lu et al.
-(2022); Yu et al.
-(2021); Kidambi et al.
-(2020); Yu et al.
-(2020)), respectively.
-We first evaluate our algorithms on a widely-used continuous control benchmark for offline RL methods – D4RL MuJoCo (Fu et al.
-(2020)).
-The evaluation results for three types of robotic agents, each with offline datasets of four different qualities, are presented in Table 1.
-(1) Compared to SOTA offline MBRL methods, our algorithms achieve superior performance on nine out of twelve tasks.
-In terms of average performance, BA-MBRL significantly outperforms the baselines, demonstrating the effectiveness of using BAMDPs to handle model uncertainties in offline MBRL.
-(2) Both BA-MCTS and BA-MCTS-SL further improve upon BA-MBRL, highlighting the enhancement brought by deep search in policy learning.
-Notably, we apply Continuous BAMCP to only 10% of states when collecting training trajectories, while for the remaining states, we sample actions directly from the policy, i.e., $a\sim\pi(\cdot|s)$.
-Increasing the search ratio could further enhance policy performance at the cost of increased computation.
-(3) BA-MCTS-SL performs similarly to BA-MCTS, validating the effectiveness of both policy update mechanisms.
-However, BA-MCTS-SL struggles on Walker2d, where a warm-up training phase (using BA-MBRL) is required to establish a better initial policy.
-On the other hand, the advantage of the SL-based policy update is evident in the training plots of our algorithms in Figure 3, where BA-MCTS-SL exhibits much smoother learning curves compared to the other two algorithms, indicating greater robustness in model selection.
-(4) We further compare our algorithms with model-free offline policy learning methods, as shown in Appendix D.
-The performance improvement is even greater than that over model-based methods, highlighting the necessity of model-based learning.
-Particularly, when data quality is low, merely mimicking or staying close to the behavior policy would result in an underperforming policy.
-(5) We provide an ablation study in Appendix F to demonstrate the necessity of incorporating the reward penalty in offline MBRL to prevent the overexploitation of inaccurate world models.
-Additionally, we find that the SL-based policy update is less sensitive to model inaccuracies.
-(a) hc-med-expert
-(b) hc-med-replay
-(c) hc-medium
-(d) hc-random
-(e) hp-med-expert
-(f) hp-med-replay
-(g) hp-medium
-(h) hp-random
-(i) wk-med-expert
-(j) wk-med-replay
-(k) wk-medium
-(l) wk-random
-Figure 1: Performance of Sampled EfficientZero on D4RL MuJoCo tasks.
-The results for HalfCheetah, Hopper, and Walker2d are presented in the three rows, respectively.
-Each subfigure depicts the change in undiscounted episodic return as a function of the number of training samples.
-Experiments are repeated three times with different random seeds, with the solid line representing the mean and the shaded area indicating the 95% confidence interval.
-For reference, the expert-level episodic returns for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively.
-Figure 2: Evaluation results for the tokamak control tasks.
-The figure shows the change in episodic returns over training epochs for the proposed algorithms and baselines across three target tracking tasks in the nuclear fusion scenario.
-Solid lines represent the average performance, while shaded areas indicate the 95% confidence intervals.
-For fair comparisons and real-time execution, we do not perform test-time search and adopt the same policy architecture as the baselines, i.e., a feedforward neural network, rather than an RNN that incorporates transition history as input.
-These alternative designs have the potential to further improve our algorithms.
-For implementation, we build on the codebase of Optimized (Lu et al.
-(2022)), which thoroughly explores design choices in offline MBRL, making minimal changes to the code and hyperparameter settings [^9] The detailed hyperparameter setups of our algorithms are provided in Appendix C..
-Therefore, we believe the performance improvements stem from the Bayesian RL framework and deep search component.
-Further, both components can be seamlessly integrated with other advancements in offline MBRL, such as more accurate world model learning and improved uncertainty quantification for constructing pessimistic MDPs.
+Section 5.1 discusses the benefit of Bayesian model adaptation.
+Section 5.2 presents comprehensive benchmarking results on D4RL MuJoCo tasks.
+Section 5.3 details our ablation studies.
+Finally, Section 5.4 validates our algorithm’s applicability on a challenging Tokamak control task.
+Table 1: Comparison of the prediction errors for next states and rewards in imaginary rollouts, with and without Bayesian belief adaptation.
+Ground truth for the imaginary rollouts is obtained by replaying the action sequences in the real simulators.
+Data Type Environment Prediction Error on Next State Prediction Error on Reward Overall Prediction Error
+Adaptive Uniform Adaptive Uniform Adaptive Uniform
+random HalfCheetah 0.339 (.021) 0.342 (.017) 0.016 (.001) 0.016 (.001) 0.177 (.011) 0.179 (.009)
+random Hopper 0.232 (.016) 0.189 (.008) 0.012 (.001) 0.022 (.008) 0.122 (.008) 0.106 (.008)
+random Walker2d 62.10 (11.0) 112.0 (19.0) 1.364 (.215) 4.908 (.708) 31.73 (5.61) 58.44 (9.83)
+medium HalfCheetah 1.387 (.040) 1.355 (.038) 0.057 (.001) 0.061 (.001) 0.722 (.021) 0.708 (.019)
+medium Hopper 0.429 (.033) 0.570 (.035) 19.03 (8.58) 68.24 (11.4) 9.727 (4.30) 34.40 (5.69)
+medium Walker2d 34.01 (.556) 34.26 (.142) 24.38 (4.31) 113.1 (3.47) 29.19 (1.89) 73.69 (1.78)
+med-replay HalfCheetah 0.677 (.027) 0.707 (.036) 0.115 (.005) 0.114 (.006) 0.396 (.016) 0.410 (.021)
+med-replay Hopper 0.170 (.022) 0.212 (.054) 1.177 (.420) 3.320 (1.14) 0.674 (.221) 1.766 (.558)
+med-replay Walker2d 66.83 (7.39) 44.61 (1.34) 27.21(3.24) 60.52 (3.17) 47.02 (5.17) 52.57 (2.25)
+med-expert HalfCheetah 1.423 (.054) 1.467 (.046) 0.071 (.002) 0.074 (.002) 0.747 (.028) 0.771 (.024)
+med-expert Hopper 3.665 (2.56) 18.27 (2.09) 71.46 (79.9) 384.7 (37.6) 37.56 (41.2) 201.5 (19.4)
+med-expert Walker2d 55.69 (3.71) 51.89 (.481) 121.9 (12.6) 186.3 (3.86) 88.77 (8.18) 119.1 (2.08)
+5.1 The Benefit of Belief Adaptation
+The core idea of belief adaptation (using the BAMDP framework detailed in Section 4.1) is that by dynamically adjusting the belief over each ensemble member based on observed transitions, the agent can rely more on models that are accurate for the current trajectory, leading to better planning.
+Our empirical results show that belief adaptation not only improves model prediction accuracy, as shown in Table 1, but also translates to better final policy performance on offline RL tasks, as shown in Tables 2 and 4.
+Specifically, Table 1 highlights that the adaptive ensemble achieves more accurate predictions for next states and rewards in 10 out of 12 benchmarking tasks, compared to treating each ensemble member uniformly as in previous offline MBRL methods.
+This suggests that Bayesian belief adaptation leads to a more accurate world model for planning, which is a key factor underlying observed performance improvements.
+A more detailed study of this aspect, including a visualization of the belief adaptation process within rollouts, is provided in Appendix F.1.
+5.2 Benchmarking Results on D4RL MuJoCo
+We evaluate our algorithm on a widely-used offline RL benchmark: D4RL MuJoCo (Fu et al., 2020), comparing its performance with state-of-the-art (SOTA) model-based/model-free methods, and search-based approaches.
+Model-Based Methods.
+In Table 2, BA-MCTS achieves the highest average score across all twelve D4RL MuJoCo tasks when compared against a range of SOTA model-based baselines.
+This result confirms that the combination of a principled Bayesian treatment of uncertainty and a powerful search-based planner leads to superior performance.
+Table 2: Comparisons of our algorithm against SOTA offline RL methods on the D4RL benchmark.
+Each value represents the normalized score, as proposed in Fu et al.
+(2020), of the policy trained by the algorithm.
+For our algorithm, we report the average score of the final ten learning epochs and its standard deviation across three random seeds.
+To ensure a fair comparison, the results for MOBILE and CBOP are from our own runs using their publicly codebases.
+The results for all other baselines are taken from their respective original papers.
+Please check Appendix F for the references.
+Data Type Environment BA-MCTS (ours) MOBILE CBOP RAMBO APE-V MAPLE Optimized
+random HalfCheetah 41.45 (1.40) 39.27 (1.38) 32.95 (0.14) 40.0 29.9 41.5 31.7
+random Hopper 31.42 (0.10) 7.965 (0.43) 23.79 (11.2) 21.6 31.3 10.7 12.1
+random Walker2d 21.54 (0.06) 21.47 (0.03) 1.391 (0.06) 11.5 15.5 22.1 21.7
+medium HalfCheetah 77.31 (1.40) 76.42 (1.00) 69.98 (0.88) 77.6 69.1 48.5 45.7
+medium Hopper 103.9 (0.33) 102.5 (1.77) 98.68 (2.30) 92.8 - 44.1 69.3
+medium Walker2d 88.23 (0.78) 87.12 (2.79) 93.92 (1.92) 86.9 90.3 81.3 79.7
+med-replay HalfCheetah 71.24 (2.65) 64.67 (8.14) 63.36 (0.47) 68.9 64.6 69.5 58.0
+med-replay Hopper 106.4 (0.53) 105.7 (1.56) 101.3 (1.74) 96.6 98.5 85.0 90.8
+med-replay Walker2d 91.42 (1.54) 85.12 (10.1) 74.56 (1.29) 85.0 82.9 75.4 65.8
+med-expert HalfCheetah 102.3 (2.27) 100.9 (4.00) 97.91 (5.26) 93.7 101.4 55.4 104.2
+med-expert Hopper 111.8 (0.82) 111.9 (0.34) 111.2 (0.77) 83.3 105.7 95.3 105.8
+med-expert Walker2d 116.0 (1.49) 115.0 (0.30) 117.6 (0.27) 68.3 110.0 107.0 97.1
+Average Score 80.25 76.50 73.87 68.85 72.65 61.32 65.16
+Model-Free Methods.
+We also compare our algorithms with a series of model-free offline policy learning methods, with detailed results presented in Appendix F.2.
+Our algorithms show significantly better performance, demonstrating the necessity of model-based learning in these environments, particularly when data quality is low.
+Notably, APE-V (see Table 2) is a model-free method built on BAMDPs, yet it shows inferior performance compared to ours in 10 out of 12 tasks.
+As a summary, in Table 3444MOReL, IQL, and the model-free offline RL baselines in Appendix F.2 did not report standard deviations.
+Additionally, IQL and DT were not evaluated on D4RL tasks with random data., we show comparisons against 19 strong model-based and model-free baselines, where BA-MCTS achieves the highest average score.
+(See Appendix F.3 for details of the selected baselines.) To quantify the significance of this improvement, we compute Cohen’s $d$ (Cohen, 1988) against all baselines for which standard deviations are available.
+The results show that the Cohen’s $d$ value is greater than 1.8 in all cases.
+According to Cohen’s Rule of Thumb, a value of $d>1$ indicates a very large (statistically significant) difference.
+This comprehensive comparison demonstrates that our proposed method establishes a new SOTA in offline RL.
+Table 3: Comprehensive performance comparison on D4RL MuJoCo.
+Average scores for our algorithm and baselines are reported.
+For baselines that report standard deviations, we compute Cohen’s $d$ to quantify the statistical significance of the improvement achieved by our method.
+A value of $d>1$ indicates a very large and statistically significant difference.
+Algorithm BA-MCTS Optimized COMBO MOReL MOPO CQL BEAR
+Avg. Score 80.3 (1.1) 65.2 (5.4) 66.8 (3.4) 64.4 36.7 (11.0) 54.9 39.8
+Avg. (No Random) 96.5 (1.3) - - - - - -
+Cohen’s $d$ - 3.5 4.6 - 4.7 - -
+Algorithm BRAC-v SAC BC MOBILE CBOP RAMBO APE-V
+Avg. Score 31.4 4.1 25.1 76.5 (2.7) 73.9 (2.2) 68.9 (2.8) 72.7 (1.8)
+Avg. (No Random) - - - - - - -
+Cohen’s $d$ - - - 1.8 3.7 4.8 4.8
+Algorithm MAPLE TD3+BC IQL SAC-N EDAC DT
+Avg. Score 61.3 (1.8) 54.6 (2.4) - 76.3 (0.6) 76.0 (2.3) -
+Avg. (No Random) - - 77.0 - - 74.7 (1.2)
+Cohen’s $d$ 12.7 12.4 - 4.8 2.2 17.4
+Table 4: Comparisons on the target tracking tasks.
+For each algorithm, we report the average return of the final ten policy learning epochs and its standard deviation across three different random seeds.
+Task BA-MCTS (Ours) BA-MCTS-SL (Ablation) BA-MBRL (Ablation) CQL Optimized
+Temperature -23.83 (9.66) -21.16 (5.00) -29.35 (4.72) -59.62 (1.57) -83.55 (10.56)
+Rotation -19.07 (5.85) -14.14 (1.88) -31.33 (11.54) -85.48 (2.72) -71.54 (9.88)
+$\beta_{n}$ -18.93 (1.75) -37.03 (17.98) -23.40 (10.77) -36.37 (1.17) -57.84 (10.27)
+Average -20.61 -24.11 -28.03 -60.49 -70.98
+Search-Based Methods.
 MuZero also applies deep search to MBRL.
-To evaluate its performance on D4RL MuJoCo tasks, we use the open-source implementation and hyperparameter configurations of Sampled EfficientZero (Ye et al.
-(2021)) provided by LightZero (Niu et al.
-(2023)).
-Benchmarking results from LightZero indicate that Sampled EfficientZero, equipped with a Gaussian policy, achieves the best performance on (online) MuJoCo locomotion tasks compared to other MuZero variants.
-To adapt Sampled EfficientZero for offline learning, we employ the reanalyse technique proposed by (Schrittwieser et al.
-(2021)).
-The evaluation results are presented in Figure 1.
-For reference, the expert-level episodic returns (corresponding to scores of 100) for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively.
-As shown, the results are significantly worse compared to the performance of offline RL methods listed in Table 1, despite Sampled EfficientZero’s higher computational cost (detailed in Appendix E).
-Notably, both Sampled EfficientZero and BA-MCTS-SL rely on supervised learning for policy improvement.
-However, for continuous control tasks, the agent can only sample a finite number of actions at a decision point, and the search result (e.g., $\pi_{\text{ret}}$ in Algorithm 1) is a distribution over this finite set, which could be a poor approximation of the optimal action distribution.
-Thus, purely mimicking the search result may be less sample-efficient than policy gradient methods, as it fails to account for the continuous nature of the action space.
-Furthermore, world model learning is the foundation of MBRL and can be particularly challenging in continuous control and offline learning settings, where the state-action space is vast but training data is limited.
-Sampled EfficientZero integrates model learning and policy training into a single stage, which significantly increases the learning difficulty (compared to BA-MCTS-SL).
-Finally, we evaluate our algorithms on three target tracking tasks in tokamak control.
-The tokamak is one of the most promising confinement devices for achieving controllable nuclear fusion, where the primary challenge lies in confining the plasma, i.e., an ionized gas of hydrogen isotopes, while heating it and increasing its pressure to initiate and sustain fusion reactions (Pironti & Walker (2005)).
-Tokamak control involves applying a series of direct actuators (e.g., neutral beam, ECH power, magnetic field) and indirect actuators (e.g., setting targets for the plasma shape and density) to confine the plasma to achieve a desired state or track a given target.
-This sophisticated physical process is an ideal test bed for our algorithms.
-Specifically, we use a well-trained data-driven dynamics model provided by Char et al.
-(2024) as a “ground truth" simulator for the nuclear fusion process during evaluation, and generate a dataset containing 725270 transitions using this model for offline RL.
-We select a reference shot (i.e., an episode of a fusion process) from DIII-D [^10] DIII-D is a tokamak device located in San Diego, California, operated by General Atomics., and use its trajectories of Ion Rotation, Electron Temperature, and $\beta_{n}$ as targets for three tracking tasks.
-These are critical quantities in tokamak control, particularly $\beta_{n}$, which serves as an economic indicator of the efficiency of nuclear fusion.
+To evaluate its performance on D4RL MuJoCo, we use the open-source implementation and hyperparameter configurations of Sampled EfficientZero (Ye et al., 2021) provided by LightZero (Niu et al., 2023).
+Benchmarking results from LightZero indicate that Sampled EfficientZero achieves the best performance on (online) MuJoCo tasks compared to other MuZero variants.
+To adapt Sampled EfficientZero for offline learning, we employ the reanalyse technique proposed by Schrittwieser et al.
+(2021).
+The evaluation results are presented in Figure 2 (Appendix F.4).
+As shown, the results are significantly worse compared to the performance of offline RL methods listed in Table 2, despite Sampled EfficientZero’s much higher computational cost.
+In Appendix F.4, we provide a detailed comparison of the computational costs and algorithm designs between our method and Sampled EfficientZero.
+World model learning is the foundation of MBRL and can be particularly challenging in continuous control and offline settings, where the state-action space is vast but training data is limited.
+Sampled EfficientZero integrates model learning and policy training into a single stage, which significantly increases the learning difficulty.
+5.3 Ablation Study
+For a principled ablation study, we reimplement BA-MCTS and its ablated variants based on the codebase of Optimized (Lu et al., 2022), which thoroughly explores design choices in offline MBRL, making minimal changes to the code and hyperparameter settings (see Appendix D).
+Specifically, we first add a belief adaptation module to Optimized to obtain BA-MBRL, and then incorporate the Continuous BAMCP planner (Section 4.2) to create BA-MCTS.
+Finally, we replace the policy distillation module of BA-MCTS from SAC to supervised learning (as detailed in Section 4.3), yielding BA-MCTS-SL.
+The detailed results are presented in Table 11, and Appendices F.5 - F.6.
+BA-MBRL significantly outperforms Optimized, demonstrating the necessity of Bayesian model adaptation.
+BA-MCTS further improves upon the performance of BA-MBRL, underscoring the effectiveness of employing deep search as a powerful policy improvement operator.
+Finally, BA-MCTS-SL achieves performance comparable to BA-MCTS.
+However, BA-MCTS-SL can be less stable on certain complex tasks: it struggles on the Walker2d environments, where a warm-up phase (using BA-MBRL) is needed to establish a good initial policy for supervised learning to be effective.
+Finally, we conduct an ablation study to assess the necessity of the reward penalty.
+By removing this penalty term (i.e., setting $\lambda=0$ in Eq.
+(5)), we observe a consistent drop in average performance.
+This demonstrates that the pessimistic reward penalty is crucial for stopping the agent exploiting inaccuracies in the learned world models.
+Results are in Table 12 in Appendix F.7.
+5.4 Application to Tokamak Control
+We evaluate our algorithm on three target tracking tasks in tokamak control.
+The tokamak is one of the most promising confinement devices for achieving controllable nuclear fusion.
+We use a well-trained data-driven dynamics model provided by Char et al.
+(2024) as a “ground truth” simulator for the nuclear fusion process during evaluation, and generate a dataset (by this dynamics model) containing 725,270 transitions for offline RL.
+We select a reference shot (i.e., an episode of a fusion process) from DIII-D (a tokamak device located in San Diego), and use its trajectories of Ion Rotation, Electron Temperature, and $\beta_{n}$ as targets for three tracking tasks.
 The tracking tasks have a 28-dimensional state space and a 14-dimensional action space, both continuous.
 Moreover, these tasks are highly stochastic, as the underlying dynamics model is a probabilistic neural network and each state transition is a sample from this model.
-For details on the simulator, and the design of the state/action spaces and reward functions, please refer to Appendix G.
-We compare our algorithms with SOTA model-free and model-based offline RL methods, specifically CQL and Optimized.
-The learning performance on the three tracking tasks is shown in Figure 2, where the x-axis and y-axis represent the training epochs and (negative) full-shot tracking errors, respectively.
-Our algorithms consistently outperform the baselines.
-Notably, the offline dataset does not include the reference shot or any similar, nearby shots.
-Therefore, restricting the policy to stay close to the behavior policy, as done in model-free methods, can be problematic.
-Also, learning dynamics models for MBRL is quite challenging in this nuclear fusion scenario.
-Our algorithms share the same ensemble of dynamics models with “Optimized" for policy learning, and the comparisons can demonstrate the superiority of Bayesian RL and deep search.
-Figure 2 has been smoothed for visualization [^11] The episodic return is plotted every 10 training epochs, with the y-axis representing the average value of a sliding window of length 5..
-We further report the average return over the final 10 training epochs in Table 2, and the conclusions align with those from the D4RL MuJoCo tasks, showing the robustness of our proposed algorithms.
-Task BA-MCTS -SL (ours) BA-MCTS (ours) BA-MBRL (ours) CQL Optimized
-Temperature -21.16 $\pm$ 5.00 -23.83 $\pm$ 9.66 -29.35 $\pm$ 4.72 -59.62 $\pm$ 1.57 -83.55 $\pm$ 10.56
-Rotation -14.14 $\pm$ 1.88 -19.07 $\pm$ 5.85 -31.33 $\pm$ 11.54 -85.48 $\pm$ 2.72 -71.54 $\pm$ 9.88
-$\beta_{n}$37.03 $\pm$ 17.98 -18.93 $\pm$ 1.75 -23.4 $\pm$ 10.77 -36.37 $\pm$ 1.17 -57.84 $\pm$ 10.27
-Average -24.11 -20.61 -28.03 -60.49 -70.98
-Table 2: Comparisons between the proposed algorithms and offline RL baselines on the target tracking tasks.
-For each algorithm, we report the average return of the final ten policy learning epochs and its standard deviation across three different random seeds.
-6 Conclusion and Discussions
-In this work, we propose framing offline model-based reinforcement learning (MBRL) as a Bayes Adaptive Markov Decision Process (BAMDP) to better address uncertainties in the world models learned from offline datasets.
-We also introduce a novel planning method for solving BAMDPs in continuous state and action spaces using Monte Carlo Tree Search.
-This planning process is integrated into a policy iteration framework, enabling the derivation of a policy suitable for real-time execution from the planning results.
+For details on the simulator, and the design of the state/action spaces and reward functions, please refer to Appendix E.
+We obtained access to the DIII-D data for a limited time window, which restricted our ability to provide evaluation results across a wide range of algorithms.
+For a principled comparison with the baselines, we utilized the codebase from the ablation study, where our algorithm implementations are primarily based on Optimized, to ensure that performance improvements are attributable to algorithmic design rather than implementation differences.
+Additionally, we included evaluation of a widely-used model-free offline RL method: CQL.
+The average tracking errors over the final 10 training epochs are reported in Table 4, where BA-MCTS outperforms all baselines.
+These results demonstrate the potential of our algorithm for application in complex, data-driven control problems.
+6 Conclusion
+We propose framing offline MBRL as a BAMDP to better address uncertainties in the world models learned from offline datasets.
+We also introduce a novel planning method for solving BAMDPs in continuous state and action spaces using MCTS.
+This planning process is integrated into a policy iteration framework, enabling the derivation of a policy function for real-time execution.
 In our evaluation, we test several variants of our algorithms to separately highlight the effectiveness of Bayesian RL and deep search.
-Additionally, we compare two different approaches for policy updates (based on the search results) in continuous control tasks: supervised learning and policy gradient methods.
-Our findings demonstrate that: (1) adapting beliefs over an ensemble of world models based on experience yields more accurate model approximations for MBRL; (2) deep search improves learning performance by incorporating planning and additional computation input; and (3) while supervised-learning-based policy updates result in smoother learning curves, they may struggle in complex continuous control tasks due to their approximation of the continuous action space as a finite set of action samples.
-For future work, our algorithms can be further improved by integrating advancements in offline MBRL and Bayesian RL, such as Bayesian RL methods that do not rely on deep ensembles, techniques to address sparse rewards in MBRL, and more principled approaches to construct pessimistic MDPs beyond those based on ensemble discrepancy.
+Additionally, we validate the superior performance of our approach on both standard offline RL benchmarks and challenging continuous control tasks in Nuclear Fusion.
 Appendix A BAMCP
@@
 Input: $\pi$, $E$, $d_{\text{max}}$, $\mathcal{R}$, $\gamma$, $\mathcal{A}$, $c$
-procedure Search($(s,h),b(\theta)$)
+[…]
   for $e=1\cdots E$ do
    $\theta\sim b(\cdot)$
-         Simulate($(s,h),\theta,d_{\text{max}}$)
+[…]
   end for
@@
 end procedure
-procedure Rollout($(s,h),\theta,d$)
+[…]
   if $d==0$ then
@@
 end procedure
-procedure Simulate($(s,h),\theta,d$)
+[…]
   if $d==0$ then
@@
   $R\leftarrow r+\gamma\textsc{Simulate}((s^{\prime},has^{\prime}),\theta,d-1)$
-[…]
+  $N((s,h))\mathrel{{+}{=}}1,N((s,h),a)\mathrel{{+}{=}}1$
   $Q((s,h),a)\mathrel{{+}{=}}\frac{R-Q((s,h),a)}{N((s,h),a)}$
@@
 end procedure
-Bayes Adaptive Monte Carlo Planning (BAMCP, Guez et al.
-(2013)) is a sample-based online planning method, aiming to find the action $a^{*}$ that approximately maximizes the expected return at a decision point $(s,h)$ under the BAMDP.
+Bayes Adaptive Monte Carlo Planning (BAMCP) (Guez et al., 2013) is a sample-based online planning method, aiming to find the action $a^{*}$ that approximately maximizes the expected return at a decision point $(s,h)$ under the BAMDP.
 Its detailed pseudo code is shown as Algorithm 3.
 BAMCP has demonstrated success in solving BAMDPs with large-scale discrete state and action spaces.
-Its key algorithmic ideas include: (1) applying MCTS with an efficient exploration strategy – UCT (Kocsis & Szepesvári (2006)) to the BAMDP in order to simulate the outcomes of different action choices; (2) utilizing root sampling to avoid frequent Bayesian posterior updates.
+Its key algorithmic ideas include: (1) applying MCTS with an efficient exploration strategy – UCT (Kocsis and Szepesvári, 2006) to the BAMDP in order to simulate the outcomes of different action choices; (2) utilizing root sampling to avoid frequent Bayesian posterior updates.
 Specifically, the UCT rule is used for selecting actions at non-leaf nodes, i.e., $a\leftarrow\arg\max_{x}Q((s,h),x)+c\sqrt{\frac{\log N((s,h))}{N((s,h),x)}}$, managing the tradeoff between exploration and exploitation.
-Root sampling refers to sampling the dynamics model only at the root node (i.e., $\theta\sim b(\cdot)$) and not adapting the belief $b(\cdot)$ according to the Bayes rule during the search process, of which the rationality is justified in the following lemma.
+(As shown in Algorithm 3, $N$ is the visit count, while $Q$ denotes the accumulated reward.) Root sampling refers to sampling the dynamics model only at the root node (i.e., $\theta\sim b(\cdot)$) and not adapting the belief $b(\cdot)$ according to the Bayes rule and experience during the search process, of which the rationality is justified in the following lemma.
 Lemma A.1.
@@
 $=Z\tilde{b}(\theta|ha)\mathcal{P}_{\theta}(s^{\prime}|s,a)=\tilde{b}(\theta|has^{\prime})$
+Here, $Z=1/\int_{\theta}\mathcal{P}_{\theta}(s^{\prime}|s,a)b(\theta|h)d\theta=1/\int_{\theta}\mathcal{P}_{\theta}(s^{\prime}|s,a)\tilde{b}(\theta|h)d\theta=1/\int_{\theta}\mathcal{P}_{\theta}(s^{\prime}|s,a)\tilde{b}(\theta|ha)d\theta$ is the normalization constant.
+The fifth equality in Eq.
+(6) holds due to the inductive hypothesis.
+The sixth equality is based on the fact that the choice of $a$ at each node $h$ is made independently of the sample $\theta$.
+As for the last equality, to experience $\theta$ at $has^{\prime}$, the sample $\theta$ needs to traverse $ha$ (with probability $\tilde{b}(\theta|ha)$) and then the state $s^{\prime}$ needs to be sampled, which is with probability $\mathcal{P}_{\theta}(s^{\prime}|s,a)$, so $\tilde{b}(\theta|has^{\prime})\propto\tilde{b}(\theta|ha)\mathcal{P}_{\theta}(s^{\prime}|s,a)$.
+∎
+Appendix B Consistency Proof of Continuous BAMCP
+This appendix provides the detailed proof for Theorem 4.1.
+Our proof adapts the consistency proof methodology developed for Polynomial Upper Confidence Trees (PUCT) (Auger et al., 2013).
+To apply this framework, we first note that the Pessimistic Bayes-Adaptive MDP (P-BAMDP) defined in the main text can be formally treated as a standard, albeit large-state-space, MDP.
+Throughout this proof, we will refer to this mathematical object as the Pessimistic MDP (P-MDP).
+Our goal is to show that Continuous BAMCP is a consistent solver for this P-MDP.
+The remainder of this appendix is organized as follows.
+Section B.1 formally defines the P-MDP and introduces the key notations, assumptions, and the induction property central to our proof.
+Section B.2 establishes the base case for our backward induction.
+Sections B.3 and B.4 present the two core inductive steps, showing how the consistency property propagates upward through the random and decision nodes of the search tree, respectively.
+Finally, Section B.5 concludes the proof.
+B.1 Preamble and Definitions
+P-MDP Formalization
+The P-MDP that our planner solves is defined by the following components:
+  • State Space $\mathcal{S}^{+}$: The augmented information state space, where each state is $(s,h)$.
+  • Action Space $\mathcal{A}$: The action space of the original problem.
+  • Transition Function $\mathcal{P}^{+}$: The belief-weighted transition function over information states.
+  • Reward Function $\tilde{r}$: The fixed, bounded, penalized reward function from Eq. (5)
+Relevant Notations
+We adopt the following notations:
+  • $z$: A decision node, corresponding to an information state $(s,h)$.
+  • $w$: A random node, corresponding to a state-action pair $(z,a)$.
+  • $V^{*}_{p}(z),V^{*}_{p}(w)$: The optimal Bellman value of node $z$ and $w$ under the P-MDP.
+  • $n(z),n(w)$: The total number of visits to node $z$ and $w$ during MCTS.
+  • $\hat{V}(z),\hat{V}(w)$: The empirical average return obtained at node $z$ and $w$.
+  • $V(z)$: The value of a node $z$ estimated by the learned value function from the previous policy iteration.
+  • $d$: The depth of a node in the search tree, integer for decision node while semi-integer for random node
+Definition B.1 (Exponentially sure in $n$).
+We say that some property $(P)$ depending on an integer $n$ is exponentially sure in $n$ (denoted e.s.) if there exists positive constant $C,h$ such that the probability that $(P)$ holds is at least
+$1-C\exp(-hn).$
+Our proof relies on the following assumptions:
+Assumption B.2 (Policy Regularity).
+For any decision node $z=(s,h)$ and any precision $\Delta>0$, the policy $\pi(\cdot|z)$ used to sample new actions has a non-zero probability of sampling a $\Delta$-optimal action within the current P-MDP.
+Formally, there exist fixed constants $\theta>0$ and $p>1$ such that,
+$V^{*}_{p}(w=(z,a))\geq V^{*}_{p}(z)-\Delta\text{ with probability at least }\min(1,\theta\Delta^{p})$ (7)
+Assumption B.3 (Bounded Value Function Error).
+In our algorithm, when a Monte Carlo Tree Search simulation reaches a leaf node $z$ of the current search tree (i.e., a node that has not yet been expanded), we use a learned value function, $V$, to estimate its value for the backup step.
+This function is trained as part of the overall policy iteration framework.
+We assume that $V$ is an imperfect approximator of the P-MDP’s true optimal value function, $V^{*}_{p}$, and that its approximation error is uniformly bounded by a constant $\epsilon_{V}\geq 0$ over the entire state space:
+$|V(z)-V^{*}_{p}(z)|\leq\epsilon_{V},\quad\forall z\in\mathcal{S}^{+}$
+Furthermore, we assume the output of $V(z)$ is bounded.
+The proof aims to establish the following property via backward induction:
+Definition B.4 (Induction Property Cons($\gamma,f,d$)).
+There exists a constant $C_{d}>0$ and a function $f_{d}(\cdot)$ such that for all decision node $z$ at depth $d$,
+$|\hat{V}(z)-V^{*}_{p}(z)|\leq C_{d}n(z)^{-\gamma_{d}}+f_{d}(\epsilon_{V})\text{ e.s. in }n(z)$
+and for all nodes $w$ at semi-integer depth $d+\frac{1}{2}$,
+$|\hat{V}(w)-V^{*}_{p}(w)|\leq C_{d+\frac{1}{2}}n(w)^{-\gamma_{d+\frac{1}{2}}}+f_{d+\frac{1}{2}}(\epsilon_{V})\text{ e.s. in }n(w)$
+where $\epsilon_{V}$ is the base approximation error of the value function (from Assumption B.3).
+B.2 Base Step
+The backward induction begins at the maximum search depth, $d_{max}$.
+We aim to establish the induction property Cons($\gamma,f,d$) for random nodes $w$ at depth $d=d_{max}-\frac{1}{2}$.
+As our algorithm uses a learned value function, $V(z_{i})$, to evaluate leaf nodes $z_{i}$ (children of $w$), this step relies on the value function error bound established in Assumption B.3.
+Our goal is to bound the total error $|\hat{V}(w)-V^{*}_{p}(w)|$. Using the triangle inequality, we decompose this error:
+$|\hat{V}(w)-V^{*}_{p}(w)|\leq|\hat{V}(w)-\mathbb{E}[V(z_{i})]|+|\mathbb{E}[V(z_{i})]-V^{*}_{p}(w)|$
+We bound the two terms on the right-hand side separately.
+(a) Bounding the Sampling Error
+The first term, $|\hat{V}(w)-\mathbb{E}[V(z_{i})]|$, is the concentration error of the MCTS sampling process.
+Since the value function output $V(z_{i})$ is bounded (per Assumption B.3), we can apply Hoeffding’s inequality.
+Following the original PUCT proof (Auger et al., 2013), we set the deviation tolerance $t=n(w)^{-\frac{1}{3}}$, which gives us the following bound:
+$|\hat{V}(w)-\mathbb{E}[V(z_{i})]|\leq O(n(w)^{-\frac{1}{3}})\text{ e.s. in }n(w)$
+(b) Bounding the Approximation Error
+The second term, $|\mathbb{E}[V(z_{i})]-V^{*}_{p}(w)|$, represents the inherent bias from using an approximate value function. We bound it using Assumption B.3:
+$|\mathbb{E}[V(z_{i})]-V^{*}_{p}(w)|$ $=|\mathbb{E}[V(z_{i})]-\mathbb{E}[V^{*}_{p}(z_{i})]|$
+$\leq\mathbb{E}[|V(z_{i})-V^{*}_{p}(z_{i})|]$
+$\leq\mathbb{E}[\epsilon_{V}]=\epsilon_{V}$
+Conclusion for the Base Step
+By combining the bounds for the sampling error and the approximation error, we obtain the bound on the total error:
+$|\hat{V}(w)-V^{*}_{p}(w)|\leq O(n(w)^{-\frac{1}{3}})+\epsilon_{V}\text{ e.s. in }n(w)$
+This result precisely matches the form required by our revised induction property, Cons($\gamma,f,d$).
+We have successfully shown that the property holds for the base case at depth $d=d_{max}-\frac{1}{2}$, with parameters:
+  • $\gamma_{d_{max}-\frac{1}{2}}=\frac{1}{3}$
+  • $f_{d_{max}-\frac{1}{2}}(\epsilon_{V})=\epsilon_{V}$
+B.3 Inductive Step: From Decision to Random Nodes
+In this section, we perform the first inductive step, moving from decision nodes up to random nodes.
+We assume the induction property Cons($\gamma,f,d$) holds for all decision nodes $z$ at depth $d$.
+Our objective is to prove that, as a consequence, the property also holds for their parent nodes, which are the random nodes $w$ at depth $d-\frac{1}{2}$.
+A key distinction in our algorithm is the special handling of the very first action sampled from a decision node.
+This requires us to categorize all random nodes $w$ (at depth $d-\frac{1}{2}$) based on their history:
+  1. 1.
+    Category 1: First-Child Nodes ($w_{0}$): A random node that was the first child ever created from its parent decision node.
+Its initial value estimate was derived directly from our learned value function, $V$.
+  2. 2.
+    Category 2: Standard Nodes ($w_{j},j>0$): All other random nodes. Their value estimates have always been computed through the standard MCTS backup process.
+We will analyze these two categories separately to show that the induction property holds universally.
+B.3.1 Analysis of Standard Random Nodes ($w_{j},j>0$)
+For a standard random node $w$ in this category, the proof structure closely follows Section 5 of Auger et al. (2013). We aim to bound the total error $|\hat{V}(w)-V^{*}_{p}(w)|$.
+Let $i_{0}=\lfloor n(w)^{\alpha}\rfloor$ be the newest child of $w$ (a decision node at depth $d$), and let $r=i_{0}-1$ be the number of “mature” children.(i.e., all children except for the newest).
+Let $R_{w}(i)=r_{i}+\gamma\hat{V}(i)$ denote the return through child $i$.
+We decompose the total error into four terms:
 […]
-Appendix B Alternative Design Choices for Continuous BAMCP
-The terms labeled in blue in Algorithm 1 have alternative design choices. Empirical comparisons among these alternatives are reserved for future work.
+$+\underbrace{|\mathbb{E}[R_{w}(j)]-V^{*}_{p}(w)|}_{\text{(c) Bias / Approximation Error}}+\underbrace{|\frac{n(i_{0})}{n(w)}R_{w}(i_{0})|}_{\text{(d) ``Newest'' Child Error}}$
+We now bound each term:
+(a) Bounding the Weight Deviation Error
+This term is bounded by leveraging the fact that the children of a random node are explored in a nearly uniform manner (a consequence of the DPW mechanism, as shown in Lemma 1 of Auger et al.
+(2013).
+Following their derivation, this term is bounded by $O(n(w)^{-(1-\alpha)}+n(w)^{-\alpha})$.
+(b) Bounding the Sampling Error of “Mature” Returns
+This term represents the concentration error for the returns of the mature child nodes.
+The returns $R_{w}(i)$ are i.i.d.
+bounded random variables.
+Therefore, we can apply Hoeffding’s inequality.
+To align with the proof by Auger et al.
+(2013) and derive the resulting convergence rate for this layer, we select the deviation tolerance $t$ as:
+$t:=n(w)^{-\frac{\gamma_{d}}{1+3\gamma_{d}}}$
+With this choice, the application of Hoeffding’s inequality shows that the error for this term is bounded by $O(n(w)^{-\frac{\gamma_{d}}{1+3\gamma_{d}}})$ and that this bound holds exponentially surely.
+(c) Bounding the Bias / Approximation Error
+This term is unique to our proof and requires our induction hypothesis.
+$|\mathbb{E}[R_{w}(j)]-V^{*}_{p}(w)|$ $=|\mathbb{E}[r_{j}+\gamma\hat{V}(j)]-\mathbb{E}[r_{j}+\gamma V^{*}_{p}(j)]|$
+$=\gamma|\mathbb{E}[\hat{V}(j)-V^{*}_{p}(j)]|$
+$\leq\gamma\mathbb{E}[|\hat{V}(j)-V^{*}_{p}(j)|]$
+By the induction hypothesis for nodes at depth $d$, we have $|\hat{V}(j)-V^{*}_{p}(j)|\leq C_{d}n(j)^{-\gamma_{d}}+f_{d}(\epsilon_{V})$. Substituting this into the inequality gives:
+$|\mathbb{E}[R_{w}(j)]-V^{*}_{p}(w)|\leq\gamma\mathbb{E}[C_{d}n(j)^{-\gamma_{d}}]+\gamma f_{d}(\epsilon_{V})\text{ e.s. in }n(w)$
+The first part of this bound is a decaying term, while the second part shows how the bias propagates.
+(d) Bounding the “Newest” Child Error
+Following the same logic as Auger et al. (2013), this term is bounded by $O(n(w)^{-\alpha})$ because the newest child has been visited only a small number of times.
+Conclusion for Standard Nodes
+Combining these bounds, the overall convergence rate is determined by the slowest-converging term.
+Following the derivation in the PUCT paper, setting the exploration parameter $\alpha=\frac{3\gamma_{d}}{1+3\gamma_{d}}$, establishes the slowest convergence rate as $\frac{\gamma_{d}}{1+3\gamma_{d}}$.
+The bias term is propagated from the layer below.
+Thus, for any standard random node, the Cons($\gamma,f,d$) property holds.
+B.3.2 Analysis of the First-Child Random Node ($w_{0}$)
+For the first child $w_{0}$, the value estimate includes a special term from its initial visit.
+On the first visit, the algorithm samples a next state, which we denote $s_{0}^{\prime}$, and receives a corresponding reward $r_{0}$.
+The next state’s value is thus initially estimated using the learned function, $V(s_{0}^{\prime})$.
+The error analysis for this random node, $|\hat{V}(w_{0})-V^{*}_{p}(w_{0})|$, will contain the same four error terms as in the standard case (for the subsequent backups), plus an additional error term originating from the first visit:
+$\text{Additional Error}=\left|\frac{1}{n(w_{0})}(r_{0}+\gamma V(s^{\prime}_{0}))-\frac{1}{n(w_{0})}V^{*}_{p}(w_{0})\right|$
+Since $r_{0}$, $V(s^{\prime}_{0})$, and $V^{*}_{p}(w_{0})$ are all bounded constants, this additional error term is bounded by $O(n(w_{0})^{-1})$.
+The overall convergence rate for $w_{0}$ is determined by the minimum of the rates of all error components.
+From the PUCT derivation, we know that $\gamma_{d}\leq 1$, which implies $\gamma_{d-1/2}=\frac{\gamma_{d}}{1+3\gamma_{d}}<1$.
+Therefore, the convergence rate of the additional error term, $O(n(w_{0})^{-1})$, is faster than the bottleneck rate of the standard MCTS error terms, $O(n(w_{0})^{-\gamma_{d-1/2}})$.
+This special term does not change the overall convergence rate.
+Thus, the Cons($\gamma,f,d$) property also holds for the first-child random node $w_{0}$ with the exact same parameters.
+B.3.3 Conclusion for this Section
+We have shown that for any random node $w$ at depth $d-\frac{1}{2}$, if its children at depth $d$ satisfy the Cons($\gamma,f,d$) property, then $w$ itself also satisfies the property.
+The parameters of the property propagate upward according to the following recursive formulas:
+  • $\gamma_{d-\frac{1}{2}}=\frac{\gamma_{d}}{1+3\gamma_{d}}$
+  • $f_{d-\frac{1}{2}}(\epsilon_{V})=\gamma f_{d}(\epsilon_{V})$
+B.4 Inductive Step: From Random to Decision Nodes
+In this section, we perform the second inductive step.
+We assume the induction property Cons($\gamma,f,d$) holds for all random nodes $w$ at depth $d+\frac{1}{2}$.
+Our objective is to prove that the property also holds for their parent nodes, which are the decision nodes $z$ at depth $d$.
+The proof adapts the structure of Section 6 in Auger et al.
+(2013), carefully accounting for the bias term $f_{d+\frac{1}{2}}(\epsilon_{V})$ from our induction hypothesis.
+First we establish an upper bound on $\hat{V}(z)-V^{*}_{p}(z)$, and then a lower bound.
+B.4.1 Upper Bound on the Error
+To establish an upper bound on $\hat{V}(z)-V^{*}_{p}(z)$, we partition the children $w_{i}$ of $z$ into two classes, with a parameter $\epsilon<1-\alpha$ to be determined later:
+  • Class I: Children $w_{i}$ with few visits, $n(w_{i})\leq n(z)^{1-\alpha-\epsilon}$.
+  • Class II: All other children, for which $n(w_{i})>n(z)^{1-\alpha-\epsilon}$.
+For Class I, the number of children is at most $n(z)^{\alpha}$, so their total contribution is bounded by $\frac{n(z)^{\alpha}\cdot n(z)^{1-\alpha-\epsilon}}{n(z)}=n(z)^{-\epsilon}$.
+For Class II, we apply the induction hypothesis for nodes at depth $d+\frac{1}{2}$:
+$\sum_{i\in\text{Class II}}\frac{n(w_{i})}{n(z)}(\hat{V}(w_{i})-V^{*}_{p}(z))$ $\leq\sum_{i\in\text{Class II}}\frac{n(w_{i})}{n(z)}(\hat{V}(w_{i})-V^{*}_{p}(w_{i}))$
+$\leq\sum_{i\in\text{Class II}}\frac{n(w_{i})}{n(z)}(C_{d+\frac{1}{2}}n(w_{i})^{-\gamma_{d+\frac{1}{2}}}+f_{d+\frac{1}{2}}(\epsilon_{V}))$
+$\leq C_{d+\frac{1}{2}}\left(n(z)^{1-\alpha-\epsilon}\right)^{-\gamma_{d+\frac{1}{2}}}+f_{d+\frac{1}{2}}(\epsilon_{V})$
+Combining the two parts, we get:
+$\hat{V}(z)-V^{*}_{p}(z)\leq n(z)^{-\epsilon}+C_{d+\frac{1}{2}}n(z)^{-\gamma_{d+\frac{1}{2}}(1-\alpha-\epsilon)}+f_{d+\frac{1}{2}}(\epsilon_{V})$
+To optimize this bound, we balance the exponents by choosing $\epsilon=\frac{\gamma_{d+\frac{1}{2}}(1-\alpha)}{1+\gamma_{d+\frac{1}{2}}}$. This leads to the final upper bound for the error:
+$\hat{V}(z)-V^{*}_{p}(z)\leq(1+C_{d+\frac{1}{2}})n(z)^{-\frac{\gamma_{d+\frac{1}{2}}(1-\alpha)}{1+\gamma_{d+\frac{1}{2}}}}+f_{d+\frac{1}{2}}(\epsilon_{V})$
+B.4.2 Lower Bound on the Error
+The proof for the lower bound is more involved.
+It relies on the parameter definitions from the PUCT paper, as well as two key results regarding the UCT exploration mechanism, which we state here for completeness before proceeding.
+Lemma B.5 (Adapted from Lemma 3, Auger et al. (2013)).
+Consider a bandit setting where the score of a child node $w_{i}$ at time $n$ is computed by the UCT rule:
+$sc_{n}(w_{i})=\hat{V}_{n}(w_{i})+\sqrt{\frac{g(n)}{n(w_{i})}}$
+where $g(n)$ is a non-decreasing exploration function. If $i$ denotes the $i^{th}$ constucted child, for all $n>i^{\frac{1}{\alpha(1-\alpha)}}$ we have
+$n(w_{i})\geq\frac{1}{4}\min\left(g(n^{1-\alpha}),n^{1-\alpha}\right).$
+Corollary B.6 (Adapted from Corollary 3, Auger et al. (2013)).
+For the exploration function $g(n)=n^{e}$ with $0<e<1$, we obtain
+$n(w_{i})\geq\frac{1}{4}n^{e(1-\alpha)}\text{ if }i\leq n^{\alpha(1-\alpha)}.$ (8)
+With these results, we can define the parameters as functions of $\gamma_{d+\frac{1}{2}}$ and $p$ (from Assumption B.2) for this layer and proceed with the three-step proof structure.
+  • Progressive widening coefficient: $\alpha:=\frac{\gamma_{d+\frac{1}{2}}}{1+4\gamma_{d+\frac{1}{2}}}$
+  • Exploration coefficient: $e:=\frac{1}{2p\left(1+4\gamma_{d+\frac{1}{2}}\right)}$
+  • Auxiliary proof parameters: $\xi:=\frac{1}{1+e\gamma_{d+\frac{1}{2}}(1-\alpha)}$ and $\Delta:=\left(\frac{1}{4}n(z)^{\xi e(1-\alpha)}\right)^{-\gamma_{d+\frac{1}{2}}}$
+Step 1: Existence of a near-optimal child
+Based on Assumption B.2 (Policy Regularity), a direct adaptation of the proof in Auger et al.
+(2013) shows that a near-optimal child node is discovered with high probability.
+Formally: Exponentially surely in $n(z)$, there exists at time $\lceil n(z)^{\xi(1-\alpha)}\rceil$ a child $w_{0}$ of $z$ such that
+$V^{*}_{p}(w_{0})\geq V^{*}_{p}(z)-\Delta\text{ and }i_{0}\leq n(z)^{\xi(1-\alpha)\alpha}$ (9)
+Step 2: Score of children selected at later stages
+The objective of this step is to show that any child selected after time $n(z)^{\xi}$ must have a high score. Let $n^{\prime}$ be such that $n^{\xi}\leq n^{\prime}\leq n$. And, by Corollary B.6,
+$n^{\prime}(w_{0})\geq\frac{1}{4}n^{\prime e(1-\alpha)}\geq\frac{1}{4}n^{\xi e(1-\alpha)}.$
+First, according to Corollary B.6, the near-optimal child $w_{0}$ discovered in Step 1 will be visited a number of times $n(w_{0})$ that grows polynomially with $n(z)$.
+We can therefore apply our induction hypothesis Cons($\gamma,f,d$) to $w_{0}$, with the Eq.(9), there exists a $C^{\prime}>0$ that:
+$\hat{V}(w_{0})$ $\geq V^{*}_{p}(w_{0})-C^{\prime}\left(\frac{1}{4}n^{\xi e(1-\alpha)}\right)^{-\gamma_{d+\frac{1}{2}}}-f_{d+\frac{1}{2}}(\epsilon_{V})$
+$\geq V^{*}_{p}(z)-(1+C^{\prime})\Delta-f_{d+\frac{1}{2}}(\epsilon_{V})$
+Now, consider any child $w_{1}$ selected by the UCT rule at a later time $n^{\prime}\geq n(z)^{\xi}$. Its score must be at least as high as the score of $w_{0}$. This gives the key inequality:
+$\hat{V}(w_{1})+\sqrt{\frac{n^{\prime e}}{n^{\prime}(w_{1})}}\geq\hat{V}(w_{0})+\sqrt{\frac{n^{\prime e}}{n^{\prime}(w_{0})}}$
+Combining these results, we can establish a lower bound on the estimated value $\hat{V}(w_{1})$ of any child selected at a late stage of the search.
+$\hat{V}(w_{1})+\sqrt{\frac{n}{n^{\prime}(w_{1})}}\geq V^{*}_{p}(z)-(1+C^{\prime})\Delta-f_{d+\frac{1}{2}}(\epsilon_{V})$
+As shown by Auger et al. (2013), this property holds exponentially surely in $n(z)$.
+Step 3: Bounding the value estimate $\hat{V}(z)$
+Consider a child $w_{1}$ selected after $n(z)^{\xi}$. By the previous step, exponentially in $n(z)$, this child must either satisfy
+$\sqrt{\frac{n^{e}}{n(w_{1})}}\geq\Delta$ (10)
+$\text{or }\hat{V}(w_{1})\geq V(z)-(2+C^{\prime})\Delta-f_{d+\frac{1}{2}}(\epsilon_{V})$ (11)
+We now bound the average value $\hat{V}(z)$ by partitioning the children of $z$ into three categories:
+  1. 1.
+    Children visited only before time $n(z)^{\xi}$.
+  2. 2.
+    Children visited after $n(z)^{\xi}$ satisfying (10)
+  3. 3.
+    Children visited after $n(z)^{\xi}$ satisfying (11).
+We bound the contribution of each category to the total error sum. The contribution from Category 1 is bounded by $O(n(z)^{\xi-1})$, since we have
+$\left|\sum_{w\text{ in cat.1}}\frac{n(w)}{n(z)}\left(\hat{V}(w)-V^{*}_{p}(z)\right)\right|\leq\frac{\sum_{w\text{ in cat.1}}n(w)}{n(z)}\leq\frac{n(z)^{\xi}}{n(z)}$
+For Category 2, the definition of the category implies $n(w_{i})\leq n(z)^{e}/\Delta^{2}$.
+Since there are at most $n(z)^{\alpha}$ children, their total contribution to the error sum is bounded by $O(n(z)^{\alpha+e-1}/\Delta^{2})$.
+For Category 3, the bound on $\hat{V}(w_{i})$ from Step 2 introduces the crucial $f_{d+1/2}(\epsilon_{V})$ term.
+Combining the bounds from all three categories gives the final lower bound on the error. Following the argument in Auger et al. (2013):
+$\hat{V}(z)-V^{*}_{p}(z)\geq\underbrace{-(2+C^{\prime})\Delta(1-n(z)^{\xi-1})-f_{d+\frac{1}{2}}(\epsilon_{V})}_{\text{cat.3}}-\underbrace{n(z)^{\xi-1}}_{\text{cat.1}}-\underbrace{\frac{n(z)^{\alpha+e-1}}{\Delta^{2}}}_{\text{cat.2}}$
+By comparing the magnitudes of the decaying terms with the chosen parameter definitions, as is done in Auger et al. (2013), this complex expression simplifies to the desired form.
+B.4.3 Conclusion for this Section
+By combining the results from the upper and lower bound analyses, we conclude that the total error at a decision node $z$ is bounded as:
+$|\hat{V}(z)-V^{*}_{p}(z)|\leq C_{d}n(z)^{-\gamma_{d}}+f_{d+\frac{1}{2}}(\epsilon_{V})$
+This demonstrates that the induction property Cons($\gamma,f,d$) holds for depth $d$.
+The parameters propagate upward according to the following recursive formulas, which are identical to those derived by Auger et al.
+(2013):
+  • $C_{d}=(5+C^{\prime})4^{\gamma_{d+\frac{1}{2}}}$
+  • $\gamma_{d}=\frac{\gamma_{d+\frac{1}{2}}}{1+7\gamma_{d+\frac{1}{2}}}$
+  • $f_{d}(\epsilon_{V})=f_{d+\frac{1}{2}}(\epsilon_{V})$
+B.5 Conclusion of the proof
+In the preceding sections, we have established all the necessary components for our backward induction argument.
+  • In Section B.2, we established the base case, proving that the induction property Cons($\gamma,f,d$) holds for random nodes at depth $d=d_{max}-\frac{1}{2}$, with initial parameters $\gamma_{d_{max}-\frac{1}{2}}=\frac{1}{3}$ and $f_{d_{max}-\frac{1}{2}}(\epsilon_{V})=\epsilon_{V}$.
+  • In Section B.3, we proved the inductive step from decision nodes up to random nodes, deriving the recursions $\gamma_{d-\frac{1}{2}}=\frac{\gamma_{d}}{1+3\gamma_{d}}$ and $f_{d-\frac{1}{2}}(\epsilon_{V})=\gamma f_{d}(\epsilon_{V})$.
+  • In Section B.4, we proved the inductive step from random nodes up to decision nodes, deriving the recursions $\gamma_{d}=\frac{\gamma_{d+\frac{1}{2}}}{1+7\gamma_{d+\frac{1}{2}}}$ and $f_{d}(\epsilon_{V})=f_{d+\frac{1}{2}}(\epsilon_{V})$.
+With the base case proven and both inductive steps validated, we conclude by backward induction that the property Cons($\gamma,f,d$) holds for all depths up to the root of the search tree ($d=0$).
+By solving the recursion for the bias term, we can find the accumulated error at the root.
+The base error at depth $d_{max}-\frac{1}{2}$ is $f_{d_{max}-\frac{1}{2}}(\epsilon_{V})=\epsilon_{V}$.
+The error term is multiplied by a discount factor $\gamma$ each time it propagates from a decision node layer up to a random node layer.
+This occurs $d_{max}$ times from the leaves to the root.
+Therefore, the bias term at the root node ($d=0$) is $f_{0}(\epsilon_{V})=\gamma^{d_{max}}\epsilon_{V}$.
+Specifically, for the root node $z_{0}$, its total error is bounded by:
+$|\hat{V}(z_{0})-V^{*}_{p}(z_{0})|\leq C_{0}n(z_{0})^{-\gamma_{0}}+\gamma^{d_{max}}\epsilon_{V}$
+where $\gamma_{0}>0$ and $C_{0}$ are constants derived from the recursions.
+This shows that as the number of simulations $E=n(z_{0})\rightarrow\infty$, the estimated value of the root node converges to a ball of radius $\gamma^{d_{max}}\epsilon_{V}$ around the true optimal value of the P-MDP.
+Notably, the size of this error ball, which stems from our function approximator, shrinks exponentially with the search depth $d_{max}$.
+This completes the proof of Theorem 4.1, establishing that Continuous BAMCP is a consistent planner for the P-MDP. ∎
+Appendix C Alternative Design Choices for Continuous BAMCP
+The terms labeled in blue in Algorithm 1 have alternative design choices, which are introduced as below. Empirical comparisons among these alternatives are reserved for future work.
 $\tilde{Q}((s,h),x)$ in the exploration strategy, i.e., $a\leftarrow\arg\max_{x\in C((s,h))}\tilde{Q}((s,h),x)$, could take various forms.
-For instance, in (Couëtoux et al.
+For instance, in Couëtoux et al.
 (2011); Guez et al.
 (2013); Lee et al.
-(2020)), $\tilde{Q}((s,h),x)=Q((s,h),x)+c\sqrt{\frac{\log N((s,h))}{N((s,h),x)}}$; in PUCT (Auger et al.
-(2013)), $\tilde{Q}((s,h),x)=Q((s,h),x)+\sqrt{\frac{N((s,h))^{e(d)}}{N((s,h),x)}}$, where $e(d)$ is a schedule of coefficients related to the search depth $d$; in Sampled MuZero (a variant of MuZero that can be applied in continuous action spaces (Hubert et al.
-(2021))), $\tilde{Q}((s,h),x)=Q((s,h),x)+\hat{\pi}(x|(s,h))\frac{\sqrt{N((s,h))}}{1+N((s,h),x)}\left(c_{1}+\log\left(\frac{N((s,h))+c_{2}+1}{c_{2}}\right)\right)$.
+(2020), $\tilde{Q}((s,h),x)=Q((s,h),x)+c\sqrt{\frac{\log N((s,h))}{N((s,h),x)}}$; in PUCT (Auger et al., 2013), $\tilde{Q}((s,h),x)=Q((s,h),x)+\sqrt{\frac{N((s,h))^{e(d)}}{N((s,h),x)}}$, where $e(d)$ is a schedule of coefficients related to the search depth $d$; in Sampled MuZero (a variant of MuZero that can be applied in continuous action spaces (Hubert et al., 2021)), $\tilde{Q}((s,h),x)=Q((s,h),x)+\hat{\pi}(x|(s,h))\frac{\sqrt{N((s,h))}}{1+N((s,h),x)}\left(c_{1}+\log\left(\frac{N((s,h))+c_{2}+1}{c_{2}}\right)\right)$.
 Here, $c$, $c_{1}$, $c_{2}$ are hyperparameters, $\hat{\pi}=\hat{\beta}\pi^{1-1/\tau}$ is a sample policy defined upon the real policy $\pi$.
@@
 Following BAMCP, we adopt the first definition of $\tilde{Q}((s,h),x)$, though it could potentially be improved with other choices.
-In addition, as an implementation trick (Hamrick et al.
-(2021)), the Q estimates are usually normalized into $\bar{Q}\in[0,1]$ before being used to calculate $\tilde{Q}$ as above.
+In addition, as an implementation trick (Hamrick et al., 2021), the Q estimates are usually normalized into $\bar{Q}\in[0,1]$ before being used to calculate $\tilde{Q}$ as above.
 The normalized estimates can be computed as $\bar{Q}((s,h),x)=\frac{Q((s,h),x)-Q_{\text{min}}}{Q_{\text{max}}-Q_{\text{min}}}$, where $Q_{\text{max}}$ and $Q_{\text{min}}$ are the maximum and minimum Q values observed in the search tree so far.
 As the planning/search result, $\pi_{\text{ret}}$ can take multiple forms.
-In (Guez et al.
-(2013); Sunberg & Kochenderfer (2018); Lee et al.
-(2020)), $\pi_{\text{ret}}((s,h))=\arg\max_{a\in C((s,h))}Q((s,h),a)$; in (Sampled) MuZero, $\pi_{\text{ret}}(a|(s,h))=\frac{N((s,h),a)^{1/\tau}}{\sum_{x\in C((s,h))}N((s,h),x)^{1/\tau}}$; in ROSMO (a variant of MuZero with improved performance in offline scenarios (Liu et al.
-(2023))), $\pi_{\text{ret}}(a|(s,h))\propto\pi(a|(s,h))\exp(Q((s,h),a)-V((s,h)))$.
+In Guez et al.
+(2013); Sunberg and Kochenderfer (2018); Lee et al.
+(2020), $\pi_{\text{ret}}((s,h))=\arg\max_{a\in C((s,h))}Q((s,h),a)$; in (Sampled) MuZero, $\pi_{\text{ret}}(a|(s,h))=\frac{N((s,h),a)^{1/\tau}}{\sum_{x\in C((s,h))}N((s,h),x)^{1/\tau}}$; in ROSMO (a variant of MuZero with improved performance in offline scenarios (Liu et al., 2023)), $\pi_{\text{ret}}(a|(s,h))\propto\pi(a|(s,h))\exp(Q((s,h),a)-V((s,h)))$.
 Here, $\tau\in(0,1]$ is a temperature parameter and decays with the training process, ensuring the action selection becomes greedier.
@@
 This is because (1) as described in PUCT, the returned action should be the most visited one, which is not necessarily the one with the highest Q value, and (2) ROSMO adopts one-step look-ahead rather than deep tree search at each root node, which does not align with our approach.
-As for the conditions of double progressive widening, PUCT designs $\alpha$ and $\beta$ to be functions of the search depth $d$, while UCT-DPW (Couëtoux et al.
-(2011)) utilizes a different set of conditions: $\left\lceil{K_{a}N((s,h))^{\alpha}}\right\rceil\geq|C((s,h))|$, $\left\lceil{K_{s}N((s,h),a)^{\beta}}\right\rceil\geq|C((s,h),a)|$, where $K_{a},K_{s},\alpha,\beta$ are all constant hyperparameters.
-When the progressive widening condition for sampling the next state is not satisfied, either the least visited node in $C((s,h),a)$ can be selected (following PUCT), or a random node can be sampled from $C((s,h),a)$ following a distribution proportional to the number of visits (following UCT-DPW).
+As for the conditions of double progressive widening, PUCT designs $\alpha$ and $\beta$ to be functions of the search depth $d$, while UCT-DPW (Couëtoux et al., 2011) utilizes a different set of conditions: $\left\lceil{K_{a}N((s,h))^{\alpha}}\right\rceil\geq|C((s,h))|$, $\left\lceil{K_{s}N((s,h),a)^{\beta}}\right\rceil\geq|C((s,h),a)|$, where $K_{a},K_{s},\alpha,\beta$ are all constant hyperparameters.
+Moreover, when the progressive widening condition for sampling the next state is not satisfied, either the least visited node in $C((s,h),a)$ can be selected (following PUCT), or a random node can be sampled from $C((s,h),a)$ according to a distribution proportional to the number of visits (following UCT-DPW).
 As shown in Algorithm 1, we follow the designs of PUCT, but keep $\alpha$ and $\beta$ as constants for simplicity in hyperparameter fine-tuning.
@@
 At the end of the simulation procedure, we can either apply rollouts, i.e., simulating a single path until the end of an episode, to estimate the expected value for a leaf node $(s,h)$, or directly use $V((s,h))$ as the estimation.
-The former approach is widely used in online planning algorithms (Guez et al.
-(2013); Sunberg & Kochenderfer (2018); Lee et al.
-(2020)), while the latter is used in iterative frameworks like MuZero.
-Appendix C Key Hyperparameter Setup
+The former approach is widely used in online planning algorithms (Guez et al., 2013; Sunberg and Kochenderfer, 2018; Lee et al., 2020), while the latter is used in policy iteration frameworks like ours.
+Appendix D Key Hyperparameter Setup
 Data Type Environment BA-MBRL BA-MCTS BA-MCTS-SL
-$K$ $\lambda$ $H$ $N$ $K$ $\lambda$ $H$ $N$ $K$ $\lambda$ $H$ $N$
+$K$ $\lambda$ $H$ NN $K$ $\lambda$ $H$ NN $K$ $\lambda$ $H$ NN
 random HalfCheetah 10 7 6 200 10 7 6 800 10 7 6 500
@@
 med-expert Walker2d 6 20 37 400 6 20 37 800 6 20 37 500
-Table 3: Key hyperparameters of the proposed algorithms for each evaluation task. $K$: ensemble size, $\lambda$: reward penalty coefficient, $H$: rollout horizon, $N$: number of training epochs.
-In Table 3, we list the key hyperparameters of the proposed algorithms.
+Table 5: Key hyperparameters of the proposed algorithms for each evaluation task in the ablation study.
+$K$: ensemble size, $\lambda$: reward penalty coefficient, $H$: rollout horizon, NN: number of training epochs.
+Table 5 lists the key hyperparameters of BA-MCTS and its ablated variants used in the ablation study.
 For each task, an ensemble of $K$ dynamics and reward models is trained using the provided offline dataset.
@@
 To mitigate overestimation, a reward penalty based on the discrepancy among the ensemble members is applied with a coefficient $\lambda$, as shown in Equation (5).
-The setups for $K$, $\lambda$, and $H$ are almost the same across the three algorithms and primarily inherited from the baseline – “Optimized" (Lu et al.
-(2022)), to make sure the improvements are brought by the Bayesian RL and deep search components.
+The setups for $K$, $\lambda$, and $H$ are almost the same across the three algorithms and primarily inherited from the baseline – “Optimized” (Lu et al., 2022), to make sure the improvements are brought by the Bayesian RL and deep search components (rather than tricky hyperparameter setups).
 The policy is evaluated on the ground truth environment for 10 episodes at the end of each training epoch.
-We report the average scores across the final 10 training epochs of our algorithms in Tables 1 and 5.
+For all tables involving benchmarking results on D4RL, we report the average scores across the final 10 training epochs of our algorithms.
 It is important to note that increasing the number of training epochs $N$ does not necessarily lead to better policy performance, since the training is based on learned dynamics and reward models rather than the ground truth.
-According to (Lu et al.
-(2022)) and our experiments, the hyperparameters listed above can significantly influence the performance of model-based RL.
+According to Lu et al.
+(2022) and our experiments, the hyperparameters listed above can significantly influence the performance of model-based RL.
 Adjusting these hyperparameters could either enhance or impair the learning performance of our algorithms.
-We also suspect that the performance of the baselines listed in Tables 1 and 5, which are from their original papers, could be further improved by fine-tuning the relevant hyperparameters.
+We also suspect that the performance of the baselines listed in Table 9, which are from their original papers, could be further improved by fine-tuning the relevant hyperparameters.
 Data Type Environment BA-MCTS BA-MCTS-SL
@@
 Table 6: Important hyperparameters used in the search process.
-(a) hc-med-expert
-(b) hc-med-replay
-(c) hc-medium
-(d) hc-random
-(e) hp-med-expert
-(f) hp-med-replay
-(g) hp-medium
-(h) hp-random
-(i) wk-med-expert
-(j) wk-med-replay
-(k) wk-medium
-(l) wk-random
-Figure 3: Performance of our proposed algorithms on D4RL MuJoCo tasks.
-The results for HalfCheetah, Hopper, and Walker2d are presented in the three rows, respectively.
-Each subfigure depicts the change in the undiscounted episodic return as a function of training epochs.
-Experiments are repeated three times with different random seeds, with the solid line representing the mean and the shaded area indicating the 95% confidence interval.
-For reference, the expert-level episodic returns for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively.
-Note that the training epochs for each algorithm, as listed in Table 3, have been linearly scaled to 800 for better visualization.
-BA-MCTS and BA-MCTS-SL utilize Bayes Adaptive Monte Carlo Tree Search to collect samples for offline model-based RL.
-Instead of performing a tree search at every state, we randomly select a proportion (i.e., $\rho$) of states from the available 50000 states at each rollout time step as root nodes for tree search.
+BA-MCTS and BA-MCTS-SL utilize Continuous BAMCP (i.e., Algorithm 1) to collect samples for subsequent policy distillation.
+Instead of performing a tree search at every state, we randomly select a proportion (i.e., $\rho$) of states from the available 50000 states at each rollout time step as root nodes for tree search to decide on the optimal actions.
 For the remaining states, actions are sampled directly from the policy, i.e., $a\sim\pi(\cdot|s)$.
@@
 Also, in the Walker2d environment, BA-MCTS-SL requires a warm-up training phase of $N_{P}$ epochs using BA-MBRL, allowing the initial policy to generate effective signals for supervised learning.
-Appendix D Comparisons with Model-free Methods on D4RL MuJoCo
-Data Type Environment BA-MCTS -SL (ours) BA-MCTS (ours) BA-MBRL (ours) CQL BEAR BRAC-v SAC BC
-random HalfCheetah 29.20 $\pm$ 2.00 36.23 $\pm$ 1.04 32.76 $\pm$ 1.16 35.4 25.1 31.2 30.5 2.1
-random Hopper 33.83 $\pm$ 0.10 31.56 $\pm$ 0.12 31.47 $\pm$ 0.03 10.8 11.4 12.2 11.3 1.6
-random Walker2d 21.89 $\pm$ 0.07 21.59 $\pm$ 0.32 21.45 $\pm$ 0.53 7.0 7.3 1.9 4.1 9.8
-medium HalfCheetah 70.47 $\pm$ 3.52 75.84 $\pm$ 3.81 56.54 $\pm$ 5.20 44.4 41.7 46.3 -4.3 36.1
-medium Hopper 97.75 $\pm$ 7.09 96.70 $\pm$ 14.0 98.25 $\pm$ 3.42 86.6 52.1 31.1 0.8 29.0
-medium Walker2d 82.24 $\pm$ 1.85 74.73 $\pm$ 3.25 75.41 $\pm$ 4.17 74.5 59.1 81.1 0.9 6.6
-med-replay HalfCheetah 61.16 $\pm$ 1.60 65.45 $\pm$ 0.81 62.50 $\pm$ 0.18 46.2 38.6 47.7 -2.4 38.4
-med-replay Hopper 106.3 $\pm$ 0.13 101.8 $\pm$ 3.46 93.91 $\pm$ 4.25 48.6 33.7 0.6 3.5 11.8
-med-replay Walker2d 92.13 $\pm$ 5.13 95.06 $\pm$ 2.11 97.54$\pm$ 1.93 32.6 19.2 0.9 1.9 11.3
-med-expert HalfCheetah 80.53 $\pm$ 6.63 76.16 $\pm$ 10.3 90.52 $\pm$ 4.13 62.4 53.4 41.9 1.8 35.8
-med-expert Hopper 112.2 $\pm$ 0.29 108.3 $\pm$ 0.22 107.8 $\pm$ 0.37 111 96.3 0.8 1.6 111.9
-med-expert Walker2d 107.7 $\pm$ 0.82 110.0 $\pm$ 1.74 84.71 $\pm$ 0.87 98.7 40.1 81.6 -0.1 6.4
-Average Score 74.62 74.45 71.06 54.85 39.83 31.44 4.13 25.07
-Table 5: Comparisons between the proposed algorithms and model-free offline policy learning methods on the D4RL benchmark suite.
-Each value represents the normalized score, as proposed in (Fu et al.
-(2020)), of the policy trained by the corresponding algorithm.
-These scores are undiscounted returns normalized to approximately range between 0 and 100, where a score of 0 corresponds to a random policy and a score of 100 corresponds to an expert-level policy.
-For our algorithms, we report the average score of the final ten policy learning epochs and its standard deviation across three different random seeds.
-As a complement to Table 1, we compare our algorithms with a series of model-free offline policy learning (Chen et al.
-(2024)) methods.
-We include SOTA model-free offline RL methods: CQL (Kumar et al.
-(2020)), BEAR (Kumar et al.
-(2019)), and BRAC-v (Wu et al.
-(2019)).
-Additionally, we show the performance of directly applying SAC or Behavioral Cloning (BC, Chen et al.
-(2024)) to the provided offline dataset in the last two columns.
-The mean performance of the baselines are taken from related works (Yu et al.
-(2020); Kidambi et al.
-(2020); Fu et al.
-(2020)).
-Our algorithms show significantly better performance, demonstrating the necessity of model-based learning in these environments.
-The training plots of our proposed algorithms in each environment is further detailed in Figure 3.
-Appendix E Computation Cost of Sampled EfficientZero on D4RL MuJoCo
-Here, we report the training time of Sampled EfficientZero on the D4RL MuJoCo tasks in Table 6.
-This set of experiments is run on a server with 40 Intel(R) Xeon(R) Gold 5215 CPUs and 4 Tesla V100-SXM2-32GB GPUs.
-Despite the substantial computational input, no corresponding performance improvement is observed.
-We welcome researchers to use our released code for further hyperparameter fine-tuning and to improve the results we have provided.
-hc-med-expert hc-med-replay hc-medium hc-random
-55.4 $\pm$ 1.3 54.8 $\pm$ 1.8 54.7 $\pm$ 1.9 54.8 $\pm$ 2.6
-hp-med-expert hp-med-replay hp-medium hp-random
-66.1 $\pm$ 0.9 126.6 $\pm$ 12.0 70.5 $\pm$ 1.0 153.7 $\pm$ 19.4
-wk-med-expert wk-med-replay wk-medium wk-random
-60.8 $\pm$ 1.8 134.9 $\pm$ 4.2 63.7 $\pm$ 1.5 175.7 $\pm$ 28.2
-Table 6: Training time (in hours) of Sampled EfficientZero for each evaluation task. The mean and standard deviation from repeated experiments are reported.
-Appendix F Ablation Study on the Reward Penalty
-Data Type Environment BA-MCTS -SL BA-MCTS BA-MBRL BA-MCTS -SL ($\lambda=0$) BA-MCTS ($\lambda=0$) BA-MBRL ($\lambda=0$)
-random HalfCheetah 29.20 $\pm$ 2.00 36.23 $\pm$ 1.04 32.76 $\pm$ 1.16 34.80 $\pm$ 1.39 38.78 $\pm$ 1.65 39.64 $\pm$ 2.86
-random Hopper 33.83 $\pm$ 0.10 31.56 $\pm$ 0.12 31.47 $\pm$ 0.03 9.16 $\pm$ 0.16 7.44 $\pm$ 0.14 6.97 $\pm$ 0.07
-random Walker2d 21.89 $\pm$ 0.07 21.59 $\pm$ 0.32 21.45 $\pm$ 0.53 17.53 $\pm$ 6.16 21.53 $\pm$ 0.42 21.41 $\pm$ 0.64
-medium HalfCheetah 70.47 $\pm$ 3.52 75.84 $\pm$ 3.81 56.54 $\pm$ 5.20 61.64 $\pm$ 4.58 60.84 $\pm$ 2.00 41.49 $\pm$ 2.29
-medium Hopper 97.75 $\pm$ 7.09 96.70 $\pm$ 14.0 98.25 $\pm$ 3.42 102.8 $\pm$ 2.29 104.4 $\pm$ 1.88 93.68 $\pm$ 11.4
-medium Walker2d 82.24 $\pm$ 1.85 74.73 $\pm$ 3.25 75.41 $\pm$ 4.17 82.61 $\pm$ 0.86 57.01 $\pm$ 7.24 57.97 $\pm$ 15.4
-med-replay HalfCheetah 61.16 $\pm$ 1.60 65.45 $\pm$ 0.81 62.50 $\pm$ 0.18 42.10 $\pm$ 2.85 36.65 $\pm$ 2.39 44.03 $\pm$ 7.35
-med-replay Hopper 106.3 $\pm$ 0.13 101.8 $\pm$ 3.46 93.91 $\pm$ 4.25 107.9 $\pm$ 0.07 84.11 $\pm$ 2.97 91.81 $\pm$ 11.5
-med-replay Walker2d 92.13 $\pm$ 5.13 95.06 $\pm$ 2.11 97.54$\pm$ 1.93 88.61 $\pm$ 5.21 97.33 $\pm$ 3.51 98.19 $\pm$ 1.23
-med-expert HalfCheetah 80.53 $\pm$ 6.63 76.16 $\pm$ 10.3 90.52 $\pm$ 4.13 51.76 $\pm$ 5.31 26.60 $\pm$ 1.46 29.88 $\pm$ 2.28
-med-expert Hopper 112.2 $\pm$ 0.29 108.3 $\pm$ 0.22 107.8 $\pm$ 0.37 106.8 $\pm$ 6.34 81.76 $\pm$ 6.45 86.79 $\pm$ 18.7
-med-expert Walker2d 107.7 $\pm$ 0.82 110.0 $\pm$ 1.74 84.71 $\pm$ 0.87 110.8 $\pm$ 1.72 110.2 $\pm$ 0.91 53.35 $\pm$ 38.0
-Average Score 74.62 74.45 71.06 68.04 60.55 55.43
-Table 7: Comparison of the proposed algorithms with their corresponding versions without the reward penalty (i.e., $\lambda=0$).
-The definitions of the values in this table are consistent with those in Table 5.
-To demonstrate the necessity of incorporating the reward penalty in offline MBRL, we conduct an ablation study by setting $\lambda$ in Eq.
-(5) to 0, resulting in ablated versions of our proposed three algorithms.
-The results are presented in Table 7.
-First, the average performance of the algorithms with the reward penalty is consistently better, demonstrating the importance of using reward penalties in offline MBRL to prevent the overexploitation of the learned world models (which can be inaccurate).
-Second, the supervised-learning-based algorithm (i.e., BA-MCTS-SL ($\lambda=0$)) is less affected by the absence of the reward penalty, compared to the policy-gradient-based methods.
-Notably, BA-MCTS-SL ($\lambda=0$) and BA-MCTS-SL achieve comparable performance in the Hopper and Walker2d tasks.
-This shows an additional advantage of BA-MCTS-SL – its reduced sensitivity to model inaccuracies.
-Lastly, there are instances where superior performance is achieved with $\lambda$ set to 0.
-For example, BA-MCTS-SL ($\lambda=0$) performs better than BA-MCTS-SL in 5 out of 12 tasks.
-This suggests that the performance of our algorithms in Tables 1 and 5 could be further improved by adjusting hyperparameters such as $\lambda$ [^12] As mentioned in Section 5, we retain most of the hyperparameter settings from Optimized (Lu et al.
-(2022)) to ensure that the performance improvements are attributed to our algorithm design..
-Appendix G Details of the Tokamak Control Tasks
+Appendix E Details of the Tokamak Control Tasks
 STATE SPACE
@@
 The tokamak is one of the most promising confinement devices.
-It uses magnetic fields acting on hydrogen atoms that have been ionized (given a charge) so that the magnetic fields can exert a force on the moving particles (Pironti & Walker (2005)).
+It uses magnetic fields acting on hydrogen atoms that have been ionized (given a charge) so that the magnetic fields can exert a force on the moving particles (Pironti and Walker, 2005).
 Char et al.
@@
 The authors demonstrate that the learned model predicts these measurements for entire shots with remarkable accuracy.
-Thus, we use this model as a “ground truth" simulator for tokamak control tasks.
+Thus, we use this model as a “ground truth” simulator for tokamak control tasks.
 Specifically, we generate a dataset of 725270 transitions for offline RL and evaluate the learned policy using this data-driven simulator.
 The state and action spaces for the tokamak control tasks are outlined in Table 7.
-For detailed physical explanations of their components, please refer to (Abbate et al.
+For detailed physical explanations of their components, please refer to Abbate et al.
 (2021); Char et al.
 (2023); Ariola et al.
-(2008)).
+(2008).
 The state space consists of five scalar values and six profiles which are discretized measurements of physical quantities along the minor radius of the toroid.
-After applying principal component analysis (Maćkiewicz & Ratajczak (1993)), the pressure profile is reduced to two dimensions, while the other profiles are reduced to four dimensions each.
+After applying principal component analysis (Maćkiewicz and Ratajczak, 1993), the pressure profile is reduced to two dimensions, while the other profiles are reduced to four dimensions each.
 In total, the state space comprises 27 dimensions.
@@
 The reward function for each task is defined as the negative squared tracking error of the corresponding component (i.e., temperature, rotation, or $\beta_{n}$) at each time step, and the reward is normalized by the episode horizon (i.e., 251 time steps).
-Notably, for policy learning, the reward function is provided rather than learned from the offline dataset as in D4RL tasks; and the dataset does not include the reference shot or any nearby, similar shots.
+Notably, for policy learning, the reward function is provided rather than learned from the offline dataset as in D4RL tasks; and the dataset (for offline RL) does not include the reference shot or any nearby, similar shots, making pure imitation infeasible.
+Appendix F Additional Evaluation Results
+This appendix provides detailed empirical results and analyses to supplement the findings presented in Section 5.
+Appendix F.1 offers both quantitative and qualitative analyses of the benefits of belief adaptation, presenting additional results for Section 5.1.
+The following appendices report additional experimental results for benchmarking on MuJoCo D4RL tasks (i.e., Section 5.2):
+  • Appendix F.2: Evaluates our algorithm against a suite of model-free offline RL methods.
+  • Appendix F.3: Presents a comprehensive comparison against 19 offline RL baselines, including a statistical significance analysis.
+  • Appendix F.4: Provides a detailed comparison with MuZero and its variants, focusing on their algorithm design, empirical performance, and computational cost.
+Finally, we provide complete results for the ablation study (i.e., Section 5.3):
+  • Appendix F.5: Demonstrates the significant performance enhancement achieved by the belief adaptation module and our MCTS-based planner.
+  • Appendix F.6: Discusses the trade-offs between the two policy distillation mechanisms investigated (i.e., SAC and supervised learning).
+  • Appendix F.7: Presents an ablation study on the necessity of using reward penalty in offline MBRL.
+F.1 Quantitative and Qualitative Analysis of the Effectiveness of Belief Adaptation
+To understand why belief adaptation over an ensemble of models using a Bayes-Adaptive MDP (BAMDP) is effective, we analyze its impact on the prediction accuracy of the learned world models.
+The Bayesian adaptation, as defined in Eq.
+(4), uses observed state transition sequences to adjust the belief over each ensemble member, thereby enhancing the quality of predicted rollout trajectories.
+For each benchmarking task in D4RL MuJoCo, we compute the average transition likelihood (i.e., $\mathcal{P}(s^{\prime}|s,a)\mathcal{R}(r|s,a)$) for the provided offline rollouts, comparing cases with and without belief adaptation.
+The ratios of these likelihoods are presented in Table 8.
+Results show that the offline rollouts, collected from real environments, are more likely under the adapted ensemble.
+This demonstrates that Bayesian belief adaptation serves as an effective calibration method for the learned world models.
+hc-med-expert hc-med-replay hc-medium hc-random
+0.7515 2.6315 2.2674 2.4561
+hp-med-expert hp-med-replay hp-medium hp-random
+14.064 4.2214 3.5913 1.7400
+wk-med-expert wk-med-replay wk-medium wk-random
+2.0898 3.8169 34.264 1.0341
+Table 8: Ratios of average transition likelihoods in offline data with and without belief adaptation across ensemble members.
+A ratio greater than 1 indicates that the adapted ensemble assigns higher likelihood to the true transitions.
+Furthermore, we evaluate the quality of imaginary rollouts generated by the learned models.
+Table 1 in the main content presents the prediction errors for next states and rewards in these rollouts.
+The results reveal that the adaptive ensemble achieves more accurate predictions in 10 out of 12 benchmarking tasks, indicating that belief adaptation leads to a more accurate model for planning.
+(a) Offline rollout
+(b) Imaginary rollout (at step 0)
+(c) Imaginary rollout (at step 22)
+(d) Imaginary rollout (at step 495)
+Figure 1: Belief adaptation during offline and imaginary rollouts.
+(a) shows the belief over twelve ensemble members, adapting to an offline trajectory of Hopper-med-expert.
+(b), (c), and (d) illustrate the belief changes during imaginary rollouts which start from different points of the offline trajectory shown in (a).
+As a qualitative analysis, we plot the belief adaptation over an offline trajectory (from $\mathcal{D}_{\mu}$) in Hopper-med-expert in Figure 1(a).
+Initially, all twelve ensemble members have the same belief.
+As the trajectory progresses, the beliefs of each member are updated based on the transition history, with the dominant model (the one with the highest belief) continuously changing.
+Figure 1(b)-1(d) further show the belief adaption during imaginary rollouts that start from different time steps of that offline trajectory.
+F.2 Comparison with Model-free Methods on D4RL MuJoCo
+Data Type Environment BA-MCTS (ours) CQL BEAR BRAC-v SAC BC
+random HalfCheetah 41.45 (1.40) 35.4 25.1 31.2 30.5 2.1
+random Hopper 31.42 (0.10) 10.8 11.4 12.2 11.3 1.6
+random Walker2d 21.54 (0.06) 7.0 7.3 1.9 4.1 9.8
+medium HalfCheetah 77.31 (1.40) 44.4 41.7 46.3 -4.3 36.1
+medium Hopper 103.9 (0.33) 86.6 52.1 31.1 0.8 29.0
+medium Walker2d 88.23 (0.78) 74.5 59.1 81.1 0.9 6.6
+med-replay HalfCheetah 71.24 (2.65) 46.2 38.6 47.7 -2.4 38.4
+med-replay Hopper 106.4 (0.53) 48.6 33.7 0.6 3.5 11.8
+med-replay Walker2d 91.42 (1.54) 32.6 19.2 0.9 1.9 11.3
+med-expert HalfCheetah 102.3 (2.27) 62.4 53.4 41.9 1.8 35.8
+med-expert Hopper 111.8 (0.82) 111 96.3 0.8 1.6 111.9
+med-expert Walker2d 116.0 (1.49) 98.7 40.1 81.6 -0.1 6.4
+Average Score 80.25 54.85 39.83 31.44 4.13 25.07
+Table 9: Comparisons between the proposed algorithms and model-free offline policy learning methods on the D4RL benchmark suite.
+Each value represents the normalized score, as proposed in Fu et al.
+(2020), of the policy trained by the corresponding algorithm.
+These scores are undiscounted returns normalized to approximately range between 0 and 100, where a score of 0 corresponds to a random policy and a score of 100 corresponds to an expert-level policy.
+For our algorithms, we report the average score of the final ten policy learning epochs and its standard deviation across three different random seeds.
+Here, we compare our algorithms with a series of model-free offline policy learning (Chen et al., 2024) methods in Table 9.
+We include SOTA model-free offline RL methods: CQL (Kumar et al., 2020), BEAR Kumar et al.
+(2019), and BRAC-v (Wu et al., 2019).
+Additionally, we show the performance of directly applying SAC or Behavioral Cloning (BC) (Chen et al., 2024) to the provided offline dataset in the last two columns.
+The mean performance of the baselines are taken from related works (Yu et al., 2020; Kidambi et al., 2020; Fu et al., 2020).
+Our algorithm shows significantly better performance, demonstrating the necessity of model-based learning in these environments.
+(a) hc-med-expert
+(b) hc-med-replay
+(c) hc-medium
+(d) hc-random
+(e) hp-med-expert
+(f) hp-med-replay
+(g) hp-medium
+(h) hp-random
+(i) wk-med-expert
+(j) wk-med-replay
+(k) wk-medium
+(l) wk-random
+Figure 2: Performance of Sampled EfficientZero on D4RL MuJoCo tasks.
+The results for HalfCheetah, Hopper, and Walker2d are presented in the three rows, respectively.
+Each subfigure depicts the change in undiscounted episodic return as a function of the number of training samples.
+Experiments are repeated three times with different random seeds, with the solid line representing the mean and the shaded area indicating the 95% confidence interval.
+For reference, the expert-level episodic returns (corresponding to scores of 100) for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively.
+F.3 Comprehensive Benchmarking on D4RL MuJoCo
+To firmly establish the state-of-the-art (SOTA) performance of our algorithm, this section provides a comprehensive comparison of BA-MCTS against 19 offline RL methods on the 12 D4RL MuJoCo tasks.
+To provide context, we briefly categorize the baselines:
+  • The model-based (MBRL) methods include classic approaches like MOPO (Yu et al., 2020), MOReL (Kidambi et al., 2020), and COMBO (Yu et al., 2021), as well as more recent/SOTA methods such as MOBILE (Sun et al., 2023), CBOP (Jeong et al., 2023), RAMBO (Rigter et al., 2022), and MAPLE (Chen et al., 2021b).
+  • The model-free (MFRL) methods cover a wide range of techniques.
+We include prominent SOTA methods such as the value-based algorithm CQL (Kumar et al., 2020), the implicit Q-learning method IQL (Kostrikov et al., 2021), ensemble-based methods like SAC-N and EDAC (An et al., 2021), the policy constraint method TD3+BC (Fujimoto and Gu, 2021), and the sequence modeling approach Decision Transformer (DT) (Chen et al., 2021a).
+We also compare against classic MFRL baselines, including BEAR (Kumar et al., 2019), BRAC-v (Wu et al., 2019), and the direct application of SAC and Behavioral Cloning (BC) (Chen et al., 2024).
+Finally, APE-V (Ghosh et al., 2022) is included as a conceptually related model-free method that also leverages the BAMDP framework.
+The results, including a statistical significance analysis using Cohen’s d (Cohen, 1988), are summarized in Table 3 in the main content.
+Note that MOReL, IQL, and the model-free offline RL baselines in Table 9 did not report standard deviations.
+Additionally, IQL and Decision Transformer (DT) were not evaluated on D4RL tasks with random data.
+As shown in Table 3, BA-MCTS achieves the highest average score among all 20 algorithms, both including and excluding the challenging ‘random‘ datasets.
+To quantify the significance of this improvement, we compute Cohen’s d against all baselines for which standard deviations are available.
+The results show that the Cohen’s d value is greater than 1.8 in all cases, and for most baselines, it is significantly larger (e.g., >> 3.5).
+This indicates that the performance improvement of our algorithm is not only substantial in magnitude but also statistically significant.
+This comprehensive comparison provides strong evidence that our proposed method establishes a new state-of-the-art in offline RL on this widely-used benchmark.
+F.4 Comparison with MuZero and its Variants
+Monte-Carlo Tree Search (MCTS) (Browne et al., 2012) has been successfully integrated with RL, as exemplified by AlphaZero (Silver et al., 2017) and MuZero (Schrittwieser et al., 2020).
+These methods have achieved superhuman performance in domains requiring highly sophisticated decision-making processes.
+AlphaZero relies on a given world model, whereas MuZero learns the world model and policy simultaneously by interacting with the environment.
+Although there have been various extensions of MuZero (Hubert et al., 2021; Schrittwieser et al., 2021; Ye et al., 2021; Antonoglou et al., 2022; Oren et al., 2022; Zhao et al., 2024), most are designed for online MBRL.
+According to Niu et al.
+(2023), applying MuZero to offline learning, especially for continuous control in highly stochastic environments, remains a significant open challenge.
+Our algorithm design differs from MuZero (and its variants) in several key aspects: (1) MuZero integrates model learning and policy training into a single stage, using a world model defined in a latent state space.
+Our algorithm separately learns a world model and then trains the policy on top of it, aligning with the widely-adopted offline MBRL framework.
+(2) MuZero employs a single latent world model (rather than an ensemble) and does not explicitly account for uncertainty in dynamics or reward predictions.
+This makes it particularly vulnerable to modeling errors in offline MBRL.
+(3) We introduce double progressive widening (Auger et al., 2013) and Bayes-adaptive planning into MCTS, resulting in a fundamentally different planning algorithm compared to the one used in MuZero.
+To evaluate the performance of MuZero-type methods on D4RL MuJoCo tasks, we use the open-source implementation and hyperparameter configurations of Sampled EfficientZero (Ye et al., 2021) provided by LightZero (Niu et al., 2023).
+Benchmarking results from LightZero indicate that Sampled EfficientZero achieves the best performance on (online) MuJoCo locomotion tasks compared to other MuZero variants.
+To adapt it for offline learning, we employ the reanalyse technique proposed by Schrittwieser et al.
+(2021).
+The evaluation results are presented in Figure 2.
+As shown, the performance is significantly worse compared to our methods listed in Figure 3, despite Sampled EfficientZero’s higher computational cost.
+For reference, the expert-level episodic returns (corresponding to scores of 100) for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively.
+In Table 10, we report the training time of our proposed algorithms and Sampled EfficientZero on the D4RL MuJoCo tasks.
+The experiments were conducted on a server with 40 Intel(R) Xeon(R) Gold 5215 CPUs and 4 Tesla V100-SXM2-32GB GPUs.
+Our search-based algorithm, BA-MCTS, requires considerably less training time than Sampled EfficientZero to achieve its superior performance.
+While BA-MCTS requires more computation than the non-search variant BA-MBRL, this extra cost is limited to the offline training phase; no MCTS is performed during deployment, ensuring real-time execution.
+Data Type Environment BA-MCTS -SL (ours) BA-MCTS (ours) BA-MBRL (ours) Sampled EfficientZero
+random HalfCheetah 11.2 (2.6) 14.6 (2.2) 1.2 (0.4) 54.8 (2.6)
+random Hopper 48.7 (2.4) 65.2 (1.9) 5.6 (0.8) 153.7 (19.4)
+random Walker2d 25.7 (2.0) 38.1 (1.1) 5.2 (1.1) 175.7 (28.2)
+medium HalfCheetah 12.6 (3.0) 15.3 (1.3) 1.9 (0.2) 54.7 (1.9)
+medium Hopper 18.9 (4.8) 67.5 (0.2) 1.8 (0.2) 70.5 (1.0)
+medium Walker2d 24.0 (1.8) 33.3 (4.0) 4.8 (1.0) 63.7 (1.5)
+med-replay HalfCheetah 24.7 (0.9) 22.4 (1.2) 2.1 (0.4) 54.8 (1.8)
+med-replay Hopper 17.9 (0.0) 12.7 (1.2) 5.4 (0.2) 126.6 (12.0)
+med-replay Walker2d 40.3 (7.2) 77.0 (2.4) 8.0 (0.3) 134.9 (4.2)
+med-expert HalfCheetah 32.6 (0.1) 12.6 (0.8) 4.7 (1.1) 55.4 (1.3)
+med-expert Hopper 61.5 (4.6) 76.9 (11.6) 5.3 (0.5) 66.1 (0.9)
+med-expert Walker2d 35.0 (6.5) 55.9 (1.8) 2.9 (0.9) 60.8 (1.8)
+Table 10: Training time (in hours) of our algorithms and Sampled EfficientZero for each evaluation task. Results are presented as the mean and standard deviation from three repeated experiments.
+The performance gap may be attributed to the challenges of applying MuZero’s design to the offline setting.
+World model learning is particularly difficult when the state-action space is vast but training data is limited.
+Sampled EfficientZero integrates model learning and policy training into a single stage, which can significantly increase the learning difficulty compared to our decoupled approach.
+F.5 Ablation Study on the Enhancement from Belief Adaptation and MCTS-based Planning
+Data Type Environment Optimized BA-MBRL BA-MCTS BA-MCTS-SL
+random HalfCheetah 31.7 32.76 (1.16) 36.23 (1.04) 29.20 (2.00)
+random Hopper 12.1 31.47 (0.03) 31.56 (0.12) 33.83 (0.10)
+random Walker2d 21.7 21.45 (0.53) 21.59 (0.32) 21.89 (0.07)
+medium HalfCheetah 45.7 56.54 (5.20) 75.84 (3.81) 70.47 (3.52)
+medium Hopper 69.3 98.25 (3.42) 96.70 (14.0) 97.75 (7.09)
+medium Walker2d 79.7 75.41 (4.17) 74.73 (3.25) 82.84 (1.85)
+med-replay HalfCheetah 58.0 62.50 (0.18) 65.45 (0.81) 61.16 (1.60)
+med-replay Hopper 90.8 93.91 (4.25) 101.8 (3.46) 106.3 (0.13)
+med-replay Walker2d 65.8 97.54 (1.93) 95.06 (2.11) 92.13 (5.13)
+med-expert HalfCheetah 104.2 90.52 (4.13) 76.16 (10.3) 80.53 (6.63)
+med-expert Hopper 105.8 107.8 (0.37) 108.3 (0.22) 112.2 (0.29)
+med-expert Walker2d 97.1 84.71 (0.87) 110.0 (1.74) 107.7 (0.82)
+Average Score 65.2 71.06 74.45 74.62
+Table 11: Performance comparison among our algorithm: BA-MCTS, its ablated versions: BA-MBRL and BA-MCTS-SL, as well as the baseline: Optimized (Lu et al., 2022).
+To ensure a principled ablation study, we reimplement BA-MCTS and its ablated variants using the codebase of Optimized, which systematically investigates design choices in offline MBRL.
+We make only minimal modifications to the code and hyperparameter settings.
+This approach ensures that any observed performance improvements are attributable to algorithmic design, rather than implementation differences or tuning tricks.
+For a principled ablation study, we reimplement BA-MCTS and its ablated variants based on the codebase of Optimized (Lu et al., 2022), which thoroughly explores design choices in offline MBRL, making minimal changes to the code and hyperparameter settings (see Appendix D).
+The comparison results are presented in Table 11555The performance of BA-MCTS decreases, relative to the results presented in Table 2, for two main reasons: (1) The original implementation in Table 2 is based on a different codebase than the one used in Optimized; and (2) for the ablation results shown in Table 11, we adopt the reward penalty design from Optimized to enable exact ablation comparisons, rather than using the one described in Eq.
+(5)., and the training curves of BA-MCTS and its ablated variants are shown in Figure 3.
+Firstly, by comparing the following two methods, any observed performance difference can be directly attributed to the benefits of belief adaptation in a BAMDP formulation.
+  • Optimized: A strong, non-adaptive offline model-based RL baseline.
+  • BA-MBRL: Our ablation variant, which is a minimal modification of Optimized, differing only by the integration of our BAMDP framework that dynamically adapts the belief over the model ensemble via Bayesian updates (Eq.
+(4)).
+As demonstrated, BA-MBRL outperforms the Optimized baseline on 8 out of 12 tasks and achieves a higher average score, providing direct evidence that Bayesian belief adaptation leads to improved final policies.
+Secondly, by comparing the following two methods, any observed performance difference can be directly attributed to the benefits of incorporating MCTS-based planning:
+  • BA-MBRL: Our ablation variant that incorporates the BAMDP belief adaptation framework but does not use MCTS. Its policy improvement relies solely on Soft Actor Critic.
+  • BA-MCTS: Our final algorithm, which builds upon BA-MBRL by integrating the Continuous BAMCP deep search planner.
+This substantial gain highlights the effectiveness of using Continuous BAMCP as a powerful policy improvement operator.
+While BA-MBRL relies on the implicit policy improvement provided by the temporal difference learning process, BA-MCTS leverages deep lookahead search to explicitly find a stronger, non-parametric policy ($\pi_{\text{ret}}$ returned by Algorithm 1) at each decision point.
+The actor-critic learner is then trained on the high-quality trajectories generated from this improved policy.
+This “RL+Search” paradigm, where the learner distills knowledge from a stronger planner, is the primary reason for the observed performance leap.
+This result strongly supports our claim that the deep search component is crucial for achieving state-of-the-art performance in the offline model-based setting.
+(a) hc-med-expert
+(b) hc-med-replay
+(c) hc-medium
+(d) hc-random
+(e) hp-med-expert
+(f) hp-med-replay
+(g) hp-medium
+(h) hp-random
+(i) wk-med-expert
+(j) wk-med-replay
+(k) wk-medium
+(l) wk-random
+Figure 3: Performance of our proposed algorithms on D4RL MuJoCo tasks in the ablation study.
+The results for HalfCheetah, Hopper, and Walker2d are presented in the three rows, respectively.
+Each subfigure depicts the change in the undiscounted episodic return as a function of training epochs.
+Experiments are repeated three times with different random seeds, with the solid line representing the mean and the shaded area indicating the 95% confidence interval.
+For reference, the expert-level episodic returns for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively.
+Note that the training epochs for each algorithm, as listed in Table 5, have been linearly scaled to 800 for better visualization.
+F.6 Comparison of Policy Distillation Mechanisms
+This section addresses the research question: How can the search outcomes (i.e., $\pi_{\text{ret}}$) be effectively used for policy improvement?
+We investigate this by comparing two different mechanisms for distilling the knowledge from the MCTS planner back into the parameterized policy network $\pi$.
+  • BA-MCTS: Our primary algorithm, which uses a policy gradient method (i.e., SAC) to update the policy network using trajectories generated by the planner.
+  • BA-MCTS-SL: An alternative variant where the policy network is updated via supervised learning (SL) to directly mimic the planner’s output distribution $\pi_{\text{ret}}$ by minimizing a cross-entropy loss.
+Performance Comparison: As shown in Table 11, BA-MCTS-SL performs similarly to BA-MCTS in terms of the final average score, validating the effectiveness of both policy update mechanisms.
+This indicates that both policy gradient and supervised learning are viable methods for distilling the results of the deep search.
+However, we note that BA-MCTS-SL can be less stable on certain complex tasks; for example, it struggles on the Walker2d environments, where a warm-up training phase (using BA-MBRL) is required to establish a good initial policy for the supervised learning to be effective.
+Training Stability and Robustness: On the other hand, a key advantage of the SL-based policy update is its training stability.
+As evident in the training plots (Figure 3), BA-MCTS-SL often exhibits much smoother learning curves compared to the other two algorithms, indicating greater robustness in model selection for deployment.
+Furthermore, as shown in our ablation study on the reward penalty (Appendix F.7), we found that the SL-based policy update is less sensitive to the degree of pessimism.
+F.7 Ablation Study on the Reward Penalty
+Data Type Environment BA-MCTS -SL BA-MCTS BA-MBRL BA-MCTS -SL ($\lambda=0$) BA-MCTS ($\lambda=0$) BA-MBRL ($\lambda=0$)
+random HalfCheetah 29.20 (2.00) 36.23 (1.04) 32.76 (1.16) 34.80 (1.39) 38.78 (1.65) 39.64 (2.86)
+random Hopper 33.83 (0.10) 31.56 (0.12) 31.47 (0.03) 9.16 (0.16) 7.44 (0.14) 6.97 (0.07)
+random Walker2d 21.89 (0.07) 21.59 (0.32) 21.45 (0.53) 17.53 (6.16) 21.53 (0.42) 21.41 (0.64)
+medium HalfCheetah 70.47 (3.52) 75.84 (3.81) 56.54 (5.20) 61.64 (4.58) 60.84 (2.00) 41.49 (2.29)
+medium Hopper 97.75 (7.09) 96.70 (14.0) 98.25 (3.42) 102.8 (2.29) 104.4 (1.88) 93.68 (11.4)
+medium Walker2d 82.24 (1.85) 74.73 (3.25) 75.41 (4.17) 82.61 (0.86) 57.01 (7.24) 57.97 (15.4)
+med-replay HalfCheetah 61.16 (1.60) 65.45 (0.81) 62.50 (0.18) 42.10 (2.85) 36.65 (2.39) 44.03 (7.35)
+med-replay Hopper 106.3 (0.13) 101.8 (3.46) 93.91 (4.25) 107.9 (0.07) 84.11 (2.97) 91.81 (11.5)
+med-replay Walker2d 92.13 (5.13) 95.06 (2.11) 97.54 (1.93) 88.61 (5.21) 97.33 (3.51) 98.19 (1.23)
+med-expert HalfCheetah 80.53 (6.63) 76.16 (10.3) 90.52 (4.13) 51.76 (5.31) 26.60 (1.46) 29.88 (2.28)
+med-expert Hopper 112.2 (0.29) 108.3 (0.22) 107.8 (0.37) 106.8 (6.34) 81.76 (6.45) 86.79 (18.7)
+med-expert Walker2d 107.7 (0.82) 110.0 (1.74) 84.71 (0.87) 110.8 (1.72) 110.2 (0.91) 53.35 (38.0)
+Average Score 74.62 74.45 71.06 68.04 60.55 55.43
+Table 12: Comparison of the proposed algorithms with their corresponding versions without the reward penalty (i.e., $\lambda=0$).
+The definitions of the scores in this table are consistent with those in Table 9.
+To demonstrate the necessity of incorporating the reward penalty in offline MBRL, we conduct an ablation study by setting $\lambda$ in Eq.
+(5) to 0, resulting in ablated versions of our proposed three algorithms.
+The results are presented in Table 12.
+First, the average performance of the algorithms with the reward penalty is consistently better, demonstrating the importance of using reward penalties in offline MBRL to prevent the overexploitation of the learned world models (which can be inaccurate).
+Second, the supervised-learning-based algorithm (i.e., BA-MCTS-SL ($\lambda=0$)) is less affected by the absence of the reward penalty, compared to the policy-gradient-based methods.
+Notably, BA-MCTS-SL ($\lambda=0$) and BA-MCTS-SL achieve comparable performance in the Hopper and Walker2d tasks.
+This shows an additional advantage of BA-MCTS-SL – its reduced sensitivity to model inaccuracies.
+Lastly, there are instances where superior performance is achieved with $\lambda$ set to 0.
+For example, BA-MCTS-SL ($\lambda=0$) performs better than BA-MCTS-SL in 5 out of 12 tasks.
+This suggests that the performance of our algorithms in Tables 11 could be further improved by adjusting hyperparameters such as $\lambda$.
+Appendix G Statement on the Use of Large Language Models
+In the spirit of transparency and in accordance with the ICLR disclosure policy, we detail the use of Large Language Models (LLMs) in the preparation of this manuscript.
+LLMs were utilized as an assistive tool in two main capacities: language refinement and technical support during the formalization of proofs. The specific roles were:
+  • Language and Readability Enhancement: We employed an LLM to polish the manuscript’s language.
+This included improving grammar, rephrasing sentences for better clarity, and ensuring a consistent and professional tone throughout the paper.
+  • Assistance with Theoretical Proofs: An LLM provided support in the technical writing and verification of our theoretical proofs. This assistance was twofold:
+    1. 1.
+      Generating source code for complex mathematical equations.
+    2. 2.
+      Acting as a verification aid by reviewing the logical steps in our derivations to help identify potential inconsistencies or errors.
+It is crucial to state that the LLM did not contribute to the original ideation, conceptual development, or the core derivations of the proofs themselves.
+The intellectual ownership and the scientific reasoning behind all theoretical results remain entirely with the authors.
+The authors have meticulously reviewed, validated, and take full and final responsibility for all content, including the correctness and rigor of the proofs that were checked with LLM assistance.